### v1

# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

Abstract
We study a new class of MDPs that employs multinomial logit function approximation to ensure valid probability distributions over the state space. Despite its benefits, introducing the non-linear function raises significant challenges in both computational and statistical efficiency. The best-known method of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity, $d$ is the feature space dimension, $H$ is the episode length, and $K$ is the number of episodes. While this result attains the same rate in $K$ as the linear cases, the method requires storing all historical data and suffers from an $\mathcal{O}(K)$ computation cost per episode. Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases. In this work, we first address the computational issue by proposing an online algorithm that achieves the same regret with only $\mathcal{O}(1)$ computation cost. Then, we design two algorithms that leverage local information to enhance statistical efficiency. They not only maintain an $\mathcal{O}(1)$ computation cost per episode but achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ respectively. Finally, we establish a lower bound, justifying the optimality of our results in $d$ and $K$. To the best of our knowledge, this is the first work that achieves almost the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation for reinforcement learning.
1 Introduction
Reinforcement Learning (RL) with function approximation has achieved great success in many applications with large state and action space, such as games (Mnih et al., 2013), robotics (Kober et al., 2013), large language models (Ouyang et al., 2022), etc. Therefore, establishing the theoretical foundation for RL with function approximation is of great importance. Recently, there have been many efforts devoted to understanding the linear function approximation due to its significance and simplicity, yielding fruitful results (Yang and Wang, 2019; Jin et al., 2020; Ayoub et al., 2020).
While these works make important steps toward understanding reinforcement learning with function approximation, there are still some challenges to be solved. In RL with linear function approximation, transitions are presumed to be linear in specified feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs. Here $\mathbb{P}(s^{\prime}|s,a)$ is the probability from state $s$ to state $s^{\prime}$ taking action $a$, $\phi(s^{\prime}|s,a)$ and $\phi(s,a)$ are feature mappings, $\theta^{*}$ and $\mu^{*}(s^{\prime})$ are unknown parameters. However, a transition function is a probability distribution over states, i.e., the magnitude falls within the range of $[0,1]$, and the aggregate probability equals $1$. Thus, for certain feature mappings, linear transitions may not yield a valid probability distribution with arbitrary $\theta$ and $\mu$. In general, an algorithm designed for linear MDPs could break down entirely if the underlying MDP is not linear as noted by Jin et al. (2020). While there are some works exploring generalized linear function approximation (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Du et al., 2021; Foster et al., 2021), they make assumptions over value functions rather than transitions, hence do not tackle this challenge.
Table 1: Comparison between previous work and ours in terms of the regret bound, computation cost and storage cost. Here $\kappa$ and $\kappa^{*}$ are exponentially small problem-dependent quantities, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes. The computational cost and storage cost per episode indicate the dependence on episode count $k$.
Reference Regret Comp. Stor. MDP model
Hwang and Oh (2023) $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ $\mathcal{O}(k)$ $\mathcal{O}(k)$ homogeneous
UCRL-MNL-OL (Theorem 1) $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
UCRL-MNL-LL (Theorem 2) $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
UCRL-MNL-LL+ (Theorem 3) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
Lower Bound  (Corollary 1) $\Omega(dH\sqrt{K\kappa^{*}})$ / / infinite action space
Towards addressing the limitation of linear function approximation, a new class of MDPs that utilizes multinomial logit (MNL) function approximation has been proposed by Hwang and Oh (2023) recently. Despite its benefits, introducing non-linear functions raises significant challenges in both computational and statistical efficiency. Specifically, the best-known method of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret. While matching the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ as linear function approximation in terms of $K$ (Zhou et al., 2021; He et al., 2023), their method requires storing all historical data and the computational cost per episode grows linearly with the episode count, which is inefficient. Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases. Thus, a natural question arises:
Can we achieve (almost) the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation for RL?
In this paper, we answer this question affirmatively by making significant advancements in both computational and statistical efficiency for RL with MNL function approximation, almost matching the results of linear function approximation. Our contributions can be summarized as follows:
  • We first propose the UCRL-MNL-OL algorithm based on a variant of Online Newton Step, attaining an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. It matches the best-known regret by Hwang and Oh (2023), yet achieves the same computational and storage efficiency as linear function approximation (Jin et al., 2020; Zhou et al., 2021).
  • As the quantity $\kappa$ can be exponentially small, we propose two algorithms UCRL-MNL-LL and UCRL-MNL-LL+, which leverage local information to improve the statistical efficiency. They not only maintain $\mathcal{O}(1)$ computational and storage costs but also achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, respectively. The latter result moves the $\kappa$ dependence to the lower-order term which is independent of $K$ and the higher-order term matches the optimal regret $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023) in both $d$ and $K$, differing only by an $\mathcal{O}(H^{1/2})$ factor.
  • We establish the first lower bound for MDPs with MNL function approximation by employing a reduction to the logistic bandits. We prove a lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for the infinite action space setting. Though this does not constitute a strict lower bound for the finite actions case studied in this work, it justifies that our results may be optimal in $d$ and $K$.
Table 1 shows comparisons between previous works and ours. From a technical perspective, inspired by recent advances in multinomial logistic bandits (Zhang and Sugiyama, 2023), we estimate the parameter by running an online algorithm instead of performing maximum likelihood estimation over all historical data. Differently, we need to design value functions carefully that utilize local information to address the unique challenges of MDPs, which is highly non-trivial.
Organization.
We first introduce the related work in Section 2 and present the setup in Section 3. Then, we introduce an algorithm with constant computational and storage costs in Section 4. Next, we present two improved algorithms in Section 5. Finally, we establish the lower bound in Section 6. Section 7 concludes the paper. Due to space constraints, we defer all proofs to the appendixes.
Notations.
We denote by $[n]$ the set $\{1,\ldots,n\}$ and define $[x]_{[a,b]}=\min(\max(x,a),b)$. For $X,Y\in\mathbb{R}^{d\times d}$, $X\succeq Y$ means $X-Y$ is positive semi-definite. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, denote $\lVert\mathbf{x}\rVert_{A}=\sqrt{\mathbf{x}^{\top}A\mathbf{x}}$. The $\widetilde{\mathcal{O}}(\cdot)$ hides all polylogarithmic factors.
2 Related Work
In this section, we review related works, both in terms of settings and techniques.
RL with Generalized Linear Function Approximation.
There are recent efforts devoted to investigating function approximation beyond the linear models. Wang et al. (2021) investigated RL with generalized linear function approximation. Notably, unlike our approach which models transitions using a generalized linear model, they apply this approximation directly to the value function. Another line of works (Chowdhury et al., 2021; Li et al., 2022; Ouhamma et al., 2023) has studied RL with exponential function approximation and also aimed to ensure that transitions constitute valid probability distributions. The MDP model can be viewed as an extension of bilinear MDPs in their work while our setting extends linear mixture MDPs. These studies are complementary to ours and not directly comparable. Moreover, these works also enter the computational and statistical challenges arising from non-linear function approximation that remain to be addressed. The most relevant work to ours is the recent work by Hwang and Oh (2023), which firstly explored a similar setting to ours, where the transition is modeled using a multinomial logit model. We significantly improve their results by providing computationally and statistically efficient algorithms.
RL with General Function Approximation.
There have also been some works studies reinforcement learning with general function approximation. Russo and Roy (2013) and Osband and Roy (2014) initiated the study on the minimal structural assumptions that render sample-efficient learning by proposing a structural condition called Eluder dimension. Recently, several works have investigated different conditions for sample-efficient interactive learning, such as Bellman Eluder (BE) dimension (Jin et al., 2021), Bilinear classes (Du et al., 2021), Decision-Estimation Coefficient (DEC) (Foster et al., 2021), and Admissible Bellman Characterization (ABC) (Chen et al., 2023). A notable difference is that they impose assumptions on the value functions while we study function approximation on the transitions. Moreover, the goal of these works is to study the conditions for sample-efficient reinforcement learning, but not focus on the computational efficiency.
Multinomial Logit (MNL) Contextual Bandits.
As our work shares similar techniques with MNL contextual bandits to some extent, we briefly review works on this topic. The pioneering study by Cheung and Simchi-Levi (2017) introduced an algorithm that achieves a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{K})$, where $K$ denotes the number of rounds. This result was further enhanced by subsequent studies (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). The recent work by Périvier and Goyal (2022) has significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ in the uniform revenue setting. Most of the above methods require storing all historical data and are computationally inefficient. A concurrent work with ours by Lee and Oh (2024) proposed an algorithm with constant computational and storage costs inspired by recent advancements in multinomial logistic bandits (Zhang and Sugiyama, 2023) and achieved optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K})$ and $\widetilde{\mathcal{O}}(d\sqrt{K})$ under uniform and nonuniform rewards respectively. However, although the underlying models of MNL contextual bandits and MNL mixture MDPs are similar, the challenges they pose differ significantly and the techniques developed for MNL contextual bandits can not be directly applied to the MNL mixture MDPs. Specifically, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize expected revenue, whereas in MNL mixture MDPs, the goal is to maximize cumulative rewards Therefore, while we utilize similar parameter estimation techniques as those in MNL contextual bandits, we need to construct value functions to address the unique challenges of MDPs, which is highly non-trivial and needs a careful design.
3 Problem Setup
In this section, we present the problem setup of RL with multinomial logit function approximation.
Inhomogeneous, Episodic MDPs.
An inhomogeneous, episodic MDP instance can be denoted by a tuple $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the length of each episode, $\mathbb{P}_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the transition kernel with $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)$ is being the probability of transferring to state $s^{\prime}$ from state $s$ and taking action $a$ at stage $h$, $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the deterministic reward function. A policy $\pi=\{\pi_{h}\}_{h=1}^{H}$ is a collection of mapping $\pi_{h}$, where each $\pi_{h}:\mathcal{S}\to\Delta(\mathcal{A})$ is a function maps a state $s$ to distributions over $\mathcal{A}$ at stage $h$. For any policy $\pi$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the action-value function $Q_{h}^{\pi}$ and value function $V_{h}^{\pi}$ as follows:
$Q_{h}^{\pi}(s,a)=\mathbb{E}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\left(s_{h^{\prime}},a_{h^{\prime}}\right)\Big{|}s_{h}=s,a_{h}=a\right],\quad V_{h}^{\pi}(s)=\mathbb{E}_{a\sim\pi_{h}(\cdot\mid s)}\left[Q_{h}^{\pi}(s,a)\right],$
where the expectation of $Q_{h}^{\pi}$ is taken over the randomness of the transition $\mathbb{P}$ and policy $\pi$. The optimal value function $V_{h}^{*}$ and action-value function $Q_{h}^{*}$ given by $V_{h}^{*}(s)=\sup_{\pi}V_{h}^{\pi}(s)$ and $Q_{h}^{*}(s,a)=\sup_{\pi}Q_{h}^{\pi}(s,a)$. For any function $V:\mathcal{S}\to\mathbb{R}$, we define $[\mathbb{P}_{h}V](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot\mid s,a)}V(s^{\prime})$. The Bellman equation for policy $\pi$ and the Bellman optimality equation are given by
$Q_{h}^{\pi}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{\pi}\right](s,a),\quad Q_{h}^{*}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{*}\right](s,a).$
Learning Protocol.
In the online MDP setting, the learner interacts with the environment without the knowledge of the transition $\{\mathbb{P}_{h}\}_{h=1}^{H}$. For the sake of simplicity, we assume the reward $\{r_{h}\}_{h=1}^{H}$ is known. The interaction proceeds in $K$ episodes. At the beginning of each episode $k$, the learner chooses a policy $\pi_{k}=\{\pi_{k,h}\}_{h=1}^{H}$. At each stage $h\in[H]$, starting from the initial state $s_{k,1}$, the learner observes the state $s_{k,h}$, chooses an action $a_{k,h}$ sampled from $\pi_{k,h}(\cdot\>|\>s_{k,h})$, obtains reward $r_{h}(s_{k,h},a_{k,h})$ and transits to the next state $s_{k,h+1}\sim\mathbb{P}_{h}(\cdot\>|\>s_{k,h},a_{k,h})$ for $h\in[H]$. The episode ends when $s_{H+1}$ is reached. The goal of the learner is to minimize regret, defined as the difference between the cumulative reward of the optimal policy and the learner’s policy, namely,
$\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1}).$
Multinomial Logit (MNL) Mixture MDPs.
Although significant advances have been achieved for MDPs with linear function approximation, Hwang and Oh (2023) show that there exists a set of features such that no linear transition model (including bilinear and low-rank MDPs) can induce a valid probability distribution over the state space, which limits the expressiveness of function approximation. To overcome this limitation, they propose a new class of MDPs with multinomial logit function approximation. We introduce the formal definition of MNL mixture MDPs below.
Definition 1 (Reachable States).
For any $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, we define the “reachable states” as the set of states that can be reached from state $s$ taking action $a$ at stage $h$ within a single transition, i.e., $\mathcal{S}_{h,s,a}\triangleq\{s^{\prime}\in\mathcal{S}\mid\mathbb{P}_{h}(s^{\prime}\>|\>s,a)>0\}$. Also, we define $S_{h,s,a}\triangleq|\mathcal{S}_{h,s,a}|$ and further define $U\triangleq\max_{(h,s,a)}S_{h,s,a}$ the maximum number of reachable states.
Definition 2 (MNL Mixture MDP).
An MDP instance $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$ is called an inhomogeneous, episodic $B$-bounded MNL mixture MDP if there exist a known feature mapping $\phi(s^{\prime}\>|\>s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$ with $\lVert\phi(s^{\prime}\>|\>s,a)\rVert_{2}\leq 1$ and unknown vectors $\{\theta_{h}^{*}\}_{h=1}^{H}\in\Theta$ with $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime}\in\mathcal{S}_{h,s,a}$, it holds that
$\mathbb{P}_{h}(s^{\prime}\>|\>s,a)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta_{h}^{*})}{\sum_{\widetilde{s}\in\mathcal{S}_{h,s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta_{h}^{*})}.$
For any $\theta\in\mathbb{R}^{d}$, we define the induced transition as $p_{s,a}^{s^{\prime}}(\theta)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta)}{\sum_{\widetilde{s}\in\mathcal{S}_{s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta)}$. Next, we introduce a key assumption for this problem as in Hwang and Oh (2023).
Assumption 1.
There exists $10<\kappa\leq\kappa^{*}<$ 1 such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime},s^{\prime\prime}\in\mathcal{S}_{h,s,a}$, it holds that $\inf_{\theta\in\Theta}p_{s,a}^{s^{\prime}}(\theta)p_{s,a}^{s^{\prime\prime}}(\theta)\geq\kappa$ and $p_{s,a}^{s^{\prime}}(\theta^{*})p_{s,a}^{s^{\prime\prime}}(\theta^{*})>\kappa^{*}$.
Assumption 1 is equivalent to the standard assumption in generalized linear model literature (Filippi et al., 2010; Wang et al., 2021) and logistic bandits literature (Faury et al., 2020; Abeille et al., 2021; Zhang and Sugiyama, 2023) to guarantee the Hessian matrix is non-singular.
4 Efficient Algorithm with Online Update
In this part, we design an algorithm that achieves $\mathcal{O}(1)$ computational and storage costs.
Efficiency Concern in Previous Methods.
While Hwang and Oh (2023) made the first step and proposed the UCRL-MNL algorithm with $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, the computational and storage complexity of their algorithm scales linearly with the episode count, which is expensive. Specifically, for the trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$ at episode $k$, we define the variable: $y_{k,h}\in\{0,1\}^{N_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\operatorname{\mathds{1}}{(s_{k,h+1}=s^{\prime})}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $N_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$. Furthermore, we denote by $p_{k,h}^{s^{\prime}}(\theta)=p_{s_{k,h},a_{k,h}}^{s^{\prime}}(\theta)$. Then $y_{k,h}$ is a sample from the following multinomial distribution:
$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{N_{k,h}}}(\theta^{*})]).$
where the parameter $1$ indicates that $y_{k,h}$ is a single-trial sample. We define the noise $\epsilon_{k,h}^{s^{\prime}}=y_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\theta_{h}^{*})$. It is clear that $\epsilon_{k,h}\in[-1,1]^{N_{k,h}}$, $\mathbb{E}[\epsilon_{k,h}]=\mathbf{0}$ and $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\epsilon_{i,h}^{s^{\prime}}=0$.
The UCRL-MNL algorithm estimates the unknown transition parameter $\theta_{h}^{*}$ using the regularized maximum likelihood estimation (MLE). Specifically, based on $y_{k,h}$, the estimator $\theta_{k,h}$ is defined as the solution to the following optimization problem:
$\theta_{k,h}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}-y_{i,h}^{s^{\prime}}\log p_{i,h}^{s^{\prime}}(\theta)+\frac{\lambda}{2}\lVert\theta\rVert_{2}^{2}.$ (1)
In episode $k$, the estimator $\theta_{k,h}$ in (1) is computed using all samples collected in previous episodes, resulting in the computational and storage cost per episode that scales linearly with the episode count, which is expensive in practice.
To address this issue, we introduce a novel algorithm named UCRL-MNL with Online Learning (UCRL-MNL-OL), which achieves both the same regret with constant computational and storage cost. At a high level, our algorithm can be divided into two phases: (i) efficient online estimation, and (ii) efficient optimistic value function construction. We introduce the details in the following.
Efficient Online Estimation.
Instead of using all historical data, our algorithm updates the estimator in an efficient online manner. This is achieved by efficiently exploiting particular properties of the MNL model. The key observation is that the negative log-likelihood function is exponentially concave over the bounded domain, which motivates us to apply a variant of online Newton step (Hazan et al., 2007), in particular, inspired by the works (Hazan et al., 2014; Zhang et al., 2016; Oh and Iyengar, 2021), which focus on logistic bandits and multinomial logit contextual bandits.
First, we define pe-episode loss function $f_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}},$
$H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$ (2)
We introduce the following key property of the loss function $f_{k,h}(\theta)$ by a second-order Taylor expansion, whose proof can be found in Lemma 5 in Appendix A.
Property 1.
(i). For any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, $\exists~{}\bar{\theta}=\nu\theta_{1}+(1-\nu)\theta_{2}$ such that
$f_{k,h}(\theta_{2})=f_{k,h}(\theta_{1})+\langle g_{k,h}(\theta_{1}),\theta_{2}-\theta_{1}\rangle+\frac{1}{2}\lVert\theta_{2}-\theta_{1}\rVert_{H_{k,h}(\bar{\theta})}^{2}.$
(ii). For any $\theta\in\Theta$, it holds that
$H_{k,h}(\theta)\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\triangleq\kappa W_{k,h},\forall\theta\in\Theta.$ (3)
This implies that $f_{k,h}(\theta)$ is exponentially concave over the bounded domain $\Theta$, which enables us to apply the online Newton step to construct an efficient estimator. Specifically, we construct the estimator $\widehat{\theta}_{k+1,h}$ by solving the problem:
$\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2},$ (4)
where $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ is the feature covariance matrix. This estimator can be computed efficiently in an online manner without the need to store all historical data. Then, we construct the confidence set by the following lemma.
Lemma 1.
For any $\delta\in(0,1)$, define the confidence set as
$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k,h}}\leq\sqrt{\lambda}B+\sqrt{\frac{8}{\kappa}d\log\left(1+\frac{kUH}{d\lambda\delta}\right)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$
Then, we have $\Pr[\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
Remark 1.
Notably, a similar confidence set is achieved by Hwang and Oh (2023) by using the MLE defined in (1). In contrast, we obtain the same results by using an online estimator which only suffers $\mathcal{O}(1)$ computation and storage cost per round. Besides, we identify a technical issue in their analysis. Specifically, they bound the confidence set using the self-normalized concentration for vector-valued martingales in Lemma 15. However, it is crucial to recognize that the noise is not independent and satisfies $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\epsilon_{i,h}^{s^{\prime}}=0$ since the learner visits each stage $h$ exactly once per episode. Thus, the noise $\epsilon_{i,h}^{s^{\prime}}$ becomes deterministic and non-zero given the remaining noise $\epsilon_{i,h}^{s^{\prime\prime}}$ for $s^{\prime\prime}\neq s^{\prime}$. This contravenes the zero-mean sub-Gaussian condition in Lemma 15. We observe similar oversights also appear in the works on multinomial logit contextual bandits (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). To the best of our knowledge, this issue has not been explicitly identified in previous studies. We note that this issue can be resolved with only slight modifications in constant factors by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al. (2024), a simplified version of Theorem C.6 in Périvier and Goyal (2022).
Algorithm 1 UCRL-MNL-OL
0:  Regularization parameter $\lambda$, confidence width $\widehat{\beta}_{k}$, confidence parameter $\delta$.
1:  Initialization:  $\widehat{\Sigma}_{1,h}=\lambda I,\widehat{\theta}_{1,h}=\mathbf{0}$ for all $h\in[H]$.
2:  for $k=1,\ldots,K$ do
3:     Compute $\widehat{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (6).
4:     for $h=1,\ldots,H$ do
5:        Observe current state $s_{k,h}$, select action $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
6:        Set $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$.
7:        Compute $\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ as (4).
8:     end for
9:  end for
Efficient Optimistic Value Function Construction.
Given the confidence set $\widehat{\mathcal{C}}_{k,h}$, the most direct way is to adopt the principle of “optimism in the face of uncertainty” and construct the optimistic value function as the maximum expected reward over the confidence set. Specifically, define the value function $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\max_{\theta\in\widehat{\mathcal{C}}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})\bigg{]}_{[0,H]},\widehat{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s,a).$ (5)
However, this construction is not efficient as it requires solving a maximization problem over the confidence set. Actually, the algorithm only requires the estimated action-value function to be an upper bound for the true value functions. Thus, we can use a closed-form confidence bound instead of computing the maximal value over the confidence set. To this end, we present the following lemma, which enables us to construct the optimistic value function efficiently.
Lemma 2.
Suppose Lemma 1 holds. For any function $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, with probability at least $1-\delta$, it holds that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}.$
Based on this lemma, we can replace the maximization problem in (5) with closed-form confidence bound and construct the optimistic value function $\widehat{Q}_{k,h}(s,a)$ as
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\widehat{\Sigma}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (6)
At state $s_{k,h}$, we chooses $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$. Algorithm 1 presents the detailed procedure. We show that our algorithm enjoys the following regret guarantee.
Theorem 1.
Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\widehat{\beta}_{k}$ as in Lemma 1 and $\lambda=1$, with probability at least $1-\delta$, Algorithm 1 ensures the following guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K}).$
Remark 2.
Our algorithm attains the same regret of Hwang and Oh (2023), yet for the more general inhomogeneous MDPs. More importantly, our algorithm only requires $\mathcal{O}(1)$ computational and storage costs, matching the computational efficiency of MDPs with linear function approximation (Ayoub et al., 2020; Jin et al., 2020; Zhou et al., 2021; He et al., 2023).
5 Improved Algorithm by Local Learning
While the UCRL-MNL-OL algorithm enjoys favorable computational efficiency, it suffers from a dependence on $\kappa^{-1}$ in the regret. By Assumption 1, the quantity $\kappa$ may be exponentially small, exhibiting an exponential dependence on the radius of the parameter set and linear dependence on the number of reachable states, i.e., $\kappa^{-1}=\mathcal{O}((U\exp(B))^{2})$. Thus, there is a significant gap between MNL function approximation and linear function approximation. To bridge this gap, we propose new algorithms that leverage local information to enhance the dependence on $\kappa$ significantly.
We first analyze where the dependence on $\kappa$ comes from. By the analysis of Theorem 1, we can see that the regret of the algorithm can be upper-bounded as follows,
$\operatorname*{Reg}(K)\leq 2H\widehat{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}.$ (7)
where the first term is the overestimate of the value function and the second term is the martingale difference sequence due to the stochastic transition. The second term can be upper-bounded by the Azuma-Hoeffding inequality which is independent of $\kappa$. Thus, the $\kappa$-dependence comes from the first term. By the construction of the confidence set in Lemma 1, the dependence on $\widehat{\beta}_{k}$ in the first term leads to a $\kappa^{-1/2}$ dependence. By the configuration of the covariance matrix $\widehat{\Sigma}_{k,h}$ in (4), the dependence on $\widehat{\Sigma}_{k,h}$ also leads to a $\kappa^{-1/2}$ dependence. Thus, we need to design a local learning algorithm that can adaptively use local information to enhance the dependence on these two terms.
5.1 Improved Online Estimation
In Section 4, we show the Hessian matrix $H_{k,h}$ is lower bounded by a positive definite matrix $\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ uniformly over the parameter domain in (3). This lower bound is not tight as the Hessian matrix can be much larger at certain regions rather than the global minimum. This motivates us to design a local learning algorithm that can adaptively use local information.
Inspired by recent advances in multinomial logistic bandit (Zhang and Sugiyama, 2023) and multinomial logit contextual bandit (Lee and Oh, 2024), we run an online algorithm to estimate the parameter $\theta_{h}^{*}$ similar to (4). Differently, we use the local Hessian matrix ~$\widetilde{\mathcal{H}}_{k,h}$ to update the estimation instead of the global lower bound. Specifically, we estimate parameter $\widetilde{\theta}_{k,h}$ as follows:
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$ (8)
where $\eta$ is the step size, $\mathcal{H}_{k,h}=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d}$ and ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$. Note both $\mathcal{H}_{k,h}$ and ~$\widetilde{\mathcal{H}}_{k,h}$ can be updated incrementally and the optimization problem in (8) can be solved efficiently. This step can be solved by the following two-step update.
$\theta^{\prime}_{k+1,h}=\theta_{k,h}-\eta\mathcal{H}_{k,h}^{-1}g_{k,h}(\theta_{k,h}),\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$ (9)
Based on this estimator, we can construct the $\kappa$-independent confidence set as follows.
Lemma 3.
For any $\delta\in(0,1)$, define the confidence set as
$\widehat{\mathcal{C}}_{k,h}=\big{\{}\theta\in\Theta\mid\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}\leq\widetilde{\beta}_{k}\big{\}}$
with $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$. Then, $\Pr[\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
Based on Lemma 3, we can construct the optimistic value function similar to (6) as
$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\mathcal{H}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (10)
The detailed algorithm is presented in UCRL-MNL-LL in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL by the following theorem.
Theorem 2.
Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, and $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, with probability at least $1-\delta$, UCRL-MNL-LL algorithm ensures the following regret guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K}).$
Compared with the result in Theorem 1, Theorem 2 enhances the dependence on $\kappa$ from $\kappa^{-1}$ to $\kappa^{-1/2}$. This is achieved by utilizing local information of the Hessian matrix.
Algorithm 2 UCRL-MNL-LL &\&& UCRL-MNL-LL+
0:  Step size $\eta$, Regularization parameter $\lambda$, confidence width $\widetilde{\beta}_{k}$, confidence parameter $\delta$.
1:  Initialization: $\mathcal{H}_{1,h}=\lambda I,\widehat{\theta}_{1,h}=\mathbf{0}$ for all $h\in[H]$.
2:  for $k=1,\ldots,K$ do
3:     UCRL-MNL-LL: Compute $\bar{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (10).
4:     UCRL-MNL-LL+: Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (12).
5:     for $h=1,\ldots,H$ do
6:        UCRL-MNL-LL: Observe $s_{k,h}$, select $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\bar{Q}_{k,h}(s_{k,h},a)$.
7:        UCRL-MNL-LL+: Observe $s_{k,h}$, select $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{Q}_{k,h}(s_{k,h},a)$.
8:        Update ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$.
9:        Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}$ as (8).
10:        Update $\mathcal{H}_{k,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
11:     end for
12:  end for
5.2 Improved Optimistic Value Function Construction
In Section 5.1, we improve the dependence on $\kappa$ by constructing a tighter $\kappa$-independent confidence set. However, the regret of UCRL-MNL-LL in Theorem 2 still have a dependence on $\kappa^{-1/2}$. In this part, we design an improved algorithm that removes the dependence on $\kappa$ in the main term.
By a similar analysis as (7), the regret of UCRL-MNL-LL can be bounded as:
$\operatorname*{Reg}(K)\leq 2H\widetilde{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}.$ (11)
Though $\widetilde{\beta}_{K}$ now is independent of $\kappa$ and $\mathcal{H}_{k,h}^{-1}$ preserves the local information in (11), the norm $\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is still in a global manner due to the maximum operation, leading to a $\kappa^{-1/2}$ dependence. To address this issue, we propose a new construction of the optimistic value function through a more refined analysis. Specifically, we employ a second-order Taylor expansion to more accurately bound the value difference arising from the estimation error in the following lemma.
Lemma 4.
Suppose Lemma 3 holds. For any function $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, with probability at least $1-\delta$, it holds that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}},$
where
$\epsilon_{s,a}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\quad\epsilon_{s,a}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
Based on Lemma 4, we construct the optimistic value function as follows:
$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]}.$ (12)
The detailed algorithm is presented in UCRL-MNL-LL+ in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL+ in the following theorem.
Theorem 3.
Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$ and $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, with probability at least $1-\delta$, UCRL-MNL-LL+ algorithm ensures the following regret guarantee
${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2}\big{)}.$
Remark 3.
Theorem 3 moves the exponentially small $\kappa$-dependence to the lower-order term, which is independent of $K$. Compared with the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023), the higher-order term only differs by a factor of $H^{1/2}$, almost matching the same computational and statistical efficiency simultaneously.
6 Lower Bound
In this section, we establish the first lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and logistic bandits.
Consider the following logistic bandit problem (Faury et al., 2020): at each round $k\in[K]$, the learner selects an action $x_{k}\in\mathcal{X}$ and receives a reward $r_{k}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d}\mid\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter. The learner aims to minimize the regret: ${\operatorname*{Reg}}^{\mathtt{LB}}(K)=\max_{x\in\mathcal{X}}\sum_{k=1}^{K}\mu(x^{\top}\theta^{*})-\sum_{k=1}^{K}\mu(x_{k}^{\top}\theta^{*})$.
Theorem 4.
For any logistic bandit $\mathcal{B}$, there exists an MNL mixture MDP $\mathcal{M}$ such that learning $\mathcal{M}$ is as hard as learning $H/2$ independent instances of $\mathcal{B}$ simultaneously.
Corollary 1.
For any problem instance $\{\theta_{h}^{*}\}_{h=1}^{H}$ and for $K\geq d^{2}\kappa^{*}$, there exists an MNL mixture MDP with infinite action space such that $\operatorname*{Reg}(K)\geq\Omega(dH\sqrt{K\kappa^{*}})$.
Remark 4.
Corollary 1 can be proved by combining Theorem 4 and the $\Omega(d\sqrt{K\kappa^{*}})$ lower bound for logistic bandits with infinite arms by Abeille et al. (2021). To our knowledge, a lower bound for logistic bandits with finite arms has not been established, which is beyond the scope of this work. Consequently, it remains open for the lower bound of MNL mixture MDPs with finite action space.
Remark 5.
While Corollary 1 is not a strict lower bound for MNL mixture MDPs with finite action space, it gives a basis to conjecture our results may be optimal in $d$ and $K$.
Dependence on $H$.
MNL mixture MDPs can be viewed as a generalization of linear mixture MDPs (Ayoub et al., 2020; Zhou et al., 2021). The pioneering work by Ayoub et al. (2020) achieved an $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ regret for linear mixture MDPs, which matches our results in Theorem 3, differing only on the lower-order term. Zhou et al. (2021) enhanced the dependence on $H$ and attained an optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$. This was made possible by recognizing that the value function in linear mixture MDPs is also linear, allowing for direct learning of value functions while incorporating their variance information. In contrast, the value function does not conform to a logistic model for MNL mixture MDPs, posing a significant challenge in using variance information of value functions. It remains open whether similar improvements in $H$ are attainable for MNL mixture MDPs.
Dependence on $\kappa$.
Though we fully remove the $\kappa$ dependence in Theorem 3 in the higher-order term and move it to the lower order term, Corollary 1 suggests that there may still be some room for improvement in the dependency on $\kappa$. The $\widetilde{\mathcal{O}}(d\sqrt{K\kappa^{*}})$ dependence has been established in the binary logistic bandit problem by Abeille et al. (2021). However, their techniques are limited to the binary case and are not directly applicable to MNL mixture MDPs which have multinomial rewards. For the multinomial feedback settings, Périvier and Goyal (2022) studied the MNL contextual bandit problem and achieved a regret bound of $\widetilde{\mathcal{O}}(d\sqrt{K\kappa^{*}}+\kappa^{-1})$. However, their results only hold for the scenarios with uniform revenue, where each choice yields identical rewards. In contrast, MNL mixture MDPs are more challenging as they involve diverse revenue values, represented by the value functions of different states. Thus, their techniques also can not be applied to MNL mixture MDPs. It remains open whether the $\widetilde{\mathcal{O}}(dH\sqrt{K\kappa^{*}})$ result can be achieved for MNL mixture MDPs.
7 Conclusion and Future Work
In this work, we address both the computational and statistical challenges for MDPs with multinomial logit function approximation. We first propose an algorithm that attains the previously best-known $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. Next, we identify that the statistical efficiency can be further improved by exploiting the local information. We propose two algorithms that not only maintain $\mathcal{O}(1)$ computational and storage costs but achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$, respectively. We also establish the first lower bound, justifying the optimality of our results in $d$ and $K$. To the best of our knowledge, this is the first work that achieves almost the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation.
There are several interesting directions for future work. First, there still exists a gap between the upper and lower bounds and how to close this gap is an open problem. Besides, while this work focuses on stochastic rewards, extending this model to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Fei et al., 2020; Li et al., 2023) is another interesting direction.
Appendix A Properties of the Multinomial Logit Function
This section collects several key properties of the multinomial logit function.
Without loss of generality, we assume $\forall\mathcal{S}_{h,s,a},\exists s^{\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi(s^{\prime}\>|\>s,a)=\mathbf{0}$. Otherwise, we can always define a new feature mapping $\phi^{\prime}(s^{\prime\prime}\>|\>s,a)=\phi(s^{\prime}\>|\>s,a)-\phi(s^{\prime\prime}\>|\>s,a)$ for any $s^{\prime\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi^{\prime}(s^{\prime}\>|\>s,a)=\mathbf{0}$ and the transition induced by $\phi^{\prime}$ is the same as that induced by $\phi$. We denote this state as $\dot{s}_{h,s,a}$ and $\dot{\mathcal{S}}_{h,s,a}=\mathcal{S}_{h,s,a}\backslash\{\dot{s}_{h,s,a}\}$.
Recall the following definitions in the paper:
$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$
$H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$
Lemma 5.
The following statements hold for any $k\in[K],h\in[H]$:
$H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\succeq\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}.$
Proof.
Note that
$\forall x,y\in\mathbb{R}^{d},(x-y)(x-y)^{\top}=xx^{\top}+yy^{\top}-xy^{\top}-yx^{\top}\succeq 0\Longrightarrow xx^{\top}+yy^{\top}\succeq xy^{\top}+yx^{\top}.$
Then, we have
$H_{k,h}(\theta)$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\frac{1}{2}\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\left(\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}+\phi_{k,h}^{s^{\prime\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\right)$
$\succeq$ $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\frac{1}{2}\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\left(\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}+\phi_{k,h}^{s^{\prime\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}\right)$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\Big{(}1-\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\theta)\Big{)}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
[…]
where the last inequality holds by the definition of $\kappa$ in Assumption 1. This finishes the proof. ∎
Lemma 6.
Suppose $\lambda\geq 1$, define $\widetilde{\phi}_{k,h}^{s^{\prime}}=\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime\prime}}$, for any $k\in[K],h\in[H]$, the following statements hold:
$(1)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}^{2}\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)$
$(2)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
$(3)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
$(4)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
$(5)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
Proof.
We prove the three statements individually.
Proof of statement (1).
By the definition that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$, we have
$\det(\widehat{\Sigma}_{k+1,h})=\det(\widehat{\Sigma}_{k,h})\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{k,h}^{-1}}\Big{)}.$
Then, we get
$\sum_{i=1}^{k}\log\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{i,h}^{-1}}\Big{)}\leq\log\left(\frac{\det(\widehat{\Sigma}_{k+1,h})}{\det(\widehat{\Sigma}_{1,h})}\right)\leq d\log\left(1+\frac{kU}{\lambda d}\right),$ (13)
where the last inequality holds by determinant trace inequality in Lemma 16. Since $\lambda\geq 1$, we have $\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\leq 1$, thus, we have
$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\min\bigg{\{}1,\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\bigg{\}}$
$\leq\frac{4}{\kappa}\sum_{i=1}^{k}\log\Bigg{(}1+\min\bigg{\{}1,\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\bigg{\}}\Bigg{)}$
$\leq\frac{4}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{i,h}^{-1}}\Big{)}$
$\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right),$
where the first inequality holds by the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$ and the last inequality holds by Equation (13).
Proof of statement (2).
By Lemma 5, we have $H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. Thus, we have
$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})p_{k,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
Then, we get
$\det(\mathcal{H}_{i+1,h})\geq\det(\mathcal{H}_{i,h})\Big{(}1+\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}\Big{)}.$
Since $\lambda\geq 1$, we have $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}\leq 1$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$, we get
$\sum_{i=1}^{k}\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}$
[…]
Proof of statement (3).
By Lemma 5, we have $\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k+1,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. Since $\lambda\geq 1$, we have $\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\leq\kappa$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$. By a similar analysis as the statement (2), we have
$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\Big{)}$
$\leq\frac{2}{\kappa}\log\left(\frac{\det(\mathcal{H}_{k+1,h})}{\det(\mathcal{H}_{1,h})}\right)$
$\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right).$
Proof of statement (4).
By the definition of $H_{k,h}(\theta)$, we have
$H_{i,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)p_{i,h}^{s^{\prime\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime\prime}})^{\top}$
$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}]-\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{(}\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{)}^{\top}$
$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}\big{[}(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})^{\top}\big{]}$
Thus, we have
$H_{i,h}(\widetilde{\theta}_{i+1,h})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\widetilde{\theta}_{i+1,h})(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}.$ (14)
Then, we get
$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}(\widetilde{\theta}_{k+1,h})(\widetilde{\phi}_{k,h}^{s^{\prime}})(\widetilde{\phi}_{k,h}^{s^{\prime}})^{\top}$
The remaining proof is similar to the proof of statement (2).
Proof of statement (5).
By (14), we have
$H_{i,h}(\widetilde{\theta}_{i+1,h})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\widetilde{\theta}_{i+1,h})(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}.$
Then, the remaining proof is similar to the proof of statement (3). ∎
Appendix B Proof of Lemma 1
B.1 Main Proof
Proof.
We define
$\bar{f}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[f_{k,h}(\theta)\mid\mathcal{F}_{k,h}],\quad\bar{g}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[g_{k,h}(\theta)\mid\mathcal{F}_{k,h}].$
By Property 1, we have $f_{k,h}(\theta)$ is exponential concave such that
$f_{k,h}(\widehat{\theta}_{k,h})\leq f_{k,h}(\theta_{h}^{*})+\langle g_{k,h}(\theta_{h}^{*}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert(\widehat{\theta}_{k,h}-\theta_{h}^{*})\rVert_{W_{k,h}}.$
Taking expectations on both sides, we have
$\bar{f}_{k,h}(\widehat{\theta}_{k,h})\leq\bar{f}_{k,h}(\theta_{h}^{*})+\langle\bar{g}_{k,h}(\theta_{h}^{*}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert(\widehat{\theta}_{k,h}-\theta_{h}^{*})\rVert_{W_{k,h}}.$ (15)
Based the property of KL diverge, we ensure the true parameter $\theta_{h}^{*}$ is the minimizer of the expected loss function by the following lemma.
Lemma 7.
For any $\theta\in\Theta$, we have $\bar{f}_{k,h}(\theta_{h}^{*})\leq\bar{f}_{k,h}(\theta)$.
Combining Lemma 7 and Equation (15), we have
00 0 $\leq\bar{f}_{k,h}(\widehat{\theta}_{k,h})-\bar{f}_{k,h}(\theta_{h}^{*})$
$\leq\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}$
$\leq\langle g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle.$ (16)
By standard analysis of OMD in Lemma 17, it holds
$2g_{k,h}(\widehat{\theta}_{k,h})(\widehat{\theta}_{k,h}-\theta_{h}^{*})\leq\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}}+\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}-\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}.$ (17)
Combining Equation (16) and Equation (17), we have
00 0 $\leq\frac{1}{2}\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}+\frac{1}{2}\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}-\frac{1}{2}\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$
$\hskip 85.35826pt-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle.$ (18)
First, we consider the first term . We show the gradient can be bounded by the following lemma.
Lemma 8.
For any positive semi-definite matrix $Z$, it holds that
$\lVert g_{k,h}(\theta)\rVert_{Z}^{2}\leq 4\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{Z}^{2}.$ (19)
Then, we consider the second term. Note that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}W_{k,h}$, we have
$\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}=\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{\widehat{\Sigma}_{k,h}}^{2}+\frac{\kappa}{2}\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{W_{k,h}}^{2}.$ (20)
Next, we bound $\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle$, which is a martingale difference sequence.
Lemma 9.
For any $\delta\in(0,1)$ and $\theta_{1},\ldots,\theta_{k},\theta^{*}\in[0,B]^{d}$, with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
$\sum_{i=1}^{k}\langle\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i}),\theta_{i}-\theta_{h}^{*}\rangle\leq\frac{\kappa}{4}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}+(\frac{4}{\kappa}+8B)\gamma_{k}.$ (21)
where $\gamma_{k}=\log\frac{2k^{2}H\log(kU)}{\delta}$.
Combining (18), (19), and (20), we have
$\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ $\leq\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}-\kappa\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}$
$\hskip 28.45274pt+2\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle$
$\leq\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k,h}}^{2}+4\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}+(\frac{8}{\kappa}+16B)\gamma_{k}$ (22)
where the first inequality holds by rearranging the terms in (18), the second holds by (19), (20) and (21). Summing (22) from $i=1$ to $k$, we have
$\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ $\leq\lambda B^{2}+4\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i+1,h}^{-1}}^{2}+(\frac{8}{\kappa}+16B)\gamma_{k}$
$\leq\lambda B^{2}+\frac{8}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)+(\frac{8}{\kappa}+16B)\gamma_{k},$
where the second inequality holds by $\lVert\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq 1$ and $\widehat{\Sigma}_{i,h}\geq I,\forall i\in[K]$, the last inequality is by the elliptical potential in Lemma 6. This finishes the proof. ∎
B.2 Proof of Auxiliary Lemmas
In this section, we provide the proofs of the lemmas used in Appendix B.1.
B.2.1 Proof of Lemma 7
Proof.
By the definition of $\bar{f}_{k,h}(\theta)$, we have
$\bar{f}_{k,h}(\theta)-\bar{f}_{k,h}(\theta_{h}^{*})$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log p_{k,h}^{s^{\prime}}(\theta)+\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\geq 0,$
where the last inequality is due to $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log\frac{p_{k,h}^{s^{\prime}}(\theta_{h}^{*})}{p_{k,h}^{s^{\prime}}(\theta)}$ is the Kullback-Leibler divergence between $p_{k,h}(\theta_{h}^{*})$ and $p_{k,h}(\theta)$, which always is non-negative. ∎
B.2.2 Proof of Lemma 8
Proof.
For any positive semi-definite matrix $Z$,
$(x_{i}-x_{j})^{\top}Z(x_{i}-x_{j})=x_{i}^{\top}Zx_{i}-x_{i}^{\top}Zx_{j}-x_{j}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq 0,\quad\forall x_{i},x_{j}\in\mathbb{R}^{d}.$ (23)
which implies $x_{i}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq x_{i}^{\top}Zx_{j}+x_{j}^{\top}Zx_{i},\forall x_{i},x_{j}\in\mathbb{R}^{d}$. Let $x_{i}=(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
$\lVert g_{k,h}(\theta)\rVert_{Z}^{2}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})(p_{k,h}^{s^{\prime\prime}}(\theta)-y_{k,h}^{s^{\prime\prime}})\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime\prime}}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})^{2}\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime}}$
$\hskip 56.9055pt+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})(p_{k,h}^{s^{\prime\prime}}(\theta)-y_{k,h}^{s^{\prime\prime}})(\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime\prime}}+\phi_{k,h}^{s^{\prime\prime}}Z\phi_{k,h}^{s^{\prime}})$
[…]
=\ = $2\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})^{2}\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime}}$
[…]
This finishes the proof. ∎
B.2.3 Proof of Lemma 9
Proof.
Notice that $\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$ is a martingale difference sequence. Also, we have
$\left|\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$ $\leq\left|\left(\bar{g}_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|+\left|\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
$\leq\lVert\bar{g}_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}+\lVert g_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}$
$\leq 4\sqrt{2}B,$
where the last inequality holds by $\lVert\bar{g}_{i,h}(\theta_{i})\rVert_{2}=\lVert\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta)-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq\sqrt{2}$.
We define the martingale $M_{k,h}=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$, and define the conditional variance $\sigma_{i}^{2}$ as
$\sigma_{k,h}^{2}$ $=\mathbb{E}[M_{k,h}^{2}\mid\mathcal{F}_{i-1}]$
$=\sum_{i=1}^{k}\mathbb{E}_{y_{k,h}}\left[\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})^{2}\right]$
$\leq\sum_{i=1}^{k}\mathbb{E}_{y_{k,h}}\left[\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})^{2}\right]$
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})^{2}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\triangleq A_{k,h}.$
where the first inequality is due to the fact that $\mathbb{E}[(\xi-\mathbb{E}[\xi])^{2}]\leq\mathbb{E}[\xi^{2}]$ for any random variable $\xi$. Note that $A_{k,h}$ is a random variable, so we cannot directly apply the Bernstein inequality to $M_{k,h}$. Instead, we consider the following two cases:(i) $A_{k,h}\leq\frac{4B^{2}}{kU}$ and (ii) $A_{k,h}>\frac{4B^{2}}{kU}$.
Case (i): $A_{k,h}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\leq\frac{4B^{2}}{kU}$. Then, we have
$M_{k,h}$ $=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$
$=\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta_{i})-y_{i,h}^{s^{\prime}})(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})$
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lvert(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})\rvert\leq\sqrt{kU\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}((\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*}))^{2}}\leq 2B.$
Case (ii): $A_{k,h}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{j,h}}^{2}>\frac{4B^{2}}{kU}$. We have both a lower and upper bound for $A_{k,h}$, i.e., $\frac{4B^{2}}{kU}<A_{k,h}\leq 4B^{2}kU$. Then, we can use the peeling process to bound $M_{k,h}$ as follows:
$\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3}\right]$
=\ = $\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3},\frac{4B^{2}}{kU}<A_{k,h}\leq 4kUB^{2}\right]$
=\ = $\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3},\frac{4B^{2}}{kU}<A_{k,h}\leq 4kUB^{2},\sigma_{k,h}\leq A_{k,h}\right]$
[…]
where $m=2\log_{2}(kU)$, and the last inequality follows the Bernstein inequality for martingales. Combining above two cases, letting $\tau=\log{\frac{mk^{2}}{\delta/H}}$ and taking the union bound over $k$ and $h\in[H]$, we have with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
$M_{k,h}$ $=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$
$\leq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3}+4\sqrt{2}B$
$\leq 2\sqrt{\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\log\frac{2k^{2}H\log(kU)}{\delta}}+8B\left(1+\log\frac{2k^{2}H\log(kU)}{\delta}\right)$
$=2\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}+8B\left(1+\gamma_{k}\right),$
where $\gamma_{k}=\log\frac{2k^{2}H\log(kU)}{\delta}$.
Then, applying $uv\leq cu^{2}+v^{2}/(4c)$ for any $c,u,v>0$ with $c=2/\kappa$, we have
$\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}\leq\frac{2\gamma_{k}}{\kappa}+\frac{\kappa}{8}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}$
Then we have
$\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\leq\frac{\kappa}{4}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}+(\frac{4}{\kappa}+8B)\gamma_{k}.$
This finishes the proof. ∎
Appendix C Proof of Lemma 2
Proof.
The gradient of $p_{s,a}(\theta)$ is given by $\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}$. By the mean value theorem, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}$
=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})\bigg{|}$
=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})\Big{(}V(s^{\prime})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})V(s^{\prime\prime})\Big{)}\bigg{|}$
[…]
where the first inequality is by $V:\mathcal{S}\to[0,H]$, the second inequality is by the definition of $\phi_{k,h}^{s^{\prime}}$, the third inequality is by the Holder’s inequality, the last inequality is by Lemma 1. ∎
Appendix D Proof of Theorem 1
D.1 Main Proof
To prove the theorem, we first introduce the following lemma.
Lemma 10.
Suppose for all $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\Gamma_{h,s,a}.$ (25)
Define
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{h,s,a}\bigg{]}_{[0,H]}.$ (26)
Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
$\operatorname*{Reg}(K)\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+H\sqrt{2KH\log(2/\delta)}.$
Proof of Theorem 1.
Substituting $\Gamma_{h,s,a}=H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}$ into Lemma 10, we have
$\operatorname*{Reg}(K)$ $\leq 2H\widehat{\beta}_{k}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}+H\sqrt{2KH\log(2/\delta)}$
$\leq 2H\widehat{\beta}_{k}\sum_{h=1}^{H}\sqrt{K\sum_{k=1}^{K}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}^{2}}+H\sqrt{2KH\log(2/\delta)}$
$\leq 2H^{2}\widehat{\beta}_{k}\sqrt{\frac{4dK}{\kappa}\log\left(1+\frac{KU}{d\lambda}\right)}+H\sqrt{2KH\log(2/\delta)}\leq\widetilde{\mathcal{O}}({\kappa}^{-1}dH^{2}\sqrt{K}),$
where the second inequality holds by the Cauchy-Schwarz inequality and the third inequality holds by Lemma 6. This finishes the proof. ∎
D.2 Proof of Auxiliary Lemmas
In this section, we provide the proofs of the lemmas used in Appendix D.1.
First, we introduce the following lemma.
Lemma 11.
Suppose (25) and (26) in Lemma 10 holds for all $k\in[K],h\in[H]$. Then, for any $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$, it holds that
$Q_{h}^{*}(s,a)\leq\widehat{Q}_{k,h}(s,a)\leq r_{h}(s,a)+\mathbb{P}_{h}V_{k,h+1}(s,a)+2\Gamma_{h,s,a}.$
Proof.
First, we prove the left-hand side of the lemma. We prove this by backward induction on $h$. For the stage $h=H$, by definition, we have
$\widehat{Q}_{k,H}(s,a)=r_{H}(s,a)=Q_{H}^{*}(s,a),\quad\widehat{V}_{k,H+1}(s)=0=V_{H+1}^{*}(s).$
Suppose the statement holds for $h+1$, we show it holds for $h$. By definition, if $\widehat{Q}_{k,h}(s,a)=H$, this holds trivially. Otherwise, we have
$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}+\Gamma_{h,s,a}$
$\geq r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})V_{k}^{*}+\Gamma_{h,s,a}\geq r_{h}(s,a)+p_{s,a}(\theta_{h}^{*})V_{k}^{*}=Q_{h}^{*}(s,a).$
where the first inequality is by the induction hypothesis, and the second inequality is by (25).
Then, we prove the right-hand side of the lemma. By the definition of $\widehat{Q}_{k,h}(s,a)$, we have
$\widehat{Q}_{k,h}(s,a)=r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}+\Gamma_{h,s,a}\leq r_{h}(s,a)+p_{s,a}(\theta_{h}^{*})\widehat{V}_{k,h+1}+2\Gamma_{h,s,a},$
where the inequality is by (25). ∎
D.2.1 Proof of Lemma 10
Proof.
By the definition of $\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$, we have
$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})=$ $\sum_{k=1}^{K}Q_{1}^{*}(s_{k,1},\pi^{*}(s_{k,1}))-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$
[…]
where the first inequality is by Lemma 11 and the second is by $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
By the right-hand side of Lemma 11, we have
$\widehat{Q}_{1}(s_{k,1},a_{k,1})-V_{1}^{\pi_{k}}(s_{k,1})$
=\ = $r(s_{k,1},a_{k,1})+\mathbb{P}_{1}\widehat{V}_{k,2}(s_{k,1},a_{k,1})++2\Gamma_{h,s_{k,1},a_{k,1}}-r(s_{k,1},a_{k,1})-\mathbb{P}_{1}V_{2}^{\pi_{k}}(s_{k,1},a_{k,1})$
[…]
Define $\mathcal{M}_{k,h}=\mathbb{P}_{h}(\widehat{V}_{k,h+1}-V_{h+1}^{\pi_{k}})(s_{k,h},a_{k,h})-(\widehat{V}_{k,h+1}-V_{h+1}^{\pi_{k}})(s_{k,h+1})$. Applying this recursively, we have
$\widehat{Q}_{1}(s_{k,1},a_{k,1})-V_{1}^{\pi_{k}}(s_{k,1})\leq 2\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+\sum_{h=1}^{H}\mathcal{M}_{k,h}$
Summing over $k$, we have for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
$\operatorname*{Reg}(K)\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+H\sqrt{2KH\log(2/\delta)}$
where the inequality holds by the Azuma-Hoeffding inequality as $\mathcal{M}_{k,h}$ is a martingale difference sequence with $\mathcal{M}_{k,h}\leq 2H$. This finishes the proof. ∎
Appendix E Proof of Lemma 3
Proof.
The proof is similar to the proof of Theorem 3 in Zhang and Sugiyama [2023] and Lemma 1 of Lee and Oh [2024]. Define
$\widetilde{f}_{k,h}(\theta)=f_{k,h}(\theta_{k,h})+\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{H_{k,h}(\widetilde{\theta}_{k,h})},$
which is a second order approximation of the original function $f_{k,h}(\theta)$ at $\widetilde{\theta}_{k,h}$. Then, the update rule in (8) can be rewritten as
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{f}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}.$ (27)
Then, we present the following lemma.
Lemma 12 (Lemma E.1 of Lee and Oh [2024]).
For the update rule defined in (27) as
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{f}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}},$
let $\eta=\frac{1}{2}\log(U+1)+(B+1)$, we have
$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}\leq$ $2\eta\left(\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})\right)+4\lambda B$
$+12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}-\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}.$
Then, we bound the right-hand side of the above lemma separately. First, we bound the first term. Define the softmax function as follows.
$[\sigma_{k,h}(z)]_{s}=\frac{\exp([z])_{s}}{1+\sum_{s\in\dot{\mathcal{S}}_{k,h}}\exp([z])_{s}},\quad\forall s\in\mathcal{S}_{k,h}.$
Then, the loss function in (2) can be rewritten as
$f_{k,h}(z_{k,h},y_{k,h})=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\operatorname{\mathds{1}}[y_{k,h}^{s^{\prime}}=1]\log\left(\frac{1}{[\sigma_{k,h}(z_{k,h})]_{s^{\prime}}}\right).$
Define a pseudo-inverse function of $\sigma_{k,h}(\cdot)$ as
$[\sigma_{k,h}^{-1}(p)]_{s^{\prime}}=\log\left(\frac{[p]_{s^{\prime}}}{1-\lVert p\rVert_{1}}\right),\forall p\in\{p\in[0,1]^{N_{k,h}}\mid\lVert p\rVert_{1}<1\}.$
Then, we decompose the first term as follows.
$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})=\underbrace{\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})}_{(a)}+\underbrace{\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})}_{(b)}$
where $z_{k,h}=\sigma^{-1}_{k,h}(\mathbb{E}_{\theta\sim P_{k,h}}[\sigma_{k,h}((\phi_{k,h}^{s^{\prime}})^{\top}\theta)_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}])$, $P_{k,h}\triangleq\mathcal{N}(\widetilde{\theta}_{k,h},(1+c\mathcal{H}_{k,h}^{-1}))$ is the Gaussian distribution with mean $\widetilde{\theta}_{k,h}$ and covariance $(1+c\mathcal{H}_{k,h}^{-1})$ where $c$ is a constant to be specified later.
First, we bound the term $(a)$ as follows.
Lemma 13 (Lemma E.2 of Lee and Oh [2024]).
Let $\delta\in(0,1]$ and $\lambda\geq 1$, for all $k\in[K],h\in[H]$, with probability at least $1-\delta$, we have
$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})$
[…]
Then, we bound the term $(b)$ as follows.
Lemma 14 (Lemma E.3 of Lee and Oh [2024]).
For any $c\geq 0$, let $\lambda\geq\max\{2,72cd\}$, then for all $k\in[K],h\in[H]$, we have
$\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})\leq\frac{1}{2c}\sum_{i=1}^{k}\left\|\widetilde{\theta}_{i,h}-\theta_{i+1,h}\right\|_{\mathcal{H}_{i,h}}^{2}+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right).$
Now, we are ready to prove Lemma 3. Combining Lemma 12, Lemma 13, and Lemma 14, we have
$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}$
[…]
$+$ $2+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right)\Bigg{]}+4\lambda B+12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}+(\frac{\eta}{c}-1)\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}+\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}$
[…]
$+2+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right)\Bigg{]}+4\lambda B,$
where the second inequality holds by setting $c={7\eta}/{6}$ and $\lambda\geq\max\{84\sqrt{2}\eta B,84d\eta\}$, we have
$12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}+\Big{(}\frac{\eta}{c}-1\Big{)}\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}+\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}$
=\ = $12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}-\frac{1}{7}\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}+\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}$
[…]
Thus, by setting $\eta=\frac{1}{2}\log(U+1)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, we have $\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}\leq\mathcal{O}\big{(}\sqrt{d}\log U\log(KH/\delta)\big{)}\triangleq\widetilde{\beta}_{k}$. This finishes the proof. ∎
Appendix F Proof of Theorem 2
Proof.
By the same analysis as the proof of Lemma 2 in Appendix C, we have
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq$ $H\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\phi_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h}-\theta_{h}^{*})$
[…]
where the first inequality holds by (24), the second inequality is by Holder’s inequality, and the third inequality is by Lemma 3.
Then, substituting $\Gamma_{h,s,a}=H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ into Lemma 10, we have
$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})\leq$ $2H\widetilde{\beta}_{k}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}+H\sqrt{2KH\log(2/\delta)}$
[…]
where the second inequality holds by the Cauchy-Schwarz inequality, and the third inequality is by Lemma 6. This finishes the proof. ∎
Appendix G Proof of Lemma 4
Proof.
By the second-order Taylor expansion, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widetilde{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})^{\top}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})$
The gradient of $p_{s,a}(\theta)$ is given by $\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}$. For the first-order term, we have
$\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})^{\top}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})$
[…]
where in the first inequality, we denote $\mathcal{S}_{h,s,a}^{+}$ as the subset of $\mathcal{S}_{h,s,a}$ such that $\phi_{s,a}^{s^{\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})$ is non-negative, the second inequality holds by the Holder’s inequality, and the third inequality is by the confidence set in Lemma 3.
For the second-order term, we define $u_{s,a}^{s^{\prime}}(\theta)=\phi_{s,a}^{s^{\prime}}\theta$, and $p_{s,a}^{s^{\prime}}(u)=\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}}\exp(u_{s,a}^{s^{\prime\prime}})}$, further define
$F(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})},\widetilde{F}(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})V(s^{\prime})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})}.$
Then, we have
$\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})$
=\ = $\frac{1}{2}\big{(}u(\widehat{\theta}_{k,h})-u(\theta_{h}^{*})\big{)}^{\top}\nabla^{2}\widetilde{F}(u(\bar{\theta}))\big{(}u(\widehat{\theta}_{k,h})-u(\theta_{h}^{*})\big{)}$
=\ = $\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\big{(}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{)}^{\top}\frac{\partial^{2}\widetilde{F}(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{)}$
[…]
where the inequality holds by $V(s)\in[0,H],\forall s$. According to Lemma 18, we have (omit the subscript $\mathcal{S}_{h,s,a}$ for simplicity):
$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
[…]
$+\frac{3H}{2}\sum_{s^{\prime}}\big{(}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta})).$ (29)
To bound the first term, by applying the AM-GM inequality, we obtain
$H\sum_{s^{\prime}}\sum_{s^{\prime\prime}\neq s^{\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
[…]
$\hskip 56.9055pt+\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}$
[…]
Plugging (30) into (29), we have
$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
[…]
=\ = $\frac{5H}{2}\sum_{s^{\prime}}\big{(}(\phi_{s,a}^{s^{\prime}})^{\top}(\widehat{\theta}_{k,h}-\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta}))$
[…]
where the last inequality holds by Lemma 3.
Combining (28) and (31) finishes the proof. ∎
Appendix H Proof of Theorem 3
Proof.
Combining Lemma 4 and Lemma 10, we have
$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$ $\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}(\epsilon_{k,h}^{\mathtt{fst}}+\epsilon_{k,h}^{\mathtt{snd}})+H\sqrt{2KH\log(2/\delta)}$
where
$\epsilon_{k,h}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
Next, we bound $\epsilon_{k,h}^{\mathtt{fst}}$ and $\epsilon_{k,h}^{\mathtt{snd}}$ respectively.
Bounding $\epsilon_{k,h}^{\mathtt{fst}}$.
For simplicity, we denote
$\mathbb{E}_{\theta}[\phi_{k,h}^{s^{\prime}}]=\mathbb{E}_{s^{\prime}\sim p_{k,h}^{s}(\theta)}[\phi_{k,h}^{s^{\prime}}],\quad\bar{\phi}_{s,a}^{s^{\prime}}=\phi_{s,a}^{s^{\prime}}-\mathbb{E}_{\widetilde{\theta}_{k,h}}[\phi_{k,h}^{s^{\prime}}],\quad\widetilde{\phi}_{s,a}^{s^{\prime}}=\phi_{s,a}^{s^{\prime}}-\mathbb{E}_{\widetilde{\theta}_{k+1,h}}[\phi_{k,h}^{s^{\prime}}]$
Then, we have
$\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}}=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
=\ = $\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(c)}+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(d)}$
$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(e)}$
Next, we bound these terms separately.
For the first term $(c)$, we have
$\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
=\ = $\Big{\lVert}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\left(p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k+1,h})-p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\right)\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
=\ = $\Big{\lVert}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\left(\nabla p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\right)\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
=\ = $\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\Big{\lvert}\Big{(}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime}}-p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime\prime}}\Big{)}^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\Big{\rvert}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
$\hskip 28.45274pt+\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\big{\lvert}(\phi_{k,h}^{s^{\prime\prime\prime}})^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\big{\rvert}$
[…]
$\hskip 28.45274pt+\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}}$
[…]
where the fifth inequality is by Cauchy-Schwarz inequality and the fourth inequality is because by Lemma 17, since ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
$\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}}\leq\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\widetilde{\mathcal{H}}_{k,h}}\leq 2\eta\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{\widetilde{\mathcal{H}}_{k,h}^{-1}}\leq\frac{2\eta}{\sqrt{\lambda}}\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2},$
and since $g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
$\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2}\leq\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}+\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}\leq 2\max_{s^{\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime}}\big{\rVert}_{2}\leq 2.$
Therefore, we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\frac{4\eta}{\sqrt{\lambda}}\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}$
[…]
where the last inequality holds by Lemma 16.
For the term $(d)$, by similar analysis, we have
$(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
=\ = $\nabla p_{k,h}^{s^{\prime}}(\xi_{k,h})^{\top}(\widetilde{\theta}_{k,h}-\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
=\ = $\Big{(}p_{k,h}^{s^{\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\xi_{k,h})\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{)}^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
where the third inequality holds by the AM-GM inequality. Thus, we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
[…]
where the last inequality holds by Lemma 16.
Finally, we bound the term $(e)$ as follows.
$\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\sqrt{\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})}\sqrt{\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}}$
[…]
where the last inequality holds by Lemma 16.
Thus, combining (32), (33), and (34), we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{fst}}$ $=H\widetilde{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{16\eta}{\kappa\sqrt{\lambda}}d\log(1+\frac{K}{d\lambda})\right)$ (35)
Bounding $\epsilon_{k,h}^{\mathtt{snd}}$.
For the second-order term, by Lemma 16, we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (36)
Combining (35) and (36), we have
$\operatorname*{Reg}(K)$ $\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}(\epsilon_{k,h}^{\mathtt{fst}}+\epsilon_{k,h}^{\mathtt{snd}})+H\sqrt{2KH\log(2/\delta)}$
$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{16\eta}{\kappa\sqrt{\lambda}}d\log(1+\frac{K}{d\lambda})\right)+\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right)$
$\hskip 28.45274pt+H\sqrt{2KH\log\Big{(}\frac{2}{\delta}\Big{)}}$
$\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1}\big{)}$
This finishes the proof. ∎
Appendix I Proof of Theorem 4
Proof.
Our proof is similar to the proof of the lower bound for linear mixture MDPs in Zhao et al. [2023]. We prove this lemma by translating this problem to learning $H/2$ logistic bandit problems.
Specifically, we use each three layers to construct a block. Note that the third layer of block $i$ is also the first layer of block $i+1$ and hence there are total $H/2$ blocks. In each block, both the first and third layers of this block only have one state, and the second layer has two states. Here we take block $i$ as an example. The first two layers of this block are associated with transition probability $\mathbb{P}_{i,1}$ and $\mathbb{P}_{i,2}$. Denote by $s_{i,1}$ the only state in the first layer of this block. In the second layer of the block $i$, we assume there exist two states $s_{i,2}^{*}$ and $s_{i,2}$. Let $s_{i,3}$ be the only state in the third layer of this block. Further, for any $a\in\mathcal{A}$, let $\phi(s_{i,1},a,s_{i,2})=0$. The probability of transferring to state $s_{i,2}^{*}$ when executing action $a$ at state $s_{i,1}$ is $\rho_{a}$, in particular,
$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=\rho_{a},\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)=1-\rho_{a}.$
For the second layer, $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$. The reward satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second layer for all $a\in\mathcal{A}$.
Consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from Bernoulli distributed with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*})^{-1})$. Thus, learning in this block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$. It is also clear that the optimal policy at state $s_{i,1}$ is to select action $a_{i,1}^{*}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$. Thus, learning this MDP is equal to learning $H/2$ logistic bandit problems. This finishes the proof. ∎
Appendix J Supporting Lemmas
In this part, we provide several supporting lemmas used in the proofs of the main results.
Lemma 15 (Abbasi-Yadkori et al. [2011, Theorem 1]).
Let $\left\{\mathcal{F}_{t}\right\}_{t=0}^{\infty}$ be a filtration. Let $\left\{\eta_{t}\right\}_{t=1}^{\infty}$ be a real-valued stochastic process such that $\eta_{t}$ is $\mathcal{F}_{t}$-measurable and $\eta_{t}$ is conditionally zero-mean $R$-sub-Gaussian for some $R\geq 0$ i.e. $\forall\lambda\in\mathbb{R}$, $\mathbb{E}\left[e^{\lambda\eta_{t}}\mid F_{t-1}\right]\leq\exp\left({\lambda^{2}R^{2}}/{2}\right)$. Let $\left\{X_{t}\right\}_{t=1}^{\infty}$ be an $\mathbb{R}^{d}$-valued stochastic process such that $X_{t}$ is $\mathcal{F}_{t-1}$-measurable. Assume that $V\in\mathbb{R}^{d\times d}$ is a positive definite matrix. For any $t\geq 0$, define $V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top},S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}$. Then, for any $\delta>0$, with probability at least $1-\delta$, for all $t\geq 0$,
$\left\|S_{t}\right\|_{V_{t}^{-1}}^{2}\leq 2R^{2}\log\left(\frac{\det\left(V_{t}\right)^{1/2}\det(V)^{-1/2}}{\delta}\right).$
Lemma 16 (Abbasi-Yadkori et al. [2011, Lemma 10]).
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, we have $\det(V_{t})\leq(\lambda+{tL^{2}}/{d})^{d}$.
Lemma 17 (Orabona [2019, Lemma 6.9]).
Let $Z$ be a positive define matrix and $\mathcal{X}$ be a convex set, define $\widetilde{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\langle\mathbf{g},\mathbf{x}\rangle+\frac{1}{2\eta}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{Z}^{2}$. Then, for all $\mathbf{x}^{*}\in\mathcal{X}$, we have
$\lVert\widetilde{\mathbf{x}}-\widehat{\mathbf{x}}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}\quad\mbox{and}\quad\langle\mathbf{g},\widehat{\mathbf{x}}-\mathbf{x}^{*}\rangle\leq\frac{\eta}{2}\lVert\mathbf{g}\rVert_{Z^{-1}}^{2}+\frac{1}{2\eta}\lVert\widehat{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}-\frac{1}{2\eta}\lVert\widetilde{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}.$
Lemma 18 (Lee and Oh [2024, Lemma D.3]).
Define $Q:\mathbb{R}^{K}\rightarrow\mathbb{R}$, such that for any $\mathbf{u}=\left(u_{1},\ldots,u_{K}\right)\in\mathbb{R}^{K},Q(\mathbf{u})=$ $\sum_{i=1}^{K}\frac{\exp\left(u_{i}\right)}{v+\sum_{k=1}^{K}\exp\left(u_{k}\right)}$. Let $p_{i}(\mathbf{u})=\frac{\exp\left(u_{i}\right)}{v+\sum_{k=1}^{K}\exp\left(u_{k}\right)}$. Then, for all $i\in[K]$,
$\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|\leqslant\begin{cases}3p_{i}(\mathbf{u})&\text{ if }i=j,\\ 2p_{i}(\mathbf{u})p_{j}(\mathbf{u})&\text{ if }i\neq j.\end{cases}$

### v2

@@
 Abstract
-We study a new class of MDPs that employs multinomial logit function approximation to ensure valid probability distributions over the state space.
+We study a new class of MDPs that employs multinomial logit (MNL) function approximation to ensure valid probability distributions over the state space.
 Despite its benefits, introducing the non-linear function raises significant challenges in both computational and statistical efficiency.
-The best-known method of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity, $d$ is the feature space dimension, $H$ is the episode length, and $K$ is the number of episodes.
-While this result attains the same rate in $K$ as the linear cases, the method requires storing all historical data and suffers from an $\mathcal{O}(K)$ computation cost per episode.
-Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases.
-In this work, we first address the computational issue by proposing an online algorithm that achieves the same regret with only $\mathcal{O}(1)$ computation cost.
-Then, we design two algorithms that leverage local information to enhance statistical efficiency.
-They not only maintain an $\mathcal{O}(1)$ computation cost per episode but achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ respectively.
-Finally, we establish a lower bound, justifying the optimality of our results in $d$ and $K$.
-To the best of our knowledge, this is the first work that achieves almost the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation for reinforcement learning.
+The best-known result of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity, $d$ is the feature dimension, $H$ is the episode length, and $K$ is the number of episodes.
+While this result attains the same rate in $K$ as linear cases, the method requires storing all historical data and suffers from an $\mathcal{O}(K)$ computation cost per episode.
+Moreover, the quantity $\kappa$ can be exponentially small in the worst case, leading to a significant gap for the regret compared to linear function approximation.
+In this work, we first address the computational and storage issue by proposing an algorithm that achieves the same regret with only $\mathcal{O}(1)$ cost.
+Then, we design an enhanced algorithm that leverages local information to enhance statistical efficiency.
+It not only maintains an $\mathcal{O}(1)$ computation and storage cost per episode but also achieves an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, nearly closing the gap with linear function approximation.
+Finally, we establish the first lower bound for MNL function approximation, justifying the optimality of our results in $d$ and $K$.
 1 Introduction
-Reinforcement Learning (RL) with function approximation has achieved great success in many applications with large state and action space, such as games (Mnih et al., 2013), robotics (Kober et al., 2013), large language models (Ouyang et al., 2022), etc.
+Reinforcement Learning (RL) with function approximation has achieved remarkable success in various applications involving large state and action spaces, such as games (Silver et al., 2016), algorithm discovery (Fawzi et al., 2022) and large language models (Ouyang et al., 2022).
 Therefore, establishing the theoretical foundation for RL with function approximation is of great importance.
-Recently, there have been many efforts devoted to understanding the linear function approximation due to its significance and simplicity, yielding fruitful results (Yang and Wang, 2019; Jin et al., 2020; Ayoub et al., 2020).
-While these works make important steps toward understanding reinforcement learning with function approximation, there are still some challenges to be solved.
-In RL with linear function approximation, transitions are presumed to be linear in specified feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs.
+Recently, there have been many efforts devoted to understanding the linear function approximation, yielding numerous valuable results (Yang and Wang, 2019; Jin et al., 2020; Ayoub et al., 2020).
+While these studies make important steps toward understanding RL with function approximation, there are still some challenges to be solved.
+In linear function approximation, transitions are assumed to be linear in specified feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs.
 Here $\mathbb{P}(s^{\prime}|s,a)$ is the probability from state $s$ to state $s^{\prime}$ taking action $a$, $\phi(s^{\prime}|s,a)$ and $\phi(s,a)$ are feature mappings, $\theta^{*}$ and $\mu^{*}(s^{\prime})$ are unknown parameters.
-However, a transition function is a probability distribution over states, i.e., the magnitude falls within the range of $[0,1]$, and the aggregate probability equals $1$.
-Thus, for certain feature mappings, linear transitions may not yield a valid probability distribution with arbitrary $\theta$ and $\mu$.
-In general, an algorithm designed for linear MDPs could break down entirely if the underlying MDP is not linear as noted by Jin et al.
-(2020).
-While there are some works exploring generalized linear function approximation (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Du et al., 2021; Foster et al., 2021), they make assumptions over value functions rather than transitions, hence do not tackle this challenge.
-Table 1: Comparison between previous work and ours in terms of the regret bound, computation cost and storage cost.
-Here $\kappa$ and $\kappa^{*}$ are exponentially small problem-dependent quantities, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes.
+However, a transition function is a probability distribution over states, i.e., the magnitude falls within the range of $[0,1]$, and the aggregation equals $1$.
+For certain feature mappings, linear transitions may not yield a valid probability distribution with arbitrary $\theta$ and $\mu$ as show by (Hwang and Oh, 2023).
+While there are also some works exploring generalized linear (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Foster et al., 2021; Chen et al., 2023), they made assumptions over value functions rather than transitions, hence do not tackle this challenge.
+Table 1: Comparison between previous works and ours in terms of the regret, computation cost and storage cost.
+Here $\kappa$ and $\kappa^{*}$ are exponentially small problem-dependent quantities defined in Assumption 1, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes.
 The computational cost and storage cost per episode indicate the dependence on episode count $k$.
-Reference Regret Comp. Stor. MDP model
+Reference Regret Com. Sto. MDP model
 Hwang and Oh (2023) $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ $\mathcal{O}(k)$ $\mathcal{O}(k)$ homogeneous
 UCRL-MNL-OL (Theorem 1) $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
-UCRL-MNL-LL (Theorem 2) $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
-UCRL-MNL-LL+ (Theorem 3) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
+UCRL-MNL-LL (Theorem 2) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
 Lower Bound  (Corollary 1) $\Omega(dH\sqrt{K\kappa^{*}})$ / / infinite action space
-Towards addressing the limitation of linear function approximation, a new class of MDPs that utilizes multinomial logit (MNL) function approximation has been proposed by Hwang and Oh (2023) recently.
-Despite its benefits, introducing non-linear functions raises significant challenges in both computational and statistical efficiency.
-Specifically, the best-known method of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret.
-While matching the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ as linear function approximation in terms of $K$ (Zhou et al., 2021; He et al., 2023), their method requires storing all historical data and the computational cost per episode grows linearly with the episode count, which is inefficient.
+Towards addressing the limitation of linear function approximation, a new class of MDPs that utilizes multinomial logit function approximation has been proposed by Hwang and Oh (2023) recently.
+Despite its benefits, the introduction of non-linear functions raises significant challenges in both computational and statistical efficiency.
+Specifically, the best-known approach of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret.
+While matching the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ as linear function approximation in terms of $K$ (Zhou et al., 2021; He et al., 2023), their method requires storing all historical data and the computational cost per episode grows linearly with the episode count, which is expensive.
 Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases.
 Thus, a natural question arises:
-Can we achieve (almost) the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation for RL?
-In this paper, we answer this question affirmatively by making significant advancements in both computational and statistical efficiency for RL with MNL function approximation, almost matching the results of linear function approximation.
-Our contributions can be summarized as follows:
+Is MNL function approximation more difficult than linear function approximation for RL?
+In this paper, we answer this question by making significant advancements in both computational and statistical efficiency for MDPs with MNL function approximation, nearly closing the gap with linear function approximation.
+Table 1 presents a comparison between previous studies and our work.
+Specifically, our contributions are summarized as follows:
   • We first propose the UCRL-MNL-OL algorithm based on a variant of Online Newton Step, attaining an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode.
-It matches the best-known regret by Hwang and Oh (2023), yet achieves the same computational and storage efficiency as linear function approximation (Jin et al., 2020; Zhou et al., 2021).
-  • As the quantity $\kappa$ can be exponentially small, we propose two algorithms UCRL-MNL-LL and UCRL-MNL-LL+, which leverage local information to improve the statistical efficiency.
-They not only maintain $\mathcal{O}(1)$ computational and storage costs but also achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, respectively.
-The latter result moves the $\kappa$ dependence to the lower-order term which is independent of $K$ and the higher-order term matches the optimal regret $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023) in both $d$ and $K$, differing only by an $\mathcal{O}(H^{1/2})$ factor.
-  • We establish the first lower bound for MDPs with MNL function approximation by employing a reduction to the logistic bandits.
-We prove a lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for the infinite action space setting.
-Though this does not constitute a strict lower bound for the finite actions case studied in this work, it justifies that our results may be optimal in $d$ and $K$.
-Table 1 shows comparisons between previous works and ours.
-From a technical perspective, inspired by recent advances in multinomial logistic bandits (Zhang and Sugiyama, 2023), we estimate the parameter by running an online algorithm instead of performing maximum likelihood estimation over all historical data.
-Differently, we need to design value functions carefully that utilize local information to address the unique challenges of MDPs, which is highly non-trivial.
+This result matches the best-known regret by Hwang and Oh (2023), yet achieves the same computational and storage efficiency as linear function approximation (Jin et al., 2020; Zhou et al., 2021).
+  • As the quantity $\kappa$ can be exponentially small, we propose an enhanced algorithm UCRL-MNL-LL, which leverages the local information to improve the statistical efficiency.
+It not only maintain $\mathcal{O}(1)$ computational and storage costs but achieve an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$.
+The higher-order term matches the optimal regret $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023) in $d$ and $K$, differing only by an $\mathcal{O}(H^{1/2})$ factor.
+  • We establish the first lower bound for MDPs with MNL function approximation by introducing a reduction to the logistic bandit problem.
+We prove a lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for MDPs with infinite action space.
+Though this does not constitute a strict lower bound for the finite actions case studied in this work, it suggests that our result may be optimal with respect to $d$ and $K$.
+[^1] After the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al.
+(2024) improved the lower bound to $\Omega(dH^{3/2}\sqrt{K})$.
+This confirms that our result is indeed optimal with respect to $d$ and $K$.
+From a technical perspective, inspired by recent advances in logistic bandits (Zhang et al., 2016; Faury et al., 2020; Zhang and Sugiyama, 2023), we observe that the negative log-likelihood function is exponentially concave, which motivates us to apply online Newton step to estimate the unknown transition parameter in an online manner.
+Moreover, we employ the Bernstein-like inequalities and the self-concordant-like property (Bach, 2010) of the log-loss to achieve the better dependence on $\kappa$.
 Organization.
-We first introduce the related work in Section 2 and present the setup in Section 3.
-Then, we introduce an algorithm with constant computational and storage costs in Section 4.
-Next, we present two improved algorithms in Section 5.
+We introduce the related work in Section 2 and present the setup in Section 3.
+Then, we design an algorithm with constant computational and storage costs in Section 4.
+Next, we present an algorithm with improved regret guarantee in Section 5.
 Finally, we establish the lower bound in Section 6.
 Section 7 concludes the paper.
-Due to space constraints, we defer all proofs to the appendixes.
+Due to space limits, we defer all proofs to the appendixes.
 Notations.
-We denote by $[n]$ the set $\{1,\ldots,n\}$ and define $[x]_{[a,b]}=\min(\max(x,a),b)$.
+Denote by $[n]$ the set $\{1,\ldots,n\}$ and use $[x]_{[a,b]}$ to denote $\min(\max(x,a),b)$.
 For $X,Y\in\mathbb{R}^{d\times d}$, $X\succeq Y$ means $X-Y$ is positive semi-definite.
 For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, denote $\lVert\mathbf{x}\rVert_{A}=\sqrt{\mathbf{x}^{\top}A\mathbf{x}}$.
-The $\widetilde{\mathcal{O}}(\cdot)$ hides all polylogarithmic factors.
+The $\widetilde{\mathcal{O}}(\cdot)$ hides polylogarithmic factors.
 2 Related Work
-In this section, we review related works, both in terms of settings and techniques.
+In this section, we review related works from both setup and technical perspectives.
 RL with Generalized Linear Function Approximation.
@@
 Moreover, these works also enter the computational and statistical challenges arising from non-linear function approximation that remain to be addressed.
-The most relevant work to ours is the recent work by Hwang and Oh (2023), which firstly explored a similar setting to ours, where the transition is modeled using a multinomial logit model.
-We significantly improve their results by providing computationally and statistically efficient algorithms.
+The most relevant work to ours is the recent work by Hwang and Oh (2023), which firstly explored a similar setting to ours, where the transition is characterized using a multinomial logit model.
+We significantly improve upon their results by providing computationally and statistically more efficient algorithms.
 RL with General Function Approximation.
-There have also been some works studies reinforcement learning with general function approximation.
+There have also been some works studies RL with general function approximation.
 Russo and Roy (2013) and Osband and Roy (2014) initiated the study on the minimal structural assumptions that render sample-efficient learning by proposing a structural condition called Eluder dimension.
 Recently, several works have investigated different conditions for sample-efficient interactive learning, such as Bellman Eluder (BE) dimension (Jin et al., 2021), Bilinear classes (Du et al., 2021), Decision-Estimation Coefficient (DEC) (Foster et al., 2021), and Admissible Bellman Characterization (ABC) (Chen et al., 2023).
-A notable difference is that they impose assumptions on the value functions while we study function approximation on the transitions.
+A notable difference is that they impose assumptions on the value functions while we study function approximation on the transitions to ensure valid probability distributions.
 Moreover, the goal of these works is to study the conditions for sample-efficient reinforcement learning, but not focus on the computational efficiency.
-Multinomial Logit (MNL) Contextual Bandits.
-As our work shares similar techniques with MNL contextual bandits to some extent, we briefly review works on this topic.
-The pioneering study by Cheung and Simchi-Levi (2017) introduced an algorithm that achieves a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{K})$, where $K$ denotes the number of rounds.
+Multinomial Logit Bandits.
+There are two types of multinomial logit bandits studied in the literature: the single-parameter model, where the parameter is a vector (Cheung and Simchi-Levi, 2017) and multiple-parameter model, where the parameter is a matrix (Amani and Thrampoulidis, 2021).
+We focus on the single-parameter model, which are more relevant to our setting.
+The pioneering work by Cheung and Simchi-Levi (2017) achieved a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{K})$, where $K$ denotes the number of rounds in bandits.
 This result was further enhanced by subsequent studies (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023).
-The recent work by Périvier and Goyal (2022) has significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ in the uniform revenue setting.
-Most of the above methods require storing all historical data and are computationally inefficient.
-A concurrent work with ours by Lee and Oh (2024) proposed an algorithm with constant computational and storage costs inspired by recent advancements in multinomial logistic bandits (Zhang and Sugiyama, 2023) and achieved optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K})$ and $\widetilde{\mathcal{O}}(d\sqrt{K})$ under uniform and nonuniform rewards respectively.
-However, although the underlying models of MNL contextual bandits and MNL mixture MDPs are similar, the challenges they pose differ significantly and the techniques developed for MNL contextual bandits can not be directly applied to the MNL mixture MDPs.
-Specifically, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize expected revenue, whereas in MNL mixture MDPs, the goal is to maximize cumulative rewards Therefore, while we utilize similar parameter estimation techniques as those in MNL contextual bandits, we need to construct value functions to address the unique challenges of MDPs, which is highly non-trivial and needs a careful design.
+In particular, Périvier and Goyal (2022) significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ in the uniform revenue setting.
+Most prior methods required storing all historical data and faced computational challenge.
+To address this issue, the most recent work by Lee and Oh (2024) proposed an algorithm with constant computational and storage costs building on recent advances in multiple-parameter model (Zhang and Sugiyama, 2023).
+Their algorithm achieves the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ and $\widetilde{\mathcal{O}}(d\sqrt{K}+\kappa^{-1})$ under uniform and non-uniform rewards respectively.
+However, although the underlying models of MNL bandits and MDPs share similarities, the challenges they present differ substantially, and techniques developed for MNL bandits cannot be directly applied to MNL MDPs.
+For example, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize the expected revenue, whereas in MNL MDPs, the goal is to choose one action at each stage to maximize the cumulative reward.
+Thus, it is necessary to design new algorithms tailored for MDPs to address these unique challenges.
 3 Problem Setup
@@
 For any function $V:\mathcal{S}\to\mathbb{R}$, we define $[\mathbb{P}_{h}V](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot\mid s,a)}V(s^{\prime})$.
-The Bellman equation for policy $\pi$ and the Bellman optimality equation are given by
+The Bellman equation for policy $\pi$ and Bellman optimality equation are given by
 $Q_{h}^{\pi}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{\pi}\right](s,a),\quad Q_{h}^{*}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{*}\right](s,a).$
@@
 In the online MDP setting, the learner interacts with the environment without the knowledge of the transition $\{\mathbb{P}_{h}\}_{h=1}^{H}$.
-For the sake of simplicity, we assume the reward $\{r_{h}\}_{h=1}^{H}$ is known.
+As learning rewards is no harder than transitions, we assume the reward $\{r_{h}\}_{h=1}^{H}$ is known.
 The interaction proceeds in $K$ episodes.
-At the beginning of each episode $k$, the learner chooses a policy $\pi_{k}=\{\pi_{k,h}\}_{h=1}^{H}$.
+At the beginning of episode $k$, the learner chooses a policy $\pi_{k}=\{\pi_{k,h}\}_{h=1}^{H}$.
 At each stage $h\in[H]$, starting from the initial state $s_{k,1}$, the learner observes the state $s_{k,h}$, chooses an action $a_{k,h}$ sampled from $\pi_{k,h}(\cdot\>|\>s_{k,h})$, obtains reward $r_{h}(s_{k,h},a_{k,h})$ and transits to the next state $s_{k,h+1}\sim\mathbb{P}_{h}(\cdot\>|\>s_{k,h},a_{k,h})$ for $h\in[H]$.
 The episode ends when $s_{H+1}$ is reached.
-The goal of the learner is to minimize regret, defined as the difference between the cumulative reward of the optimal policy and the learner’s policy, namely,
-$\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1}).$
+The goal of the learner is to minimize regret, defined as
+$\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1}),$
+which is the difference between the cumulative reward of the optimal policy and the learner’s policy.
 Multinomial Logit (MNL) Mixture MDPs.
@@
 To overcome this limitation, they propose a new class of MDPs with multinomial logit function approximation.
-We introduce the formal definition of MNL mixture MDPs below.
+We introduce the definition of MNL mixture MDPs below.
 Definition 1 (Reachable States).
 For any $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, we define the “reachable states” as the set of states that can be reached from state $s$ taking action $a$ at stage $h$ within a single transition, i.e., $\mathcal{S}_{h,s,a}\triangleq\{s^{\prime}\in\mathcal{S}\mid\mathbb{P}_{h}(s^{\prime}\>|\>s,a)>0\}$.
-Also, we define $S_{h,s,a}\triangleq|\mathcal{S}_{h,s,a}|$ and further define $U\triangleq\max_{(h,s,a)}S_{h,s,a}$ the maximum number of reachable states.
+Furthermore, we define $S_{h,s,a}\triangleq|\mathcal{S}_{h,s,a}|$ and denote by $U\triangleq\max_{(h,s,a)}S_{h,s,a}$ the maximum number of reachable states.
+Remark 1.
+There are many cases that even when the state space is very large, the maximum number of reachable states $U$ is small.
+This phenomenon is common in situations where the next state is close to the current state.
+An illustrative example is the RiverSwim problem (Strehl and Littman, 2004).
 Definition 2 (MNL Mixture MDP).
@@
 $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta_{h}^{*})}{\sum_{\widetilde{s}\in\mathcal{S}_{h,s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta_{h}^{*})}.$
+Remark 2.
+While the work of Hwang and Oh (2023) focuses on the homogeneous setting, where the transitions remain the same across all stages (i.e., $\mathbb{P}_{1}=...=\mathbb{P}_{H}$), we address the more general inhomogeneous setting, allowing transitions to vary across different stages.
 For any $\theta\in\mathbb{R}^{d}$, we define the induced transition as $p_{s,a}^{s^{\prime}}(\theta)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta)}{\sum_{\widetilde{s}\in\mathcal{S}_{s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta)}$.
-Next, we introduce a key assumption for this problem as in Hwang and Oh (2023).
+We introduce the key problem-dependent quantities for this problem firstly introduced by Hwang and Oh (2023).
 Assumption 1.
 There exists $10<\kappa\leq\kappa^{*}<$ 1 such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime},s^{\prime\prime}\in\mathcal{S}_{h,s,a}$, it holds that $\inf_{\theta\in\Theta}p_{s,a}^{s^{\prime}}(\theta)p_{s,a}^{s^{\prime\prime}}(\theta)\geq\kappa$ and $p_{s,a}^{s^{\prime}}(\theta^{*})p_{s,a}^{s^{\prime\prime}}(\theta^{*})>\kappa^{*}$.
-Assumption 1 is equivalent to the standard assumption in generalized linear model literature (Filippi et al., 2010; Wang et al., 2021) and logistic bandits literature (Faury et al., 2020; Abeille et al., 2021; Zhang and Sugiyama, 2023) to guarantee the Hessian matrix is non-singular.
-4 Efficient Algorithm with Online Update
-In this part, we design an algorithm that achieves $\mathcal{O}(1)$ computational and storage costs.
-Efficiency Concern in Previous Methods.
-While Hwang and Oh (2023) made the first step and proposed the UCRL-MNL algorithm with $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, the computational and storage complexity of their algorithm scales linearly with the episode count, which is expensive.
-Specifically, for the trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$ at episode $k$, we define the variable: $y_{k,h}\in\{0,1\}^{N_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\operatorname{\mathds{1}}{(s_{k,h+1}=s^{\prime})}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $N_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$.
-Furthermore, we denote by $p_{k,h}^{s^{\prime}}(\theta)=p_{s_{k,h},a_{k,h}}^{s^{\prime}}(\theta)$.
+Assumption 1 is similar to the standard assumption in generalized linear model literature (Filippi et al., 2010; Wang et al., 2021) and logistic bandits literature (Faury et al., 2020; Abeille et al., 2021; Zhang and Sugiyama, 2023) to guarantee the Hessian matrix is non-singular in Property 1.
+4 Computationally Efficient Algorithm by Online Learning
+In this part, we design an algorithm that achieves $\mathcal{O}(1)$ computational and storage costs per episode.
+Since the transition parameter $\theta_{h}^{*}$ is unknown, we need to estimate it using the historical data.
+At episode $k$, we collect a trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$, then define the variable: $y_{k,h}\in\{0,1\}^{N_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\operatorname{\mathds{1}}{(s_{k,h+1}=s^{\prime})}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $N_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$.
+We denote by $p_{k,h}^{s^{\prime}}(\theta)=p_{s_{k,h},a_{k,h}}^{s^{\prime}}(\theta)$.
 Then $y_{k,h}$ is a sample from the following multinomial distribution:
-$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{N_{k,h}}}(\theta^{*})]).$
+$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{N_{k,h}}}(\theta^{*})]),$
 where the parameter $1$ indicates that $y_{k,h}$ is a single-trial sample.
-We define the noise $\epsilon_{k,h}^{s^{\prime}}=y_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\theta_{h}^{*})$.
+Furthermore, we define the noise $\epsilon_{k,h}^{s^{\prime}}=y_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\theta_{h}^{*})$.
 It is clear that $\epsilon_{k,h}\in[-1,1]^{N_{k,h}}$, $\mathbb{E}[\epsilon_{k,h}]=\mathbf{0}$ and $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\epsilon_{i,h}^{s^{\prime}}=0$.
-The UCRL-MNL algorithm estimates the unknown transition parameter $\theta_{h}^{*}$ using the regularized maximum likelihood estimation (MLE).
-Specifically, based on $y_{k,h}$, the estimator $\theta_{k,h}$ is defined as the solution to the following optimization problem:
+Efficiency Concern.  Hwang and Oh (2023) made the first step by proposing an approach using the maximum likelihood estimation (MLE). Specifically, their estimator $\theta_{k,h}$ is defined as
 $\theta_{k,h}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}-y_{i,h}^{s^{\prime}}\log p_{i,h}^{s^{\prime}}(\theta)+\frac{\lambda}{2}\lVert\theta\rVert_{2}^{2}.$ (1)
-In episode $k$, the estimator $\theta_{k,h}$ in (1) is computed using all samples collected in previous episodes, resulting in the computational and storage cost per episode that scales linearly with the episode count, which is expensive in practice.
-To address this issue, we introduce a novel algorithm named UCRL-MNL with Online Learning (UCRL-MNL-OL), which achieves both the same regret with constant computational and storage cost.
+Despite favorable theoretical guarantee, the computational and storage cost of this method is expensive.
+First, in episode $k$, the estimator $\theta_{k,h}$ in (1) is computed using all samples collected in previous episodes, resulting an $\mathcal{O}(k)$ storage cost.
+Second, to solve this optimization problem, the projected gradient descent (Boyd and Vandenberghe, 2004) are usually applied.
+However, as discussed in Faury et al.
+(2022), the optimization of the MLE problem typically require $\mathcal{O}(k\log(1/\epsilon))$ iterations to achieve an $\epsilon$-accurate solution, which is computationally expensive.
+To address this issue, we introduce a novel algorithm named UCRL-MNL with Online Learning (UCRL-MNL-OL), which achieves both the same regret with $\mathcal{O}(1)$ computational and storage cost.
 At a high level, our algorithm can be divided into two phases: (i) efficient online estimation, and (ii) efficient optimistic value function construction.
@@
 Efficient Online Estimation.
-Instead of using all historical data, our algorithm updates the estimator in an efficient online manner.
-This is achieved by efficiently exploiting particular properties of the MNL model.
-The key observation is that the negative log-likelihood function is exponentially concave over the bounded domain, which motivates us to apply a variant of online Newton step (Hazan et al., 2007), in particular, inspired by the works (Hazan et al., 2014; Zhang et al., 2016; Oh and Iyengar, 2021), which focus on logistic bandits and multinomial logit contextual bandits.
-First, we define pe-episode loss function $f_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
+Instead of using all historical data, our algorithm updates the estimator in an online manner.
+Specifically, inspired by the works (Hazan et al., 2014; Zhang et al., 2016) on logistic bandit, we find the negative log-likelihood function is exponentially concave, which motivates us to apply a variant of online Newton step (Zhang et al., 2016; Oh and Iyengar, 2021).
+First, we define per-episodic loss function $f_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
 $f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}},$
 $H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$ (2)
-We introduce the following key property of the loss function $f_{k,h}(\theta)$ by a second-order Taylor expansion, whose proof can be found in Lemma 5 in Appendix A.
+We introduce the following key property of the loss function $f_{k,h}(\theta)$ by a second-order Taylor expansion and mean value theorem, whose proof can be found in Lemma 5 in Appendix A.
 Property 1.
@@
 (ii). For any $\theta\in\Theta$, it holds that
-$H_{k,h}(\theta)\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\triangleq\kappa W_{k,h},\forall\theta\in\Theta.$ (3)
-This implies that $f_{k,h}(\theta)$ is exponentially concave over the bounded domain $\Theta$, which enables us to apply the online Newton step to construct an efficient estimator.
-Specifically, we construct the estimator $\widehat{\theta}_{k+1,h}$ by solving the problem:
+$H_{k,h}(\theta)\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\triangleq\kappa W_{k,h}.$ (3)
+Property 1 implies that $f_{k,h}(\theta)$ is exponentially concave, which enables us to apply the online Newton step to construct an efficient estimator.
+Specifically, we construct $\widehat{\theta}_{k+1,h}$ by solving the problem:
 $\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2},$ (4)
 where $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ is the feature covariance matrix.
-This estimator can be computed efficiently in an online manner without the need to store all historical data.
-Then, we construct the confidence set by the following lemma.
+We show the online estimator $\widehat{\theta}_{k,h}$ in (4) enjoys computational and storage efficiency simultaneously.
+As the optimization problem exhibits a standard online mirror descent formulation, it can be solved with a projected gradient step with the following equivalent formulation by
+$\theta^{\prime}_{k+1,h}=\widehat{\theta}_{k,h}-\widehat{\Sigma}_{k+1,h}^{-1}g_{k,h}(\widehat{\theta}_{k,h}),\quad\mbox{and}\quad\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widehat{\Sigma}_{k+1,h}}.$
+This update enjoys a computational cost of only $\mathcal{O}(d^{3}U)$, independent of episode count $k$ (Zhang and Sugiyama, 2023; Lee and Oh, 2024).
+As for storage costs, it avoids the need to store all historical data by updating the feature covariance matrix $\widehat{\Sigma}_{k,h}$ incrementally, requiring only $\mathcal{O}(d^{2})$ storage cost.
+We note that a dependence on $U$ is introduced in the computational cost.
+However, as we have discussed in Remark 1, $U$ can be much smaller than the state space size $S$, which is acceptable.
+This dependency is typical for model-based methods that directly learn transitions, as it need to control the estimation error of transitions, which typically involves a total of $U$ elements.
+Similar dependencies have been observed in the literature (Hwang and Oh, 2023).
+Additionally, even for the model-free method (Yang and Wang, 2020; Zhou et al., 2021) which learn value functions, a common assumption is the value $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}V(s^{\prime})$ can be obtained by an Oracle, which depends on $U$ implicitly.
+Then, we show the estimator $\widehat{\theta}_{k,h}$ is close to the true parameter $\theta_{h}^{*}$ by the following lemma.
 Lemma 1.
 For any $\delta\in(0,1)$, define the confidence set as
-$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k,h}}\leq\sqrt{\lambda}B+\sqrt{\frac{8}{\kappa}d\log\left(1+\frac{kUH}{d\lambda\delta}\right)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$
+$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k,h}}\leq\sqrt{\lambda}B+\sqrt{\frac{8}{\kappa}d\log\left(1+\frac{kUH}{d\lambda\delta}\right)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$ (5)
 Then, we have $\Pr[\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
@@
 We observe similar oversights also appear in the works on multinomial logit contextual bandits (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023).
-To the best of our knowledge, this issue has not been explicitly identified in previous studies.
-We note that this issue can be resolved with only slight modifications in constant factors by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al.
-(2024), a simplified version of Theorem C.6 in Périvier and Goyal (2022).
+To our knowledge, this issue has not been explicitly identified in previous studies.
+We note that this issue can be resolved by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al.
+(2024b) with only slight modifications in constant factors.
 Algorithm 1 UCRL-MNL-OL
@@
 6:         Set $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$.
-7:        Compute $\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ as (4).
+7:         Compute $\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$.
 8:     end for
@@
 Given the confidence set $\widehat{\mathcal{C}}_{k,h}$, the most direct way is to adopt the principle of “optimism in the face of uncertainty” and construct the optimistic value function as the maximum expected reward over the confidence set.
-Specifically, define the value function $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
+Specifically, we define the optimistic value function $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
 $\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\max_{\theta\in\widehat{\mathcal{C}}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})\bigg{]}_{[0,H]},\widehat{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s,a).$ (6)
@@
 Thus, we can use a closed-form confidence bound instead of computing the maximal value over the confidence set.
-To this end, we present the following lemma, which enables us to construct the optimistic value function efficiently.
+To this end, we present the following lemma that enables us to construct the optimistic value function more efficiently.
 Lemma 2.
-Suppose Lemma 1 holds. For any function $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, with probability at least $1-\delta$, it holds that
+Suppose Lemma 1 holds. For any $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds
 $\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}.$
@@
 $\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\widehat{\Sigma}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (7)
-At state $s_{k,h}$, we chooses $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
+At state $s_{k,h}$, our algorithm chooses $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
 Algorithm 1 presents the detailed procedure.
-We show that our algorithm enjoys the following regret guarantee.
+We show that our algorithm achieves the following regret guarantee.
 Theorem 1.
-Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\widehat{\beta}_{k}$ as in Lemma 1 and $\lambda=1$, with probability at least $1-\delta$, Algorithm 1 ensures the following guarantee
-${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K}).$
+For any $\delta\in(0,1)$, set $\widehat{\beta}_{k}$ as in Lemma 1 and $\lambda=1$, with probability at least $1-\delta$, Algorithm 1 ensures the following regret guarantee
+${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}\big{(}\kappa^{-1}dH^{2}\sqrt{K}\big{)}.$
 Remark 4.
 Our algorithm attains the same regret of Hwang and Oh (2023), yet for the more general inhomogeneous MDPs.
-More importantly, our algorithm only requires $\mathcal{O}(1)$ computational and storage costs, matching the computational efficiency of MDPs with linear function approximation (Ayoub et al., 2020; Jin et al., 2020; Zhou et al., 2021; He et al., 2023).
-5 Improved Algorithm by Local Learning
-While the UCRL-MNL-OL algorithm enjoys favorable computational efficiency, it suffers from a dependence on $\kappa^{-1}$ in the regret.
-By Assumption 1, the quantity $\kappa$ may be exponentially small, exhibiting an exponential dependence on the radius of the parameter set and linear dependence on the number of reachable states, i.e., $\kappa^{-1}=\mathcal{O}((U\exp(B))^{2})$.
-Thus, there is a significant gap between MNL function approximation and linear function approximation.
-To bridge this gap, we propose new algorithms that leverage local information to enhance the dependence on $\kappa$ significantly.
+Importantly, our algorithm only requires constant computational and storage costs, matching the computational efficiency of the linear cases (Jin et al., 2020; Zhou et al., 2021).
+5 Statistically Improved Algorithm by Local Learning
+In this section, we present an enhanced algorithm, UCRL-MNL-LL, that leverages local information to improve the statistical efficiency for MDPs with MNL function approximation.
+While the UCRL-MNL-OL algorithm in Section 4 offers favorable computational and storage efficiency, it suffers from a dependence on $\kappa^{-1}$ in the regret.
+By Assumption 1, the quantity $\kappa$ may be exponentially small in the worst case, exhibiting an exponential dependence on the radius of the parameter set and linear in the number of reachable states, i.e., $\kappa^{-1}=\mathcal{O}((U\exp(B))^{2})$.
+This creates a significant gap between MNL and linear function approximation.
+Recently, the improved dependence on $\kappa$ has been achieved in the logistic bandit literature (Faury et al., 2020; Abeille et al., 2021; Périvier and Goyal, 2022) by the use of generalization of the Bernstein-like tail inequality (Faury et al., 2020) and the self-concordant-like property of the log loss (Bach, 2010).
+Thus, a natural question then arises: Can we improve the statistical efficiency of MNL function approximation for MDPs?
+We answer this question affirmatively by proposing an enhanced algorithm that reduces the dependence on $\kappa$ significantly through the use of local information.
+Though the achievement has been made in the logistic bandit, the extension to MDPs with MNL function approximation is non-trivial and new techniques specific to MDPs are required.
 We first analyze where the dependence on $\kappa$ comes from. By the analysis of Theorem 1, we can see that the regret of the algorithm can be upper-bounded as follows,
 $\operatorname*{Reg}(K)\leq 2H\widehat{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}.$ (8)
-where the first term is the overestimate of the value function and the second term is the martingale difference sequence due to the stochastic transition.
-The second term can be upper-bounded by the Azuma-Hoeffding inequality which is independent of $\kappa$.
-Thus, the $\kappa$-dependence comes from the first term.
-By the construction of the confidence set in Lemma 1, the dependence on $\widehat{\beta}_{k}$ in the first term leads to a $\kappa^{-1/2}$ dependence.
-By the configuration of the covariance matrix $\widehat{\Sigma}_{k,h}$ in (4), the dependence on $\widehat{\Sigma}_{k,h}$ also leads to a $\kappa^{-1/2}$ dependence.
-Thus, we need to design a local learning algorithm that can adaptively use local information to enhance the dependence on these two terms.
+Here, the first term corresponds to the overestimation of the value function, and the second term represents the martingale difference sequence arising from the stochastic transition dynamics.
+The second term can be bounded using the Azuma-Hoeffding inequality, which is independent of $\kappa$.
+For the first term, by the construction of the confidence set in Lemma 1, $\widehat{\beta}_{k}$ represents the width of the confidence set, contributing to a $\kappa^{-1/2}$ dependence.
+The feature covariance matrix $\widehat{\Sigma}_{k,h}$ defined in (4) reflects the degree of exploration across different states, which also results in a $\kappa^{-1/2}$ dependence.
+To mitigate these dependencies, we design a local learning algorithm that: (i) constructs a confidence set independent of $\kappa$; and (ii) builds a $\kappa$-independent feature covariance matrix that effectively captures the exploration degree across different states.
+We introduce the details in the following.
 5.1 Improved Online Estimation
-In Section 4, we show the Hessian matrix $H_{k,h}$ is lower bounded by a positive definite matrix $\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ uniformly over the parameter domain in (3).
-This lower bound is not tight as the Hessian matrix can be much larger at certain regions rather than the global minimum.
-This motivates us to design a local learning algorithm that can adaptively use local information.
-Inspired by recent advances in multinomial logistic bandit (Zhang and Sugiyama, 2023) and multinomial logit contextual bandit (Lee and Oh, 2024), we run an online algorithm to estimate the parameter $\theta_{h}^{*}$ similar to (4).
-Differently, we use the local Hessian matrix ~$\widetilde{\mathcal{H}}_{k,h}$ to update the estimation instead of the global lower bound.
-Specifically, we estimate parameter $\widetilde{\theta}_{k,h}$ as follows:
+In Property 1, we show the Hessian matrix $H_{k,h}$ is lower bounded by a positive definite matrix $\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ uniformly over the parameter domain, as in (3).
+This quantity measures the exploration degree across different states and is used to update the parameter estimation and construct the confidence set.
+However, the bound is not tight, as the Hessian matrix can be significantly larger in certain regions, away from the global minimum.
+This observation motivates us to design a local learning algorithm that can adaptively leverage local information for improved guarantee.
+Inspired by recent advances in multinomial logistic bandit (MLogB) (Zhang and Sugiyama, 2023) and multinomial logit contextual bandit (MNL) (Lee and Oh, 2024), we run an online mirror descent algorithm to estimate the parameter $\theta_{h}^{*}$.
+Differently from (4), we use the local Hessian matrix ~$\widetilde{\mathcal{H}}_{k,h}$ to update the estimation instead of the global lower bound.
+Specifically, we estimate $\widetilde{\theta}_{k,h}$ as follows:
 $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$ (9)
 where $\eta$ is the step size, $\mathcal{H}_{k,h}=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d}$ and ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$.
-Note both $\mathcal{H}_{k,h}$ and ~$\widetilde{\mathcal{H}}_{k,h}$ can be updated incrementally and the optimization problem in (8) can be solved efficiently.
-This step can be solved by the following two-step update.
-$\theta^{\prime}_{k+1,h}=\theta_{k,h}-\eta\mathcal{H}_{k,h}^{-1}g_{k,h}(\theta_{k,h}),\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$ (9)
+Note that both $\mathcal{H}_{k,h}$ and ~$\widetilde{\mathcal{H}}_{k,h}$ can be updated incrementally.
+Similar to the update in (4), the optimization problem in (9) can be efficiently solved using the following two-step update:
+$\theta^{\prime}_{k+1,h}=\theta_{k,h}-\eta\widetilde{\mathcal{H}}_{k,h}^{-1}g_{k,h}(\theta_{k,h}),\quad\mbox{and}\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$
+This two-step update procedure incurs a computational cost of $\mathcal{O}(d^{3}U)$ and a storage cost of $\mathcal{O}(d^{2})$ per episode, both independent of the episode count $k$.
 Based on this estimator, we can construct the $\kappa$-independent confidence set as follows.
@@
 For any $\delta\in(0,1)$, define the confidence set as
-$\widehat{\mathcal{C}}_{k,h}=\big{\{}\theta\in\Theta\mid\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}\leq\widetilde{\beta}_{k}\big{\}}$
-with $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$. Then, $\Pr[\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
-Based on Lemma 3, we can construct the optimistic value function similar to (6) as
-$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\mathcal{H}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (10)
-The detailed algorithm is presented in UCRL-MNL-LL in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL by the following theorem.
-Theorem 2.
-Suppose Assumption 1 holds.
-For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, and $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, with probability at least $1-\delta$, UCRL-MNL-LL algorithm ensures the following regret guarantee
-${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K}).$
-Compared with the result in Theorem 1, Theorem 2 enhances the dependence on $\kappa$ from $\kappa^{-1}$ to $\kappa^{-1/2}$. This is achieved by utilizing local information of the Hessian matrix.
-Algorithm 2 UCRL-MNL-LL &\&& UCRL-MNL-LL+
+$\widetilde{\mathcal{C}}_{k,h}=\big{\{}\theta\in\Theta\mid\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}\leq\widetilde{\beta}_{k}\big{\}},$
+where $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$. Then, we have $\Pr[\theta_{h}^{*}\in\widetilde{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
+Remark 5.
+Zhang and Sugiyama (2023) studied the multiple-parameter MLogB model, where the unknown parameter is a matrix.
+Consequently, the confidence set in Theorem 3 of their work exhibits a polynomial dependence on the number of possible outcomes, which corresponds to the number of reachable states $U$ in our setting.
+This dependence is acceptable in the bandit setting while is not suitable for the MDP setting.
+In contrast, Lee and Oh (2024) focused on the single-parameter MNL model and revisited the self-concordant-like property, demonstrating that the log-loss of the single-parameter MNL model is $3\sqrt{2}$-self-concordant-like (Proposition B.1 in Lee and Oh (2024)).
+This property is crucial for the improved confidence set in Lemma 3 that is independent of $\kappa$ and $U$.
+Algorithm 2 UCRL-MNL-LL
 0:  Step size $\eta$, regularization parameter $\lambda$, confidence width $\widetilde{\beta}_{k}$, confidence parameter $\delta$.
@@
 2:  for $k=1,\ldots,K$ do
-3:     UCRL-MNL-LL: Compute $\bar{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (10).
-4:     UCRL-MNL-LL+: Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (12).
-5:     for $h=1,\ldots,H$ do
-6:        UCRL-MNL-LL: Observe $s_{k,h}$, select $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\bar{Q}_{k,h}(s_{k,h},a)$.
-7:        UCRL-MNL-LL+: Observe $s_{k,h}$, select $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{Q}_{k,h}(s_{k,h},a)$.
-8:        Update ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$.
-9:        Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}$ as (8).
-10:        Update $\mathcal{H}_{k,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
-11:     end for
-12:  end for
+3:     Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (11).
+4:     for $h=1,\ldots,H$ do
+5:         Observe state $s_{k,h}$, select action $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{Q}_{k,h}(s_{k,h},a)$.
+6:         Update ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$.
+7:         Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}$.
+8:         Update $\mathcal{H}_{k,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
+9:     end for
+10:  end for
 5.2 Improved Optimistic Value Function Construction
-In Section 5.1, we improve the dependence on $\kappa$ by constructing a tighter $\kappa$-independent confidence set.
-However, the regret of UCRL-MNL-LL in Theorem 2 still have a dependence on $\kappa^{-1/2}$.
-In this part, we design an improved algorithm that removes the dependence on $\kappa$ in the main term.
-By a similar analysis as (7), the regret of UCRL-MNL-LL can be bounded as:
-$\operatorname*{Reg}(K)\leq 2H\widetilde{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}.$ (11)
-Though $\widetilde{\beta}_{K}$ now is independent of $\kappa$ and $\mathcal{H}_{k,h}^{-1}$ preserves the local information in (11), the norm $\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is still in a global manner due to the maximum operation, leading to a $\kappa^{-1/2}$ dependence.
-To address this issue, we propose a new construction of the optimistic value function through a more refined analysis.
-Specifically, we employ a second-order Taylor expansion to more accurately bound the value difference arising from the estimation error in the following lemma.
+Based on the confidence set in Lemma 3, a natural choice for the optimistic value function construction is analogous to (7) and can be expressed as:
+$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\mathcal{H}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (10)
+However, though $\widetilde{\beta}_{K}$ now is independent of $\kappa$ and $\mathcal{H}_{k,h}^{-1}$ preserves the local information in (10), the norm $\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is still in a global manner due to the maximum operation, leading to a $\kappa^{-1/2}$ dependence.
+To address this issue, we propose a new construction of the optimistic value function.
+Specifically, we employ a second-order Taylor expansion to more accurately bound the value difference arising from transition estimation errors.
+Although the idea of using second-order Taylor expansion has been explored in bandits (Périvier and Goyal, 2022; Zhang and Sugiyama, 2023; Lee and Oh, 2024), fundamental differences arise in the MDP setting.
+Specifically, Périvier and Goyal (2022) studied the uniform revenue setting, where the reward is identical for all actions.
+Zhang and Sugiyama (2023) and Lee and Oh (2024) focused on the non-uniform setting, but the rewards for different actions are known a priori.
+However, in the MDP setting, the value function is state-dependent and unknown to the learner, leading to a more challenging problem.
+Moreover, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize the expected revenue, whereas in MNL MDPs, the goal is to choose one action at each stage to maximize the cumulative reward.
+Due to these differences, the analyses used in the bandit setting cannot be directly applied to the MDP setting.
+For MDPs, we show the value difference arising from the transition estimation error as follows.
 Lemma 4.
-Suppose Lemma 3 holds. For any function $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, with probability at least $1-\delta$, it holds that
-$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}},$
+Suppose Lemma 3 holds. For any $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds
+$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}.$
 where
-$\epsilon_{s,a}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\quad\epsilon_{s,a}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
+$\epsilon_{s,a}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\epsilon_{s,a}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
 Based on Lemma 4, we construct the optimistic value function as follows:
-$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]}.$ (12)
-The detailed algorithm is presented in UCRL-MNL-LL+ in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL+ in the following theorem.
+$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]}.$ (11)
+In contrast to the value function specified in (10), where the term $\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is utilized, the refined value function introduced in (11) substitutes this term with $\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}$.
+This adjustment better preserves local information, offering a more precise and $\kappa$-independent estimation error bound.
+At state $s_{k,h}$, our algorithm chooses action $a_{k,h}=\max\widetilde{Q}_{k,h}(s_{k,h},\cdot)$.
+The detailed algorithm is presented in Algorithm 2.
+We show the regret guarantee of UCRL-MNL-LL in the following theorem.
 Theorem 2.
-Suppose Assumption 1 holds.
-For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$ and $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, with probability at least $1-\delta$, UCRL-MNL-LL+ algorithm ensures the following regret guarantee
+For any $\delta\in(0,1)$, set $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, $\eta=\frac{1}{2}\log(1+U)+(B+1)$ and $\lambda=84\sqrt{2}\eta(B+d)$, with probability at least $1-\delta$, UCRL-MNL-LL algorithm ensures the following regret guarantee
 ${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2}\big{)}.$
 Remark 6.
-Theorem 3 moves the exponentially small $\kappa$-dependence to the lower-order term, which is independent of $K$.
-Compared with the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023), the higher-order term only differs by a factor of $H^{1/2}$, almost matching the same computational and statistical efficiency simultaneously.
+The high-order term in Theorem 2 is now independent $\kappa$, significantly improving the statistical efficiency compared with Theorem 1.
+In comparison with the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear cases (Zhou et al., 2021; He et al., 2023), the higher-order term in Theorem 2 only differs by a factor of $H^{1/2}$, almost matching the same computational and statistical efficiency simultaneously.
 6 Lower Bound
-In this section, we establish the first lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and logistic bandits.
-Consider the following logistic bandit problem (Faury et al., 2020): at each round $k\in[K]$, the learner selects an action $x_{k}\in\mathcal{X}$ and receives a reward $r_{k}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d}\mid\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter.
-The learner aims to minimize the regret: ${\operatorname*{Reg}}^{\mathtt{LB}}(K)=\max_{x\in\mathcal{X}}\sum_{k=1}^{K}\mu(x^{\top}\theta^{*})-\sum_{k=1}^{K}\mu(x_{k}^{\top}\theta^{*})$.
-Theorem 4.
-For any logistic bandit $\mathcal{B}$, there exists an MNL mixture MDP $\mathcal{M}$ such that learning $\mathcal{M}$ is as hard as learning $H/2$ independent instances of $\mathcal{B}$ simultaneously.
+In this section, we establish a lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and the logistic bandit problem.
+Consider the following logistic bandit problem (Faury et al., 2020; Abeille et al., 2021): at each round $k\in[K]$, the learner selects an action $x_{k}\in\mathcal{X}$ and receives a reward $r_{k}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter.
+The learner aims to to minimize the regret:
+${\operatorname*{Reg}}^{\mathtt{LogB}}(K)=\max_{x\in\mathcal{X}}\sum_{k=1}^{K}\mu(x^{\top}\theta^{*})-\sum_{k=1}^{K}\mu(x_{k}^{\top}\theta^{*}).$
+Theorem 3 (Lower Bound).
+For any logistic bandit problem $\mathcal{B}$, there exists an MNL mixture MDP $\mathcal{M}$ such that learning $\mathcal{M}$ is as hard as learning $H/2$ independent instances of $\mathcal{B}$ simultaneously.
 Corollary 1.
@@
 (2021).
-To our knowledge, a lower bound for logistic bandits with finite arms has not been established, which is beyond the scope of this work.
-Consequently, it remains open for the lower bound of MNL mixture MDPs with finite action space.
-Remark 5.
-While Corollary 1 is not a strict lower bound for MNL mixture MDPs with finite action space, it gives a basis to conjecture our results may be optimal in $d$ and $K$.
+To the best of our knowledge, a lower bound for logistic bandits with finite arms has not been established, which is beyond the scope of this work.
+This absence leaves the lower bound for MNL mixture MDPs with a finite action space open through this reduction.
+However, after the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al.
+(2024) proposed a new reduction that bridges MNL mixture MDPs with linear mixture MDPs by approximating multinomial logit functions to linear functions, employing the mean value theorem.
+Leveraging this new reduction, they established an $\Omega(dH^{3/2}\sqrt{K})$ lower bound for MNL mixture MDPs with the finite action space.
+This achievement confirms that our result is indeed optimal with respect to the dependence on $d$ and $K$, only differing by a $\mathcal{O}(H^{1/2})$ factor.
 Dependence on $H$.
-MNL mixture MDPs can be viewed as a generalization of linear mixture MDPs (Ayoub et al., 2020; Zhou et al., 2021).
+By the discussion in Remark 7, we note that our result is optimal with respect to $d$ and $K$, but loosing by a $\mathcal{O}(H^{1/2})$ factor.
+We discuss the challenges in improving the dependence on $H$.
+Notably, MNL mixture MDPs can be viewed as a generalization of linear mixture MDPs (Ayoub et al., 2020; Zhou et al., 2021).
 The pioneering work by Ayoub et al.
-(2020) achieved an $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ regret for linear mixture MDPs, which matches our results in Theorem 3, differing only on the lower-order term.
-Zhou et al.
-(2021) enhanced the dependence on $H$ and attained an optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$.
-This was made possible by recognizing that the value function in linear mixture MDPs is also linear, allowing for direct learning of value functions while incorporating their variance information.
-In contrast, the value function does not conform to a logistic model for MNL mixture MDPs, posing a significant challenge in using variance information of value functions.
-It remains open whether similar improvements in $H$ are attainable for MNL mixture MDPs.
-Dependence on $\kappa$.
-Though we fully remove the $\kappa$ dependence in Theorem 3 in the higher-order term and move it to the lower order term, Corollary 1 suggests that there may still be some room for improvement in the dependency on $\kappa$.
-The $\widetilde{\mathcal{O}}(d\sqrt{K\kappa^{*}})$ dependence has been established in the binary logistic bandit problem by Abeille et al.
-(2021).
-However, their techniques are limited to the binary case and are not directly applicable to MNL mixture MDPs which have multinomial rewards.
-For the multinomial feedback settings, Périvier and Goyal (2022) studied the MNL contextual bandit problem and achieved a regret bound of $\widetilde{\mathcal{O}}(d\sqrt{K\kappa^{*}}+\kappa^{-1})$.
-However, their results only hold for the scenarios with uniform revenue, where each choice yields identical rewards.
-In contrast, MNL mixture MDPs are more challenging as they involve diverse revenue values, represented by the value functions of different states.
-Thus, their techniques also can not be applied to MNL mixture MDPs.
-It remains open whether the $\widetilde{\mathcal{O}}(dH\sqrt{K\kappa^{*}})$ result can be achieved for MNL mixture MDPs.
+(2020) achieved a regret bound of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ for linear mixture MDPs, which matches our results in Theorem 2, differing only on the lower order term.
+Later, Zhou et al.
+(2021) enhanced the dependence on $H$ and attained an optimal regret bound of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$.
+This was made possible by recognizing that the value function in linear mixture MDPs is linear, allowing for direct learning of the value function while incorporating variance information.
+In contrast, the value function for MNL mixture MDPs does not conform to a specific structure, posing a significant challenge in using the variance information of value functions.
+Thus, it remains open whether similar improvements on $H$ are attainable for MNL mixture MDPs.
 7 Conclusion and Future Work
-In this work, we address both the computational and statistical challenges for MDPs with multinomial logit function approximation.
-We first propose an algorithm that attains the previously best-known $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode.
-Next, we identify that the statistical efficiency can be further improved by exploiting the local information.
-We propose two algorithms that not only maintain $\mathcal{O}(1)$ computational and storage costs but achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$, respectively.
-We also establish the first lower bound, justifying the optimality of our results in $d$ and $K$.
-To the best of our knowledge, this is the first work that achieves almost the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation.
+In this work, we study MNL mixture MDPs that employ multinomial logit function approximation to ensure valid probability distributions over the state space.
+We address both the computational and statistical challenges for this problem.
+Specifically, we first propose an algorithm based on the online Newton step that attains the $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode.
+Next, we propose an enhanced algorithm that leverages local information to improve the statistical efficiency.
+It not only maintains $\mathcal{O}(1)$ computational and storage costs but also achieves an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, nearly matching the result of linear function approximation from both computational and statistical perspectives.
+Finally, we establish the first lower bound for MNL mixture MDPs, justifying the optimality of our results in $d$ and $K$.
 There are several interesting directions for future work.
 First, there still exists a gap between the upper and lower bounds and how to close this gap is an open problem.
-Besides, while this work focuses on stochastic rewards, extending this model to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Fei et al., 2020; Li et al., 2023) is another interesting direction.
+Besides, while this work focuses on stochastic rewards, extending this model to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Zhao et al., 2022; Li et al., 2023) is an another important direction.
+Acknowledgments
+This research was supported by National Science and Technology Major Project (2022ZD0114800) and NSFC (U23A20382, 62206125). Peng Zhao was supported in part by the Xiaomi Foundation.
 Appendix A Properties of the Multinomial Logit Function
-This section collects several key properties of the multinomial logit function.
+This section collects several key properties of the multinomial logit function used in the paper.
 Without loss of generality, we assume $\forall\mathcal{S}_{h,s,a},\exists s^{\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi(s^{\prime}\>|\>s,a)=\mathbf{0}$.
@@
 Recall the following definitions in the paper:
-$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$
+$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),$
+$g_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$
 $H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$
+$\widehat{\Sigma}_{k,h}$ $=\frac{\kappa}{2}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}+\lambda I_{d},$
+$\mathcal{H}_{k,h}$ $=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d},$
 Lemma 5.
@@
 Proof.
-Note that
+First, note that
 $\forall x,y\in\mathbb{R}^{d},(x-y)(x-y)^{\top}=xx^{\top}+yy^{\top}-xy^{\top}-yx^{\top}\succeq 0\Longrightarrow xx^{\top}+yy^{\top}\succeq xy^{\top}+yx^{\top}.$
@@
 Suppose $\lambda\geq 1$, define $\widetilde{\phi}_{k,h}^{s^{\prime}}=\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime\prime}}$, for any $k\in[K],h\in[H]$, the following statements hold:
-$(1)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}^{2}\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)$
-$(2)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
-$(3)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
-$(4)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
-$(5)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
+$(\textnormal{I})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}^{2}\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)$
+(II$)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)( I$
+$(\textnormal{III})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
+$(\textnormal{IV})\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
+$(\textnormal{V})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
 Proof.
-We prove the three statements individually.
-Proof of statement (1).
+We prove the five statements individually.
+Proof of statement (I).
 By the definition that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$, we have
@@
 $\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\min\bigg{\{}1,\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\bigg{\}}$
-$\leq\frac{4}{\kappa}\sum_{i=1}^{k}\log\Bigg{(}1+\min\bigg{\{}1,\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\bigg{\}}\Bigg{)}$
 $\leq\frac{4}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{i,h}^{-1}}\Big{)}$
 $\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right),$
-where the first inequality holds by the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$ and the last inequality holds by Equation (13).
-Proof of statement (2).
+where the first inequality holds by the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$ and the last inequality holds by (12).
+Proof of statement (II).
 By Lemma 5, we have $H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$.
@@
 […]
-Proof of statement (3).
-By Lemma 5, we have $\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k+1,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$.
+Proof of statement (III).  By Lemma 5, we have
+$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k+1,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
 Since $\lambda\geq 1$, we have $\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\leq\kappa$.
@@
 $\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right).$
-Proof of statement (4).
-By the definition of $H_{k,h}(\theta)$, we have
+This finishes the proof.
+Proof of statement (IV).  By the definition of $H_{k,h}(\theta)$, we have
 $H_{i,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)p_{i,h}^{s^{\prime\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime\prime}})^{\top}$
@@
 The remaining proof is similar to the proof of statement (2).
-Proof of statement (5).
-By (14), we have
+Proof of statement (V).  By (13), we have
 $H_{i,h}(\widetilde{\theta}_{i+1,h})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\widetilde{\theta}_{i+1,h})(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}.$
@@
 Proof.
-We define
+The main proof follows the proof of Theorem 1 in Zhang et al. [2016]. We define
 $\bar{f}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[f_{k,h}(\theta)\mid\mathcal{F}_{k,h}],\quad\bar{g}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[g_{k,h}(\theta)\mid\mathcal{F}_{k,h}].$
@@
 $\leq\lambda B^{2}+\frac{8}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)+(\frac{8}{\kappa}+16B)\gamma_{k},$
-where the second inequality holds by $\lVert\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq 1$ and $\widehat{\Sigma}_{i,h}\geq I,\forall i\in[K]$, the last inequality is by the elliptical potential in Lemma 6.
+where the second inequality holds by $\lVert\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq 1$ and $\widehat{\Sigma}_{i,h}\geq I,\forall i\in[K]$, and the last inequality is by Lemma 6.
 This finishes the proof.
@@
 By the definition of $\bar{f}_{k,h}(\theta)$, we have
-$\bar{f}_{k,h}(\theta)-\bar{f}_{k,h}(\theta_{h}^{*})$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log p_{k,h}^{s^{\prime}}(\theta)+\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\geq 0,$
+$\bar{f}_{k,h}(\theta)-\bar{f}_{k,h}(\theta_{h}^{*})$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log\frac{p_{k,h}^{s^{\prime}}(\theta_{h}^{*})}{p_{k,h}^{s^{\prime}}(\theta)}\geq 0,$
 where the last inequality is due to $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log\frac{p_{k,h}^{s^{\prime}}(\theta_{h}^{*})}{p_{k,h}^{s^{\prime}}(\theta)}$ is the Kullback-Leibler divergence between $p_{k,h}(\theta_{h}^{*})$ and $p_{k,h}(\theta)$, which always is non-negative.
@@
 $(x_{i}-x_{j})^{\top}Z(x_{i}-x_{j})=x_{i}^{\top}Zx_{i}-x_{i}^{\top}Zx_{j}-x_{j}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq 0,\quad\forall x_{i},x_{j}\in\mathbb{R}^{d}.$ (22)
-which implies $x_{i}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq x_{i}^{\top}Zx_{j}+x_{j}^{\top}Zx_{i},\forall x_{i},x_{j}\in\mathbb{R}^{d}$.
+which implies $x_{i}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq x_{i}^{\top}Zx_{j}+x_{j}^{\top}Zx_{i},\quad\forall x_{i},x_{j}\in\mathbb{R}^{d}$.
 Let $x_{i}=(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
@@
 Proof.
-Notice that $\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$ is a martingale difference sequence. Also, we have
-$\left|\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$ $\leq\left|\left(\bar{g}_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|+\left|\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
+First, notice that $\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$ is a martingale difference sequence. Also, we have
+$\quad\left|\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
+$\leq\left|\left(\bar{g}_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|+\left|\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
 $\leq\lVert\bar{g}_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}+\lVert g_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}$
@@
 $\leq\sum_{i=1}^{k}\mathbb{E}_{y_{k,h}}\left[\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})^{2}\right]$
-$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})^{2}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\triangleq A_{k,h}.$
+$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})^{2}$
+$=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}$
+$\triangleq A_{k,h}.$
 where the first inequality is due to the fact that $\mathbb{E}[(\xi-\mathbb{E}[\xi])^{2}]\leq\mathbb{E}[\xi^{2}]$ for any random variable $\xi$.
@@
 $=\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta_{i})-y_{i,h}^{s^{\prime}})(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})$
-$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lvert(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})\rvert\leq\sqrt{kU\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}((\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*}))^{2}}\leq 2B.$
+$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lvert(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})\rvert$
+$\leq\sqrt{kU\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}((\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*}))^{2}}$
+$\leq 2B.$
+where the second equality is due to the definition of $\bar{g}_{i,h}(\theta_{i})$ and $g_{i,h}(\theta_{i})$, the first inequality holds by $p_{i,h}^{s^{\prime}}(\theta_{i})-y_{i,h}^{s^{\prime}}\in[-1,1]$, the second inequality holds by the Cauchy-Schwarz inequality, and the last inequality holds by the condition of case (i).
 Case (ii): $A_{k,h}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{j,h}}^{2}>\frac{4B^{2}}{kU}$.
@@
 […]
+[…]
 where $m=2\log_{2}(kU)$, and the last inequality follows the Bernstein inequality for martingales.
@@
 $\leq 2\sqrt{\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\log\frac{2k^{2}H\log(kU)}{\delta}}+8B\left(1+\log\frac{2k^{2}H\log(kU)}{\delta}\right)$
-$=2\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}+8B\left(1+\gamma_{k}\right),$
+$=2\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}+8B\left(1+\gamma_{k}\right),~{}$ (23)
 where $\gamma_{k}=\log\frac{2k^{2}H\log(kU)}{\delta}$.
 Then, applying $uv\leq cu^{2}+v^{2}/(4c)$ for any $c,u,v>0$ with $c=2/\kappa$, we have
-$\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}\leq\frac{2\gamma_{k}}{\kappa}+\frac{\kappa}{8}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}$
-Then we have
+$\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}\leq\frac{2\gamma_{k}}{\kappa}+\frac{\kappa}{8}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}.$ (24)
+Combining (23) and (24), we have
 $\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\leq\frac{\kappa}{4}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}+(\frac{4}{\kappa}+8B)\gamma_{k}.$
@@
 Proof.
-The gradient of $p_{s,a}(\theta)$ is given by $\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}$.
+The gradient of $p_{s,a}(\theta)$ is given by
+$\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}.$
 By the mean value theorem, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
 $\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}$
+=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})\bigg{|}$
 =\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})\bigg{|}$
 =\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})\Big{(}V(s^{\prime})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})V(s^{\prime\prime})\Big{)}\bigg{|}$
+[…]
 […]
@@
 […]
-[…]
 Define $\mathcal{M}_{k,h}=\mathbb{P}_{h}(\widehat{V}_{k,h+1}-V_{h+1}^{\pi_{k}})(s_{k,h},a_{k,h})-(\widehat{V}_{k,h+1}-V_{h+1}^{\pi_{k}})(s_{k,h+1})$. Applying this recursively, we have
@@
 $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{f}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}.$ (28)
-Then, we present the following lemma.
+Then, we present the following lemma that bounds the estimation error of the update rule in (28).
 Lemma 12 (Lemma E.1 of Lee and Oh [2024]).
-For the update rule defined in (27) as
-$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{f}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}},$
-let $\eta=\frac{1}{2}\log(U+1)+(B+1)$, we have
+For the update rule defined in (28), it holds that
 $\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}\leq$ $2\eta\left(\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})\right)+4\lambda B$
 $+12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}-\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}.$
-Then, we bound the right-hand side of the above lemma separately. First, we bound the first term. Define the softmax function as follows.
+We bound the right-hand side of the above lemma separately in the following.
+First, we bound the first term. Define the softmax function as follows.
 $[\sigma_{k,h}(z)]_{s}=\frac{\exp([z])_{s}}{1+\sum_{s\in\dot{\mathcal{S}}_{k,h}}\exp([z])_{s}},\quad\forall s\in\mathcal{S}_{k,h}.$
@@
 Then, we decompose the first term as follows.
-$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})=\underbrace{\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})}_{(a)}+\underbrace{\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})}_{(b)}$
+$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})$
+=\ = $\underbrace{\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})}_{(a)}+\underbrace{\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})}_{(b)}$
 where $z_{k,h}=\sigma^{-1}_{k,h}(\mathbb{E}_{\theta\sim P_{k,h}}[\sigma_{k,h}((\phi_{k,h}^{s^{\prime}})^{\top}\theta)_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}])$, $P_{k,h}\triangleq\mathcal{N}(\widetilde{\theta}_{k,h},(1+c\mathcal{H}_{k,h}^{-1}))$ is the Gaussian distribution with mean $\widetilde{\theta}_{k,h}$ and covariance $(1+c\mathcal{H}_{k,h}^{-1})$ where $c$ is a constant to be specified later.
@@
 […]
-Thus, by setting $\eta=\frac{1}{2}\log(U+1)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, we have $\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}\leq\mathcal{O}\big{(}\sqrt{d}\log U\log(KH/\delta)\big{)}\triangleq\widetilde{\beta}_{k}$.
-This finishes the proof.
-∎
-Appendix F Proof of Theorem 2
+[…]
+Thus, by setting $\eta=\frac{1}{2}\log(U+1)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, we have
+$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}\leq\mathcal{O}\big{(}\sqrt{d}\log U\log(KH/\delta)\big{)}\triangleq\widetilde{\beta}_{k}.$
+This finishes the proof. ∎
+Appendix F Proof of Lemma 4
 Proof.
-By the same analysis as the proof of Lemma 2 in Appendix C, we have
-$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq$ $H\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\phi_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h}-\theta_{h}^{*})$
-[…]
-[…]
-where the first inequality holds by (24), the second inequality is by Holder’s inequality, and the third inequality is by Lemma 3.
-Then, substituting $\Gamma_{h,s,a}=H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ into Lemma 10, we have
-$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})\leq$ $2H\widetilde{\beta}_{k}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}+H\sqrt{2KH\log(2/\delta)}$
-[…]
-[…]
-[…]
-where the second inequality holds by the Cauchy-Schwarz inequality, and the third inequality is by Lemma 6. This finishes the proof. ∎
-Appendix G Proof of Lemma 4
-Proof.
-By the second-order Taylor expansion, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widetilde{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
-${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}$
+By the second-order Taylor expansion, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
+${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}$
 =\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})^{\top}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})$
-The gradient of $p_{s,a}(\theta)$ is given by $\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}$.
+The gradient of $p_{s,a}(\theta)$ is given by
+$\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}.$
 For the first-order term, we have
@@
 […]
-where the inequality holds by $V(s)\in[0,H],\forall s$. According to Lemma 18, we have (omit the subscript $\mathcal{S}_{h,s,a}$ for simplicity):
+where the inequality holds by $V(s)\in[0,H],\forall s$.
+According to Lemma 18, we have (omit the subscript $\mathcal{S}_{h,s,a}$ for simplicity):
 $\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
@@
 […]
-where the last inequality holds by Lemma 3.
-Combining (28) and (31) finishes the proof. ∎
-Appendix H Proof of Theorem 3
+where the last inequality holds by Lemma 3. Combining (29) and (32) finishes the proof. ∎
+Appendix G Proof of Theorem 2
 Proof.
@@
 where
-$\epsilon_{k,h}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
+$\epsilon_{k,h}^{\mathtt{fst}}$ $=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},$
+$\epsilon_{k,h}^{\mathtt{snd}}$ $=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
 Next, we bound $\epsilon_{k,h}^{\mathtt{fst}}$ and $\epsilon_{k,h}^{\mathtt{snd}}$ respectively.
-Bounding $\epsilon_{k,h}^{\mathtt{fst}}$.
-For simplicity, we denote
+Bounding $\epsilon_{k,h}^{\mathtt{fst}}$.  For simplicity, we denote
 $\mathbb{E}_{\theta}[\phi_{k,h}^{s^{\prime}}]=\mathbb{E}_{s^{\prime}\sim p_{k,h}^{s}(\theta)}[\phi_{k,h}^{s^{\prime}}],\quad\bar{\phi}_{s,a}^{s^{\prime}}=\phi_{s,a}^{s^{\prime}}-\mathbb{E}_{\widetilde{\theta}_{k,h}}[\phi_{k,h}^{s^{\prime}}],\quad\widetilde{\phi}_{s,a}^{s^{\prime}}=\phi_{s,a}^{s^{\prime}}-\mathbb{E}_{\widetilde{\theta}_{k+1,h}}[\phi_{k,h}^{s^{\prime}}]$
@@
 =\ = $\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(c)}+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(d)}$
-$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(e)}$
-Next, we bound these terms separately.
+$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(e)}.$
+We bound these terms separately in the following.
 For the first term $(c)$, we have
@@
 […]
-where the fifth inequality is by Cauchy-Schwarz inequality and the fourth inequality is because by Lemma 17, since ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
+[…]
+where and the fifth inequality is by Cauchy-Schwarz inequality and the fourth inequality is because by Lemma 17, since ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
 $\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}}\leq\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\widetilde{\mathcal{H}}_{k,h}}\leq 2\eta\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{\widetilde{\mathcal{H}}_{k,h}^{-1}}\leq\frac{2\eta}{\sqrt{\lambda}}\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2},$
@@
 […]
+$\hskip 56.9055pt+\frac{2\eta}{\sqrt{\lambda}}\frac{\left(\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\right)^{2}+\left(\max_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\right)^{2}}{2}$
 […]
@@
 […]
+[…]
 where the last inequality holds by Lemma 6.
 Finally, we bound the term $(e)$ as follows.
-$\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\sqrt{\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})}\sqrt{\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}}$
+$\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
 […]
-where the last inequality holds by Lemma 16.
+[…]
+where the first inequality holds by the Cauchy-Schwarz inequality and the last holds by Lemma 6.
 Thus, combining (33), (34), and (35), we have
@@
 $\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{16\eta}{\kappa\sqrt{\lambda}}d\log(1+\frac{K}{d\lambda})\right)$ (36)
-Bounding $\epsilon_{k,h}^{\mathtt{snd}}$.
 For the second-order term, by Lemma 6, we have
-$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (36)
+$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}$ $=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}$
+$\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (37)
+where the last inequality holds by Lemma 6.
 Combining (36) and (37), we have
@@
 This finishes the proof. ∎
-Appendix I Proof of Theorem 4
+Appendix H Proof of Theorem 3
 Proof.
-Our proof is similar to the proof of the lower bound for linear mixture MDPs in Zhao et al. [2023]. We prove this lemma by translating this problem to learning $H/2$ logistic bandit problems.
+Our proof is similar to the proof of the lower bound for adversarial linear mixture MDPs with the unknown transition in Zhao et al.
+[2023].
+We prove this lemma by reducing the MNL mixture MDP problem to a sequence of binary logistic bandit problems.
 Specifically, we use each three layers to construct a block.
@@
 The probability of transferring to state $s_{i,2}^{*}$ when executing action $a$ at state $s_{i,1}$ is $\rho_{a}$, in particular,
-$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=\rho_{a},\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)=1-\rho_{a}.$
-For the second layer, $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$.
-The reward satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second layer for all $a\in\mathcal{A}$.
-Consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from Bernoulli distributed with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*})^{-1})$.
-Thus, learning in this block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$.
+$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a))^{\top}\theta_{i,1}^{*})}=\rho_{a},$
+$\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)$ $=\frac{1}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=1-\rho_{a}.$
+For the second layer, the MDP instance satisfies that $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$.
+The reward function satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second layer for all $a\in\mathcal{A}$.
+Consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from Bernoulli distributed with mean
+$\mu(x^{\top}\theta^{*})=\frac{1}{1+\exp(x^{\top}\theta^{*})}.$
+By this configuration, we can see that learning in this block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$.
 It is also clear that the optimal policy at state $s_{i,1}$ is to select action $a_{i,1}^{*}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$.
-Thus, learning this MDP is equal to learning $H/2$ logistic bandit problems.
+Thus, learning this MNL mixture MDP equals to learning $H/2$ logistic bandit problems.
 This finishes the proof.
 ∎
-Appendix J Supporting Lemmas
-In this part, we provide several supporting lemmas used in the proofs of the main results.
+Appendix I Supporting Lemmas
+In this section, we provide several supporting lemmas used in the proofs of the main results.
 Lemma 15 (Abbasi-Yadkori et al. [2011, Theorem 1]).
@@
 Let $\left\{X_{t}\right\}_{t=1}^{\infty}$ be an $\mathbb{R}^{d}$-valued stochastic process such that $X_{t}$ is $\mathcal{F}_{t-1}$-measurable.
-Assume that $V\in\mathbb{R}^{d\times d}$ is a positive definite matrix.
-For any $t\geq 0$, define $V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top},S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}$.
+Assume that $V$ is a $d\times d$ positive definite matrix.
+For any $t\geq 0$, define
+$V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top}\quad S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}.$
 Then, for any $\delta>0$, with probability at least $1-\delta$, for all $t\geq 0$,
-$\left\|S_{t}\right\|_{V_{t}^{-1}}^{2}\leq 2R^{2}\log\left(\frac{\det\left(V_{t}\right)^{1/2}\det(V)^{-1/2}}{\delta}\right).$
+$\left\|S_{t}\right\|_{V_{t}^{-1}}^{2}\leq 2R^{2}\log\left(\frac{\operatorname{det}\left(V_{t}\right)^{1/2}\operatorname{det}(V)^{-1/2}}{\delta}\right).$
 Lemma 16 (Abbasi-Yadkori et al. [2011, Lemma 10]).
@@
 Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$.
-Then, we have $\det(V_{t})\leq(\lambda+{tL^{2}}/{d})^{d}$.
+Then, for any $1\leq s\leq t$, we have
+$\det(V_{t})\leq\left(\lambda+\frac{tL^{2}}{d}\right)^{d}.$
 Lemma 17 (Orabona [2019, Lemma 6.9]).
-Let $Z$ be a positive define matrix and $\mathcal{X}$ be a convex set, define $\widetilde{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\langle\mathbf{g},\mathbf{x}\rangle+\frac{1}{2\eta}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{Z}^{2}$.
+Let $Z$ be a positive define matrix and $\mathcal{X}$ be a convex set,
+$\widetilde{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\langle\mathbf{g},\mathbf{x}\rangle+\frac{1}{2\eta}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{Z}^{2}.$
 Then, for all $\mathbf{x}^{*}\in\mathcal{X}$, we have
-$\lVert\widetilde{\mathbf{x}}-\widehat{\mathbf{x}}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}\quad\mbox{and}\quad\langle\mathbf{g},\widehat{\mathbf{x}}-\mathbf{x}^{*}\rangle\leq\frac{\eta}{2}\lVert\mathbf{g}\rVert_{Z^{-1}}^{2}+\frac{1}{2\eta}\lVert\widehat{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}-\frac{1}{2\eta}\lVert\widetilde{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}.$
+$\lVert\widetilde{\mathbf{x}}-\widehat{\mathbf{x}}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}.$
+and
+$\langle\mathbf{g},\widehat{\mathbf{x}}-\mathbf{x}^{*}\rangle\leq\frac{\eta}{2}\lVert\mathbf{g}\rVert_{Z^{-1}}^{2}+\frac{1}{2\eta}\lVert\widehat{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}-\frac{1}{2\eta}\lVert\widetilde{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}.$
 Lemma 18 (Lee and Oh [2024, Lemma D.3]).
@@
 Let $p_{i}(\mathbf{u})=\frac{\exp\left(u_{i}\right)}{v+\sum_{k=1}^{K}\exp\left(u_{k}\right)}$.
-Then, for all $i\in[K]$,
+Then, for all $i\in[K]$, we have
 $\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|\leqslant\begin{cases}3p_{i}(\mathbf{u})&\text{ if }i=j,\\ 2p_{i}(\mathbf{u})p_{j}(\mathbf{u})&\text{ if }i\neq j.\end{cases}$

### v3

@@
 We study a new class of MDPs that employs multinomial logit (MNL) function approximation to ensure valid probability distributions over the state space.
-Despite its benefits, introducing the non-linear function raises significant challenges in both computational and statistical efficiency.
-The best-known result of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity, $d$ is the feature dimension, $H$ is the episode length, and $K$ is the number of episodes.
-While this result attains the same rate in $K$ as linear cases, the method requires storing all historical data and suffers from an $\mathcal{O}(K)$ computation cost per episode.
-Moreover, the quantity $\kappa$ can be exponentially small in the worst case, leading to a significant gap for the regret compared to linear function approximation.
-In this work, we first address the computational and storage issue by proposing an algorithm that achieves the same regret with only $\mathcal{O}(1)$ cost.
-Then, we design an enhanced algorithm that leverages local information to enhance statistical efficiency.
-It not only maintains an $\mathcal{O}(1)$ computation and storage cost per episode but also achieves an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, nearly closing the gap with linear function approximation.
-Finally, we establish the first lower bound for MNL function approximation, justifying the optimality of our results in $d$ and $K$.
+Despite its significant benefits, incorporating the non-linear function raises substantial challenges in both statistical and computational efficiency.
+The best-known result of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret upper bound, where $\kappa$ is a problem-dependent quantity, $d$ is the feature dimension, $H$ is the episode length, and $K$ is the number of episodes.
+However, we observe that $\kappa^{-1}$ exhibits polynomial dependence on the number of reachable states, which can be as large as the state space size in the worst case and thus undermines the motivation for function approximation.
+Additionally, their method requires storing all historical data and the time complexity scales linearly with the episode count, which is computationally expensive.
+In this work, we propose a statistically efficient algorithm that achieves a regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$, eliminating the dependence on $\kappa^{-1}$ in the dominant term for the first time.
+We then address the computational challenges by introducing an enhanced algorithm that achieves the same regret guarantee but with only constant cost.
+Finally, we establish the first lower bound for this problem, justifying the optimality of our results in $d$ and $K$.
 1 Introduction
@@
 Recently, there have been many efforts devoted to understanding the linear function approximation, yielding numerous valuable results (Yang and Wang, 2019; Jin et al., 2020; Ayoub et al., 2020).
-While these studies make important steps toward understanding RL with function approximation, there are still some challenges to be solved.
-In linear function approximation, transitions are assumed to be linear in specified feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs.
-Here $\mathbb{P}(s^{\prime}|s,a)$ is the probability from state $s$ to state $s^{\prime}$ taking action $a$, $\phi(s^{\prime}|s,a)$ and $\phi(s,a)$ are feature mappings, $\theta^{*}$ and $\mu^{*}(s^{\prime})$ are unknown parameters.
-However, a transition function is a probability distribution over states, i.e., the magnitude falls within the range of $[0,1]$, and the aggregation equals $1$.
-For certain feature mappings, linear transitions may not yield a valid probability distribution with arbitrary $\theta$ and $\mu$ as show by (Hwang and Oh, 2023).
-While there are also some works exploring generalized linear (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Foster et al., 2021; Chen et al., 2023), they made assumptions over value functions rather than transitions, hence do not tackle this challenge.
-Table 1: Comparison between previous works and ours in terms of the regret, computation cost and storage cost.
-Here $\kappa$ and $\kappa^{*}$ are exponentially small problem-dependent quantities defined in Assumption 1, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes.
-The computational cost and storage cost per episode indicate the dependence on episode count $k$.
-Reference Regret Com. Sto. MDP model
+While these studies make important steps toward understanding RL with function approximation, there are still challenges to be solved.
+In linear function approximation, transitions are assumed to be linear in feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs.
+Here $\mathbb{P}(s^{\prime}|s,a)$ is the probability from state $s$ to $s^{\prime}$ taking action $a$, $\phi(s^{\prime}|s,a)$ and $\phi(s,a)$ are feature mappings, $\theta^{*}$ and $\mu^{*}(s^{\prime})$ are unknown parameters.
+However, the transition function is a probability distribution over states, meaning its values must lie within $[0,1]$ and sum to 1.
+Thus, the linearity assumption is restrictive and hard to satisfy in practice.
+An algorithm designed for linear MDPs could break down entirely if the underlying MDP is not linear (Jin et al., 2020).
+While some works explore generalized linear (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Foster et al., 2021; Chen et al., 2023), they focus on function approximation for value functions rather than transitions, hence do not tackle this challenge.
+Table 1: Comparison between previous works and ours in terms of the regret and computational cost (including storage and time complexity) .
+Here $\kappa$ and $\kappa^{*}$ are problem-dependent quantities defined in Assumption 1, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes.
+The computational cost per episode only highlight the dependence on the episode count $k$.
+Reference Regret Storage Time MDP model
 Hwang and Oh (2023) $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ $\mathcal{O}(k)$ $\mathcal{O}(k)$ homogeneous
-UCRL-MNL-OL (Theorem 1) $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
-UCRL-MNL-LL (Theorem 2) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
-Lower Bound  (Corollary 1) $\Omega(dH\sqrt{K\kappa^{*}})$ / / infinite action space
+UCRL-MNL-LL (Theorem 1) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ $\mathcal{O}(k)$ $\mathcal{O}(k)$ inhomogeneous
+UCRL-MNL-OL (Theorem 2) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
+Lower Bound  (Corollary 1) $\Omega(dH\sqrt{K\kappa^{*}})$ infinite action space
 Towards addressing the limitation of linear function approximation, a new class of MDPs that utilizes multinomial logit function approximation has been proposed by Hwang and Oh (2023) recently.
-Despite its benefits, the introduction of non-linear functions raises significant challenges in both computational and statistical efficiency.
-Specifically, the best-known approach of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret.
-While matching the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ as linear function approximation in terms of $K$ (Zhou et al., 2021; He et al., 2023), their method requires storing all historical data and the computational cost per episode grows linearly with the episode count, which is expensive.
-Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases.
+Such formula also aligns better with models like neural networks (LeCun et al., 2015), which inherently respect the probabilistic constraints through a softmax layer and allow for greater expressive power.
+However, though it offers promising benefits, the introduction of non-linear functions introduces significant challenges in both statistical and computational efficiency.
+Specifically, the best-known approach of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity that measures the effective non-linearity over the entire parameter space, $d$ is the feature dimension, $H$ is the episode length, and $K$ is the number of episodes.
+Unfortunately, as we show in Claim 1, it holds that $\kappa^{-1}>U^{2}$, where $U$ denotes the maximum number of reachable states, which equals to the size of the state space $S$ in the worst case.
+This undermines the core motivation for function approximation, which aims to mitigate dependence on large state and action spaces.
+Furthermore, the method requires storing all historical data, and its time complexity per episode grows linearly with the episode count (i.e., $\mathcal{O}(k)$ at episode $k$).
 Thus, a natural question arises:
-Is MNL function approximation more difficult than linear function approximation for RL?
-In this paper, we answer this question by making significant advancements in both computational and statistical efficiency for MDPs with MNL function approximation, nearly closing the gap with linear function approximation.
-Table 1 presents a comparison between previous studies and our work.
-Specifically, our contributions are summarized as follows:
-  • We first propose the UCRL-MNL-OL algorithm based on a variant of Online Newton Step, attaining an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode.
-This result matches the best-known regret by Hwang and Oh (2023), yet achieves the same computational and storage efficiency as linear function approximation (Jin et al., 2020; Zhou et al., 2021).
-  • As the quantity $\kappa$ can be exponentially small, we propose an enhanced algorithm UCRL-MNL-LL, which leverages the local information to improve the statistical efficiency.
-It not only maintain $\mathcal{O}(1)$ computational and storage costs but achieve an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$.
-The higher-order term matches the optimal regret $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023) in $d$ and $K$, differing only by an $\mathcal{O}(H^{1/2})$ factor.
-  • We establish the first lower bound for MDPs with MNL function approximation by introducing a reduction to the logistic bandit problem.
-We prove a lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for MDPs with infinite action space.
-Though this does not constitute a strict lower bound for the finite actions case studied in this work, it suggests that our result may be optimal with respect to $d$ and $K$.
+Is it possible to design both statistically and computationally efficient algorithms
+for RL with MNL function approximation?
+In this work, we answer this question affirmatively for the class of MNL mixture MDPs where the transition is parameterized by a multinomial logit function. Our contributions are listed as follows:
+  • For statistical efficiency, we propose the UCRL-MNL-LL algorithm, which attains a regret bound of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$.
+Our result significantly improves upon the $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ rate of Hwang and Oh (2023), making the first time to achieve a $\kappa$-independent dominant term (note that the lower-order term still scales with $\kappa^{-1}$, but does not depend on $K$, making it acceptable).
+To achieve this, we propose a tighter confidence set based on a new Bernstein-type concentration (Périvier and Goyal, 2022) instead of the standard Hoeffding-type concentration, and exploit the self-concordant-like property (Bach, 2010) of the log-loss function to better use local information.
+  • For computational efficiency, we propose the UCRL-MNL-OL algorithm, which enjoys the same regret bound as UCRL-MNL-LL, but with only constant storage and time complexity per episode.
+This is enabled by recognizing that the negative log-likelihood function is exponentially concave, which motivates the use of online mirror descent with a specifically tailored local norm (Zhang and Sugiyama, 2023) to replace the standard maximum likelihood estimation.
+Furthermore, we construct the optimistic value function by incorporating a closed-form bonus term through a second-order Taylor expansion, thus avoiding the need to solve a non-convex optimization problem.
+  • We establish the first lower bound for MNL mixture MDPs by introducing a reduction to the logistic bandit problem.
+We prove a problem-dependent lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for infinite action setting, where $\kappa^{*}$ is an another problem-dependent quantity that measures the effective non-linearity over around the ground truth parameter.
+Though this does not constitute a strict lower bound for the finite action case studied in this work, it suggests that our result may be optimal in $d$ and $K$.
 [^1] After the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al.
-(2024) improved the lower bound to $\Omega(dH^{3/2}\sqrt{K})$.
-This confirms that our result is indeed optimal with respect to $d$ and $K$.
-From a technical perspective, inspired by recent advances in logistic bandits (Zhang et al., 2016; Faury et al., 2020; Zhang and Sugiyama, 2023), we observe that the negative log-likelihood function is exponentially concave, which motivates us to apply online Newton step to estimate the unknown transition parameter in an online manner.
-Moreover, we employ the Bernstein-like inequalities and the self-concordant-like property (Bach, 2010) of the log-loss to achieve the better dependence on $\kappa$.
+(2024) proved a lower bound of $\Omega(dH^{3/2}\sqrt{K})$ for the finite action setting.
+This confirms that our result is optimal in $d$ and $K$.
+Table 1 provides a comparison between our work and previous studies, focusing on regret and computational costs, including both storage and time complexity.
 Organization.
 We introduce the related work in Section 2 and present the setup in Section 3.
-Then, we design an algorithm with constant computational and storage costs in Section 4.
-Next, we present an algorithm with improved regret guarantee in Section 5.
+Then, we design a statistically efficient algorithm in Section 4.
+Next, we present an algorithm that achieves both statistical and computational efficiency in Section 5.
 Finally, we establish the lower bound in Section 6.
@@
 Notations.
-Denote by $[n]$ the set $\{1,\ldots,n\}$ and use $[x]_{[a,b]}$ to denote $\min(\max(x,a),b)$.
-For $X,Y\in\mathbb{R}^{d\times d}$, $X\succeq Y$ means $X-Y$ is positive semi-definite.
+We use $[x]_{[a,b]}$ to denote $\min(\max(x,a),b)$.
 For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, denote $\lVert\mathbf{x}\rVert_{A}=\sqrt{\mathbf{x}^{\top}A\mathbf{x}}$.
-The $\widetilde{\mathcal{O}}(\cdot)$ hides polylogarithmic factors.
+For a strictly convex and continuously differentiable function $\psi:\mathcal{W}\mapsto\mathbb{R}$, the Bregman divergence is defined as $\mathcal{D}_{\psi}\left(\mathbf{w}_{1},\mathbf{w}_{2}\right)=\psi\left(\mathbf{w}_{1}\right)-\psi\left(\mathbf{w}_{2}\right)-$ $\left\langle\nabla\psi\left(\mathbf{w}_{2}\right),\mathbf{w}_{1}-\mathbf{w}_{2}\right\rangle$.
+We use the notation $\mathcal{O}(\cdot)$ to indicate different types of dependencies depending on the context.
+For regret analysis, $\mathcal{O}(\cdot)$ omits only constant factors.
+For computational costs, we use $\mathcal{O}(\cdot)$ to solely highlight the dependence on the number of episode as this is the primary factor influencing the complexity.
+Additionally, we employ $\widetilde{\mathcal{O}}(\cdot)$ to hide all polylogarithmic factors.
 2 Related Work
@@
 (2021) investigated RL with generalized linear function approximation.
-Notably, unlike our approach which models transitions using a generalized linear model, they apply this approximation directly to the value function.
+Notably, unlike our approach that models transitions using a generalized linear model, they apply this approximation directly to the value function.
 Another line of works (Chowdhury et al., 2021; Li et al., 2022; Ouhamma et al., 2023) has studied RL with exponential function approximation and also aimed to ensure that transitions constitute valid probability distributions.
@@
 The most relevant work to ours is the recent work by Hwang and Oh (2023), which firstly explored a similar setting to ours, where the transition is characterized using a multinomial logit model.
-We significantly improve upon their results by providing computationally and statistically more efficient algorithms.
+We significantly improve upon their results by providing statistically and computationally more efficient algorithms.
 RL with General Function Approximation.
-There have also been some works studies RL with general function approximation.
+There have also been some works that studies RL with general function approximation.
 Russo and Roy (2013) and Osband and Roy (2014) initiated the study on the minimal structural assumptions that render sample-efficient learning by proposing a structural condition called Eluder dimension.
@@
 We focus on the single-parameter model, which are more relevant to our setting.
-The pioneering work by Cheung and Simchi-Levi (2017) achieved a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{K})$, where $K$ denotes the number of rounds in bandits.
+The pioneering work by Cheung and Simchi-Levi (2017) achieved a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{T})$, where $T$ denotes the number of rounds in bandits.
 This result was further enhanced by subsequent studies (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023).
-In particular, Périvier and Goyal (2022) significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ in the uniform revenue setting.
+In particular, Périvier and Goyal (2022) significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa T}+\kappa^{-1})$ in the uniform revenue setting.
 Most prior methods required storing all historical data and faced computational challenge.
 To address this issue, the most recent work by Lee and Oh (2024) proposed an algorithm with constant computational and storage costs building on recent advances in multiple-parameter model (Zhang and Sugiyama, 2023).
-Their algorithm achieves the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ and $\widetilde{\mathcal{O}}(d\sqrt{K}+\kappa^{-1})$ under uniform and non-uniform rewards respectively.
+Their algorithm achieves the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa T}+\kappa^{-1})$ and $\widetilde{\mathcal{O}}(d\sqrt{T}+\kappa^{-1})$ under uniform and non-uniform rewards respectively.
 However, although the underlying models of MNL bandits and MDPs share similarities, the challenges they present differ substantially, and techniques developed for MNL bandits cannot be directly applied to MNL MDPs.
@@
 Inhomogeneous, Episodic MDPs.
-An inhomogeneous, episodic MDP instance can be denoted by a tuple $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the length of each episode, $\mathbb{P}_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the transition kernel with $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)$ is being the probability of transferring to state $s^{\prime}$ from state $s$ and taking action $a$ at stage $h$, $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the deterministic reward function.
+An inhomogeneous, episodic MDP instance can be denoted by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the length of each episode, $\mathbb{P}_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the transition kernel with $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)$ is being the probability of transferring to state $s^{\prime}$ from state $s$ and taking action $a$ at stage $h$, $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the deterministic reward function.
 A policy $\pi=\{\pi_{h}\}_{h=1}^{H}$ is a collection of mapping $\pi_{h}$, where each $\pi_{h}:\mathcal{S}\to\Delta(\mathcal{A})$ is a function maps a state $s$ to distributions over $\mathcal{A}$ at stage $h$.
@@
 For any function $V:\mathcal{S}\to\mathbb{R}$, we define $[\mathbb{P}_{h}V](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot\mid s,a)}V(s^{\prime})$.
-The Bellman equation for policy $\pi$ and Bellman optimality equation are given by
-$Q_{h}^{\pi}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{\pi}\right](s,a),\quad Q_{h}^{*}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{*}\right](s,a).$
 Learning Protocol.
-In the online MDP setting, the learner interacts with the environment without the knowledge of the transition $\{\mathbb{P}_{h}\}_{h=1}^{H}$.
-As learning rewards is no harder than transitions, we assume the reward $\{r_{h}\}_{h=1}^{H}$ is known.
+In the online MDP setting, the learner interacts with the environment without the knowledge of the transition kernel $\{\mathbb{P}_{h}\}_{h=1}^{H}$.
+We assume the reward function $\{r_{h}\}_{h=1}^{H}$ is deterministic and known to the learner.
 The interaction proceeds in $K$ episodes.
@@
 Multinomial Logit (MNL) Mixture MDPs.
-Although significant advances have been achieved for MDPs with linear function approximation, Hwang and Oh (2023) show that there exists a set of features such that no linear transition model (including bilinear and low-rank MDPs) can induce a valid probability distribution over the state space, which limits the expressiveness of function approximation.
+Although significant advances have been achieved for MDPs with linear function approximation, Hwang and Oh (2023) show that there exists a set of features such that no linear transition model can induce a valid probability distribution, which limits the expressiveness of function approximation.
 To overcome this limitation, they propose a new class of MDPs with multinomial logit function approximation.
-We introduce the definition of MNL mixture MDPs below.
+However, their work focuses on the homogeneous setting, where the transitions remain the same across all stages (i.e., $\mathbb{P}_{1}=...=\mathbb{P}_{H}$).
+In this work, we address the more general inhomogeneous setting, allowing transitions to vary across different stages.
+We introduce the formal definition of inhomogeneous MNL mixture MDPs below.
 Definition 1 (Reachable States).
@@
 Furthermore, we define $S_{h,s,a}\triangleq|\mathcal{S}_{h,s,a}|$ and denote by $U\triangleq\max_{(h,s,a)}S_{h,s,a}$ the maximum number of reachable states.
-Remark 1.
-There are many cases that even when the state space is very large, the maximum number of reachable states $U$ is small.
-This phenomenon is common in situations where the next state is close to the current state.
-An illustrative example is the RiverSwim problem (Strehl and Littman, 2004).
 Definition 2 (MNL Mixture MDP).
-An MDP instance $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$ is called an inhomogeneous, episodic $B$-bounded MNL mixture MDP if there exist a known feature mapping $\phi(s^{\prime}\>|\>s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$ with $\lVert\phi(s^{\prime}\>|\>s,a)\rVert_{2}\leq 1$ and unknown vectors $\{\theta_{h}^{*}\}_{h=1}^{H}\in\Theta$ with $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime}\in\mathcal{S}_{h,s,a}$, it holds that
+An MDP instance $\mathcal{M}=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$ is called an inhomogeneous, episodic $B$-bounded MNL mixture MDP if there exist a known feature mapping $\phi(s^{\prime}\>|\>s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$ with $\lVert\phi(s^{\prime}\>|\>s,a)\rVert_{2}\leq 1$ and unknown vectors $\{\theta_{h}^{*}\}_{h=1}^{H}\in\Theta$ with $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime}\in\mathcal{S}_{h,s,a}$, it holds that
 $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta_{h}^{*})}{\sum_{\widetilde{s}\in\mathcal{S}_{h,s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta_{h}^{*})}.$
 Remark 1.
-While the work of Hwang and Oh (2023) focuses on the homogeneous setting, where the transitions remain the same across all stages (i.e., $\mathbb{P}_{1}=...=\mathbb{P}_{H}$), we address the more general inhomogeneous setting, allowing transitions to vary across different stages.
+This model is consistent with models like neural networks (LeCun et al., 2015), where the feature $\phi$ is obtained by omitting the final layer, and $\theta_{h}^{*}$ represents the weights of the last layer.
+A final softmax layer is then applied to ensure that the output forms a valid probability distribution.
 For any $\theta\in\mathbb{R}^{d}$, we define the induced transition as $p_{s,a}^{s^{\prime}}(\theta)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta)}{\sum_{\widetilde{s}\in\mathcal{S}_{s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta)}$.
-We introduce the key problem-dependent quantities for this problem firstly introduced by Hwang and Oh (2023).
+We then introduce the following two key problem-dependent quantities $\kappa$ and $\kappa^{*}$ that measure the effective non-linearity over the entire parameter space and around the ground truth parameter respectively.
 Assumption 1.
-There exists $10<\kappa\leq\kappa^{*}<$ 1 such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime},s^{\prime\prime}\in\mathcal{S}_{h,s,a}$, it holds that $\inf_{\theta\in\Theta}p_{s,a}^{s^{\prime}}(\theta)p_{s,a}^{s^{\prime\prime}}(\theta)\geq\kappa$ and $p_{s,a}^{s^{\prime}}(\theta^{*})p_{s,a}^{s^{\prime\prime}}(\theta^{*})>\kappa^{*}$.
-Assumption 1 is similar to the standard assumption in generalized linear model literature (Filippi et al., 2010; Wang et al., 2021) and logistic bandits literature (Faury et al., 2020; Abeille et al., 2021; Zhang and Sugiyama, 2023) to guarantee the Hessian matrix is non-singular in Property 1.
-4 Computationally Efficient Algorithm by Online Learning
-In this part, we design an algorithm that achieves $\mathcal{O}(1)$ computational and storage costs per episode.
+There exists $10<\kappa\leq\kappa^{*}<$ 1 such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime},s^{\prime\prime}\in\mathcal{S}_{h,s,a}$, it holds that $\inf_{\theta\in\Theta}p_{s,a}^{s^{\prime}}(\theta)p_{s,a}^{s^{\prime\prime}}(\theta)\geq\kappa$ and $p_{s,a}^{s^{\prime}}(\theta_{h}^{*})p_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\geq\kappa^{*}$.
+Assumption 1 is similar to the assumption in generalized linear bandit (Filippi et al., 2010) and logistic bandit (Faury et al., 2020; Abeille et al., 2021) to guarantee the Hessian matrix is non-singular.
+Finally, we show the claim about the range of the magnitude of $\kappa$ and $\kappa^{*}$.
+Claim 1.
+It holds that ${1}/{(U\exp(2B))^{2}}\leq\kappa\leq\kappa^{*}\leq{1}/{U^{2}}$.
+4 Statistically Efficient Algorithm
+The work of Hwang and Oh (2023) first introduced the MNL mixture MDPs and proposed an algorithm with a regret bound of $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$.
+However, as discussed in Claim 1, it follows that $U^{2}\leq\kappa^{-1}\leq(U\exp(2B))^{2}$, which results in the regret bound scaling polynomially with the number of reachable states $U$.
+In the worst case, $U$ can be equal to the size of the state space $S$, thereby undermining the motivation for function approximation, which aims to mitigate the dependence on the large state and action spaces.
+In this section, we address this significant issue by proposing a statistically efficient algorithm that eliminates this dependence in the dominant term of the regret.
+4.1 Parameter Estimation
+In this section, we first present the parameter estimation method based on the maximum likelihood estimation (MLE) for MNL mixture MDPs.
+Next, we review the confidence set construction based on the estimated parameters from previous work (Hwang and Oh, 2023).
+Finally, we propose our new confidence set construction and highlight the improvements it offers over the previous approach.
 Since the transition parameter $\theta_{h}^{*}$ is unknown, we need to estimate it using the historical data.
-At episode $k$, we collect a trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$, then define the variable: $y_{k,h}\in\{0,1\}^{N_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\operatorname{\mathds{1}}{(s_{k,h+1}=s^{\prime})}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $N_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$.
+At episode $k$, we collect a trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$, then define the variable: $y_{k,h}\in\{0,1\}^{S_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\mathds{1}_{\{s^{\prime}=s_{k,h+1}\}}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $S_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$.
 We denote by $p_{k,h}^{s^{\prime}}(\theta)=p_{s_{k,h},a_{k,h}}^{s^{\prime}}(\theta)$.
 Then $y_{k,h}$ is a sample from the following multinomial distribution:
-$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{N_{k,h}}}(\theta^{*})]),$
+$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{S_{k,h}}}(\theta^{*})]),$
 where the parameter $1$ indicates that $y_{k,h}$ is a single-trial sample.
 Furthermore, we define the noise $\epsilon_{k,h}^{s^{\prime}}=y_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\theta_{h}^{*})$.
-It is clear that $\epsilon_{k,h}\in[-1,1]^{N_{k,h}}$, $\mathbb{E}[\epsilon_{k,h}]=\mathbf{0}$ and $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\epsilon_{i,h}^{s^{\prime}}=0$.
-Efficiency Concern.  Hwang and Oh (2023) made the first step by proposing an approach using the maximum likelihood estimation (MLE). Specifically, their estimator $\theta_{k,h}$ is defined as
-$\theta_{k,h}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}-y_{i,h}^{s^{\prime}}\log p_{i,h}^{s^{\prime}}(\theta)+\frac{\lambda}{2}\lVert\theta\rVert_{2}^{2}.$ (1)
-Despite favorable theoretical guarantee, the computational and storage cost of this method is expensive.
-First, in episode $k$, the estimator $\theta_{k,h}$ in (1) is computed using all samples collected in previous episodes, resulting an $\mathcal{O}(k)$ storage cost.
-Second, to solve this optimization problem, the projected gradient descent (Boyd and Vandenberghe, 2004) are usually applied.
-However, as discussed in Faury et al.
-(2022), the optimization of the MLE problem typically require $\mathcal{O}(k\log(1/\epsilon))$ iterations to achieve an $\epsilon$-accurate solution, which is computationally expensive.
-To address this issue, we introduce a novel algorithm named UCRL-MNL with Online Learning (UCRL-MNL-OL), which achieves both the same regret with $\mathcal{O}(1)$ computational and storage cost.
-At a high level, our algorithm can be divided into two phases: (i) efficient online estimation, and (ii) efficient optimistic value function construction.
-We introduce the details in the following.
-Efficient Online Estimation.
-Instead of using all historical data, our algorithm updates the estimator in an online manner.
-Specifically, inspired by the works (Hazan et al., 2014; Zhang et al., 2016) on logistic bandit, we find the negative log-likelihood function is exponentially concave, which motivates us to apply a variant of online Newton step (Zhang et al., 2016; Oh and Iyengar, 2021).
-First, we define per-episodic loss function $f_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
-$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}},$
-$H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$ (2)
-We introduce the following key property of the loss function $f_{k,h}(\theta)$ by a second-order Taylor expansion and mean value theorem, whose proof can be found in Lemma 5 in Appendix A.
-Property 1.
-(i). For any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, $\exists~{}\bar{\theta}=\nu\theta_{1}+(1-\nu)\theta_{2}$ such that
-$f_{k,h}(\theta_{2})=f_{k,h}(\theta_{1})+\langle g_{k,h}(\theta_{1}),\theta_{2}-\theta_{1}\rangle+\frac{1}{2}\lVert\theta_{2}-\theta_{1}\rVert_{H_{k,h}(\bar{\theta})}^{2}.$
-(ii). For any $\theta\in\Theta$, it holds that
-$H_{k,h}(\theta)\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\triangleq\kappa W_{k,h}.$ (3)
-Property 1 implies that $f_{k,h}(\theta)$ is exponentially concave, which enables us to apply the online Newton step to construct an efficient estimator.
-Specifically, we construct $\widehat{\theta}_{k+1,h}$ by solving the problem:
-$\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2},$ (4)
-where $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ is the feature covariance matrix.
-We show the online estimator $\widehat{\theta}_{k,h}$ in (4) enjoys computational and storage efficiency simultaneously.
-As the optimization problem exhibits a standard online mirror descent formulation, it can be solved with a projected gradient step with the following equivalent formulation by
-$\theta^{\prime}_{k+1,h}=\widehat{\theta}_{k,h}-\widehat{\Sigma}_{k+1,h}^{-1}g_{k,h}(\widehat{\theta}_{k,h}),\quad\mbox{and}\quad\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widehat{\Sigma}_{k+1,h}}.$
-This update enjoys a computational cost of only $\mathcal{O}(d^{3}U)$, independent of episode count $k$ (Zhang and Sugiyama, 2023; Lee and Oh, 2024).
-As for storage costs, it avoids the need to store all historical data by updating the feature covariance matrix $\widehat{\Sigma}_{k,h}$ incrementally, requiring only $\mathcal{O}(d^{2})$ storage cost.
-We note that a dependence on $U$ is introduced in the computational cost.
-However, as we have discussed in Remark 1, $U$ can be much smaller than the state space size $S$, which is acceptable.
-This dependency is typical for model-based methods that directly learn transitions, as it need to control the estimation error of transitions, which typically involves a total of $U$ elements.
-Similar dependencies have been observed in the literature (Hwang and Oh, 2023).
-Additionally, even for the model-free method (Yang and Wang, 2020; Zhou et al., 2021) which learn value functions, a common assumption is the value $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}V(s^{\prime})$ can be obtained by an Oracle, which depends on $U$ implicitly.
-Then, we show the estimator $\widehat{\theta}_{k,h}$ is close to the true parameter $\theta_{h}^{*}$ by the following lemma.
+It is clear that $\epsilon_{k,h}\in[-1,1]^{S_{k,h}}$, $\mathbb{E}[\epsilon_{k,h}]=\mathbf{0}$ and $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\epsilon_{i,h}^{s^{\prime}}=0$.
+We estimate the parameter $\theta_{h}^{*}$ using the MLE and construct the estimator $\widehat{\theta}_{k,h}$ as follows:
+$\widehat{\theta}_{k,h}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\mathcal{L}_{k,h}(\theta)\triangleq\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}-y_{i,h}^{s^{\prime}}\log p_{i,h}^{s^{\prime}}(\theta)+\frac{\lambda_{k}}{2}\lVert\theta\rVert_{2}^{2}.$ (1)
+where $\lambda_{k}$ is the regularization parameter.
+Though the MLE estimator $\widehat{\theta}_{k,h}$ is the same as that of Hwang and Oh (2023), the confidence set is constructed differently.
+Specifically, define the gradient $\mathcal{G}_{k,h}(\theta)$ and Hessian matrix $\mathcal{H}_{k,h}(\theta)$ of the MLE loss by
+$\mathcal{G}_{k,h}(\theta)$ $=\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta)-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}+\lambda_{k}\theta,$
+$\mathcal{H}_{k,h}(\theta)$ $=\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}-\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)p_{i,h}^{s^{\prime\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime\prime}})^{\top}+\lambda_{k}I.$
+Furthermore, we define the feature covariance matrix $A_{k,h}=\kappa^{-1}\lambda_{k}I+\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}$.
+By demonstrating $\mathcal{H}_{k,h}(\theta)\succeq\kappa A_{k,h},\forall\theta\in\Theta$, Hwang and Oh (2023) construct the confidence set as
+$\mathcal{C}_{k,h}=\Big{\{}\big{\lVert}\theta-\widehat{\theta}_{k,h}\big{\rVert}_{A_{k,h}}\leq\kappa^{-1}\sqrt{d\log(kH/\delta)}\triangleq\beta_{k}\Big{\}}.$ (2)
+Since the radius of the confidence set depends on $\kappa^{-1}$, the final regret bound also exhibits a dependence on $\kappa^{-1}$.
+To eliminates this dependence, we construct a $\kappa$-independent confidence set based on new Bernstein-like inequalities in Lemma 13, following recent advances in logistic bandits (Faury et al., 2020; Périvier and Goyal, 2022).
+Specifically, we show the following lemma.
 Lemma 1.
-For any $\delta\in(0,1)$, define the confidence set as
-$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k,h}}\leq\sqrt{\lambda}B+\sqrt{\frac{8}{\kappa}d\log\left(1+\frac{kUH}{d\lambda\delta}\right)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$ (5)
+For any $\delta\in(0,1)$, set $\lambda_{k}=d\log(kH/\delta)$ and define the confidence set as
+$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\big{\lVert}\mathcal{G}_{k,h}(\theta)-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta)}\leq(B+3)\sqrt{d\log(kH/\delta)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$ (3)
 Then, we have $\Pr[\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}]\geq 1-\delta,\forall k\in[K],h\in[H]$.
+Comparison to prior work.  We compare the confidence sets defined in Eqs. (2) and (3).
+For the confidence set in (3), by the self-concordance property of log-loss in Lemma 10, we have:
+$\big{\lVert}\theta-\widehat{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}(\theta)}\leq(1+3\sqrt{2})\big{\lVert}\mathcal{G}_{k,h}(\theta)-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta)}\leq(1+3\sqrt{2})\widehat{\beta}_{k}.$
+Then, note that $\mathcal{H}_{k,h}(\theta)\succeq\kappa A_{k,h}$ for all $\theta\in\Theta$, we have
+$\big{\lVert}\theta-\widehat{\theta}_{k,h}\big{\rVert}_{A_{k,h}}\leq\kappa^{-1/2}\big{\lVert}\theta-\widehat{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}(\theta)}\leq\kappa^{-1/2}(1+3\sqrt{2})(B+3)\sqrt{d\log(kH/\delta)}.$ (4)
+Thus, compared to the confidence set in Eq.
+(2) from Hwang and Oh (2023), our confidence set in Eq.
+(3) provides a strict improvement by at least a factor of $\kappa^{-1/2}$.
+This improvement is one of the key components to eliminate the dependence on $\kappa^{-1}$ in the dominant term of the final regret bound.
+Additionally, we identify a technical issue of Hwang and Oh (2023).
+Specifically, they bound the confidence set in Eq.
+(2) using the self-normalized concentration in Lemma 12.
+However, the noise is not independent, and since $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\epsilon_{i,h}^{s^{\prime}}=0$ (due to the learner visiting each stage $h$ exactly once per episode), it does not satisfy the zero-mean sub-Gaussian condition in Lemma 12.
+We observe similar oversights in multinomial logit contextual bandits (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023), an issue that, to our knowledge, has not been explicitly addressed in prior work.
+This issue can be resolved with only slight modifications in constant factors by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al.
+(2024b), a simplified version of Lemma 13.
+4.2 Optimistic Value Function Construction
+Given the confidence set $\widehat{\mathcal{C}}_{k,h}$, it is natural to follow the principle of “optimism in the face of uncertainty” and construct the optimistic value function.
+Hwang and Oh (2023) constructed the optimistic value function $\bar{Q}_{k,h}(s,a)$ by adding a closed-form upper confidence bound as follows:
+$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+2H\beta_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{A_{k,h}^{-1}}\bigg{]}_{[0,H]},$ (5)
+where $\bar{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\bar{Q}_{k,h}(s,a)$.
+Then, a naive idea to compute the optimistic value function is replacing the radius of the confidence set $\beta_{k}$ with $\widehat{\beta}_{k}$ and the feature covariance matrix $A_{k,h}$ with the Hessian matrix $\mathcal{H}_{k,h}(\theta_{h}^{*})$.
+However, there are two issues with this approach.
+First, the true parameter $\theta_{h}^{*}$ is unknown thus the Hessian matrix $\mathcal{H}_{k,h}(\theta_{h}^{*})$ is not computable in the algorithmic updates.
+Second, though $\widehat{\beta}_{k}$ is independent of $\kappa$ and the Hessian matrix $\mathcal{H}_{k,h}(\theta_{h}^{*})$ captures local information, the bonus term $\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}(\theta)}$ remains in a global form.
+This term involves taking the maximum over all states $s^{\prime}\in\mathcal{S}_{h,s,a}$, which prevents fully utilizing the local information.
+To address these challenges, we construct the optimistic value function by directly taking the maximum expected reward over the confidence set.
+Specifically, we define $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
+$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\max_{\theta\in\widehat{\mathcal{C}}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})\bigg{]}_{[0,H]},\widehat{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s,a).$ (6)
+This construction addresses the first challenge by eliminating the need for the Hessian matrix $\mathcal{H}_{k,h}(\theta_{h}^{*})$ and directly leveraging the local information embedded in the confidence set $\widehat{\mathcal{C}}_{k,h}$.
+For the second challenge, although we bypass this issue in the construction of the optimistic value function, we still need to address it in the analysis.
+To tackle this, we employ a second-order Taylor expansion, in contrast to the first-order expansion used in the analysis of Hwang and Oh (2023).
+This allows for a more precise capture of local information.
+Further details are provided in Lemma 7 in the appendix.
+4.3 Regret Guarantee
+Algorithm 1 UCRL-MNL-LL
+0:  Regularization parameter $\lambda$, confidence width $\widehat{\beta}_{k}$, confidence parameter $\delta$.
+1:  Initialization: Set $\widehat{\theta}_{1,h}=\mathbf{0},\widehat{Q}_{1,h}(\cdot,\cdot)=0,\widehat{V}_{1,h}(\cdot)=0$ for all $h\in[H]$.
+2:  for $k=1,\ldots,K$ do
+3:     for $h=1,\ldots,H$ do
+4:        Observe current state $s_{k,h}$ and select action $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
+5:     end for
+6:     Set $\widehat{V}_{k+1,H+1}(\cdot)=0$.
+7:     for $h=H,\ldots,1$ do
+8:        Compute the estimator $\widehat{\theta}_{k+1,h}$ by Eq. (1) and update the confidence set $\widehat{\mathcal{C}}_{k+1,h}$ by Eq. (3).
+9:        Compute $\widehat{Q}_{k+1,h}(\cdot,\cdot)$ and $\widehat{V}_{k+1,h}(\cdot)$ as in Eq. (6).
+10:     end for
+11:  end for
+Based on the parameter estimation in Section 4.1 and the construction of the optimistic value function in Section 4.2, we propose the UCRL-MNL-LL algorithm.
+At each stage $h$ of episode $k$, the algorithm observes the current state $s_{k,h}$ and selects the action that maximizes the value function, i.e., $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$, and transits to next state $s_{k,h+1}$.
+After collecting the trajectory $\{s_{k,h},a_{k,h}\}_{h=1}^{H}$, the estimator $\widehat{\theta}_{k+1,h}$ is updated using Eq.
+(1), and the confidence set $\widehat{\mathcal{C}}_{k+1,h}$ is updated according to Eq.
+(3).
+Then, the value function $\widehat{Q}_{k+1,h}$ and $\widehat{V}_{k+1,h}$ are updated using Eq.
+(6).
+The detailed procedure is outlined in Algorithm 1.
+We show it achieves the following regret guarantee.
+Theorem 1.
+For any $\delta\in(0,1)$, set $\lambda_{k}=d\log(kH/\delta)$, and $\widehat{\beta}_{k}=(B+3)\sqrt{d\log(kH/\delta)}$.
+With probability at least $1-\delta$, UCRL-MNL-LL algorithm (Algorithm 1) ensures the following guarantee:
+${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2}\big{)}.$
 Remark 2.
-Notably, a similar confidence set is achieved by Hwang and Oh (2023) by using the MLE defined in (1).
-In contrast, we obtain the same results by using an online estimator which only suffers $\mathcal{O}(1)$ computation and storage cost per round.
-Besides, we identify a technical issue in their analysis.
-Specifically, they bound the confidence set using the self-normalized concentration for vector-valued martingales in Lemma 15.
-However, it is crucial to recognize that the noise is not independent and satisfies $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\epsilon_{i,h}^{s^{\prime}}=0$ since the learner visits each stage $h$ exactly once per episode.
-Thus, the noise $\epsilon_{i,h}^{s^{\prime}}$ becomes deterministic and non-zero given the remaining noise $\epsilon_{i,h}^{s^{\prime\prime}}$ for $s^{\prime\prime}\neq s^{\prime}$.
-This contravenes the zero-mean sub-Gaussian condition in Lemma 15.
-We observe similar oversights also appear in the works on multinomial logit contextual bandits (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023).
-To our knowledge, this issue has not been explicitly identified in previous studies.
-We note that this issue can be resolved by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al.
-(2024b) with only slight modifications in constant factors.
-Algorithm 1 UCRL-MNL-OL
-0:  Regularization parameter $\lambda$, confidence width $\widehat{\beta}_{k}$, confidence parameter $\delta$.
-1:  Initialization: $\widehat{\Sigma}_{1,h}=\lambda I,\widehat{\theta}_{1,h}=\mathbf{0}$ for all $h\in[H]$.
-2:  for $k=1,\ldots,K$ do
-3:     Compute $\widehat{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (7).
-4:     for $h=1,\ldots,H$ do
-5:         Observe current state $s_{k,h}$, select action $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
-6:         Set $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$.
-7:         Compute $\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$.
-8:     end for
-9:  end for
-Efficient Optimistic Value Function Construction.
-Given the confidence set $\widehat{\mathcal{C}}_{k,h}$, the most direct way is to adopt the principle of “optimism in the face of uncertainty” and construct the optimistic value function as the maximum expected reward over the confidence set.
-Specifically, we define the optimistic value function $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
-$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\max_{\theta\in\widehat{\mathcal{C}}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})\bigg{]}_{[0,H]},\widehat{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s,a).$ (6)
-However, this construction is not efficient as it requires solving a maximization problem over the confidence set.
-Actually, the algorithm only requires the estimated action-value function to be an upper bound for the true value functions.
-Thus, we can use a closed-form confidence bound instead of computing the maximal value over the confidence set.
-To this end, we present the following lemma that enables us to construct the optimistic value function more efficiently.
+Focusing on the dominant term, our guarantee eliminates the problematic dependence on $\kappa^{-1}$, in stark contrast to the $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ result of Hwang and Oh (2023).
+As noted in Claim 1, such undesirable dependence has polynomial scaling with the number of reachable states $U$, which can be as large as the entire state space $S$ in the worst case.
+This renders the guarantee for function approximation—designed for settings with large state and action spaces—essentially vacuous.
+5 Computationally Efficient Algorithm
+While the UCRL-MNL-LL algorithm is the first statistically efficient algorithm for MNL mixture MDPs, it is computationally expensive dur the the optimization of the MLE in Eq.
+(1) and non-convex optimization in Eq.
+(6).
+To address these challenges, we propose a computationally efficient algorithm in this section, which attains the same regret but with constant computational costs per episode.
+5.1 Efficient Online Parameter Estimation
+In this section, we focus on estimating the unknown parameter $\theta_{h}^{*}$ in a computationally efficient manner.
+We first discuss the storage and time complexities of the MLE optimization in Eq.
+(1).
+Next, we introduce an efficient online parameter estimation based on online mirror descent that provides similar guarantees to the MLE, but with constant storage and time complexity per episode.
+For the storage complexity, the optimization problem defined in Eq.
+(1) requires storing all historical data, resulting in a storage complexity of $\mathcal{O}(k)$ at episode $k$.
+In terms of time complexity, the problem does not have a closed-form solution and can only be solved to within an $\varepsilon$-accuracy, such as using projected gradient descent.
+As discussed in Faury et al.
+(2022), optimizing the MLE typically requires $\mathcal{O}(\log(1/\varepsilon))$ iterations to achieve an $\varepsilon$-accurate solution.
+Since the loss function is defined over all historical data, each gradient step incurs a query complexity of $\mathcal{O}(k)$.
+As $\varepsilon$ is usually chosen as $1/k$ for episode $k$, the total time complexity is $\mathcal{O}(k\log k)$ at episode $k$.
+Consequently, both storage and time complexities scale linearly with the episode count, which is computationally expensive.
+To improve the computational efficiency, the basic idea is to estimate the unknown parameter with the online mirror descent (OMD) update instead of the MLE as defined in Eq.
+(1).
+To this end, we first define per-episode loss function $\ell_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
+$\ell_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\nabla\ell_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$ (7)
+$H_{k,h}(\theta)$ $=\nabla^{2}\ell_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.\vspace{-1mm}$
+Then, the design of the OMD algorithm can be conceptually divided into two parts: the approximation of the past losses and the approximation of the current loss. We provide the details of each below.
+Approximate the past losses.
+To integrate historical information from previous iterations while avoiding the use of MLE in Eq.
+(1), we construct the estimator $\bar{\theta}_{k+1,h}$ using the implicit OMD form:
+$\bar{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\ell_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\bar{\theta}_{k,h}\rVert_{\bar{\mathcal{H}}_{k,h}}^{2}\Big{\}},$ (8)
+where $\eta$ is a step size and ¯$\bar{\mathcal{H}}_{k,h}\triangleq\bar{\mathcal{H}}_{k,h}(\bar{\theta}_{k+1,h})=\sum_{i=1}^{k-1}H_{i,h}(\bar{\theta}_{i+1,h})+\lambda_{k}I$.
+The optimization problem can be decomposed in two terms.
+The first term is the instantaneous log-loss $\ell_{k,h}(\theta)$, which accounts for the information of the current episode.
+The second is a regularization term that ensures the current model remains close to the previous one, $\bar{\theta}_{k,h}$, thereby incorporating the historical information acquired so far.
+The most critical aspect in the above is the design of the local norm ¯$\bar{\mathcal{H}}_{k,h}$, which intentionally approximate the per-episode Hessian matrix by $H_{i}(\bar{\theta}_{i+1,h})$ at a look ahead point $\bar{\theta}_{i+1,h}$.
+Such a Hessian matrix, originally introduced by Faury et al.
+(2022), effectively captures the local curvature of the loss function and is crucial for ensuring statistical efficiency.
+The update rule in Eq.
+(8) is storage efficient, as it only requires storing the Hessian matrix ¯$\bar{\mathcal{H}}_{k,h}$, which can be updated incrementally, resulting in an $\mathcal{O}(1)$ storage cost.
+In terms of time complexity, the optimization problem in Eq.
+(8) suffers an $\mathcal{O}(\log k)$ time complexity at episode $k$, since the loss function is defined only over the current episode.
+While this represents a significant improvement over the $\mathcal{O}(k\log k)$ time complexity of the MLE in Eq.
+(1), there is still a need to reduce the cost further to $\mathcal{O}(1)$ per episode, particularly given the potentially large number of episodes.
+Approximate the current loss.
+To achieve $\mathcal{O}(1)$ time complexity per episode, we can further approximate the current loss with a second order approximation.
+Drawing inspiration from Zhang and Sugiyama (2023), we define the second-order approximation of the original loss function $\ell_{k,h}(\theta)$ at $\widetilde{\theta}_{k,h}$ as ~$\widetilde{\ell}_{k,h}(\theta)=\ell_{k,h}(\widetilde{\theta}_{k,h})+\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{H_{k,h}(\widetilde{\theta}_{k,h})}^{2}$, where $\widetilde{\theta}_{k,h}$ is the current estimate.
+Then, we can replace $\ell_{k,h}(\theta)$ with its second-order approximation ~$\widetilde{\ell}_{k,h}(\theta)$ in the optimization problem in Eq.
+(8).
+This leads to the following approximate optimization problem:
+$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}\Big{\}}.$ (9)
+where $\eta$ is the step size, ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$ and $\mathcal{H}_{k,h}=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda_{k}I$.
+Then, Eq.
+(9) can be solved with a single projected gradient step with the following equivalent formulation:
+$\widetilde{\theta}^{\prime}_{k+1,h}=\widetilde{\theta}_{k,h}-\eta\widetilde{\mathcal{H}}_{k,h}^{-1}\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\widetilde{\theta}^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}.$
+Thus, Eq. (9) is computationally efficient, as it only suffers an $\mathcal{O}(1)$ storage and time complexity.
+Notice that the update rule in Eq. (9) is actually a standard online mirror descent (OMD) formula,
+$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta\rangle+\frac{1}{\eta}\mathcal{D}_{\psi_{k}}(\theta,\widetilde{\theta}_{k,h})\Big{\}}.$ (10)
+where the regularizer is $\psi_{k}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}$ and $\mathcal{D}_{\psi_{k}}(\cdot,\cdot)$ is the induced Bregman divergence.
+Therefore, we can construct the confidence set building upon the modern analysis of OMD (Orabona, 2019; Zhao et al., 2024).
+Specifically, we can construct the $\kappa$-independent confidence set as follows.
 Lemma 2.
-Suppose Lemma 1 holds. For any $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds
-$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}.$
-Based on this lemma, we can replace the maximization problem in (6) with closed-form confidence bound and construct the optimistic value function $\widehat{Q}_{k,h}(s,a)$ as
-$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\widehat{\Sigma}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (7)
-At state $s_{k,h}$, our algorithm chooses $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
-Algorithm 1 presents the detailed procedure.
-We show that our algorithm achieves the following regret guarantee.
-Theorem 1.
-For any $\delta\in(0,1)$, set $\widehat{\beta}_{k}$ as in Lemma 1 and $\lambda=1$, with probability at least $1-\delta$, Algorithm 1 ensures the following regret guarantee
-${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}\big{(}\kappa^{-1}dH^{2}\sqrt{K}\big{)}.$
-Remark 4.
-Our algorithm attains the same regret of Hwang and Oh (2023), yet for the more general inhomogeneous MDPs.
-Importantly, our algorithm only requires constant computational and storage costs, matching the computational efficiency of the linear cases (Jin et al., 2020; Zhou et al., 2021).
-5 Statistically Improved Algorithm by Local Learning
-In this section, we present an enhanced algorithm, UCRL-MNL-LL, that leverages local information to improve the statistical efficiency for MDPs with MNL function approximation.
-While the UCRL-MNL-OL algorithm in Section 4 offers favorable computational and storage efficiency, it suffers from a dependence on $\kappa^{-1}$ in the regret.
-By Assumption 1, the quantity $\kappa$ may be exponentially small in the worst case, exhibiting an exponential dependence on the radius of the parameter set and linear in the number of reachable states, i.e., $\kappa^{-1}=\mathcal{O}((U\exp(B))^{2})$.
-This creates a significant gap between MNL and linear function approximation.
-Recently, the improved dependence on $\kappa$ has been achieved in the logistic bandit literature (Faury et al., 2020; Abeille et al., 2021; Périvier and Goyal, 2022) by the use of generalization of the Bernstein-like tail inequality (Faury et al., 2020) and the self-concordant-like property of the log loss (Bach, 2010).
-Thus, a natural question then arises: Can we improve the statistical efficiency of MNL function approximation for MDPs?
-We answer this question affirmatively by proposing an enhanced algorithm that reduces the dependence on $\kappa$ significantly through the use of local information.
-Though the achievement has been made in the logistic bandit, the extension to MDPs with MNL function approximation is non-trivial and new techniques specific to MDPs are required.
-We first analyze where the dependence on $\kappa$ comes from. By the analysis of Theorem 1, we can see that the regret of the algorithm can be upper-bounded as follows,
-$\operatorname*{Reg}(K)\leq 2H\widehat{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}.$ (8)
-Here, the first term corresponds to the overestimation of the value function, and the second term represents the martingale difference sequence arising from the stochastic transition dynamics.
-The second term can be bounded using the Azuma-Hoeffding inequality, which is independent of $\kappa$.
-For the first term, by the construction of the confidence set in Lemma 1, $\widehat{\beta}_{k}$ represents the width of the confidence set, contributing to a $\kappa^{-1/2}$ dependence.
-The feature covariance matrix $\widehat{\Sigma}_{k,h}$ defined in (4) reflects the degree of exploration across different states, which also results in a $\kappa^{-1/2}$ dependence.
-To mitigate these dependencies, we design a local learning algorithm that: (i) constructs a confidence set independent of $\kappa$; and (ii) builds a $\kappa$-independent feature covariance matrix that effectively captures the exploration degree across different states.
-We introduce the details in the following.
-5.1 Improved Online Estimation
-In Property 1, we show the Hessian matrix $H_{k,h}$ is lower bounded by a positive definite matrix $\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ uniformly over the parameter domain, as in (3).
-This quantity measures the exploration degree across different states and is used to update the parameter estimation and construct the confidence set.
-However, the bound is not tight, as the Hessian matrix can be significantly larger in certain regions, away from the global minimum.
-This observation motivates us to design a local learning algorithm that can adaptively leverage local information for improved guarantee.
-Inspired by recent advances in multinomial logistic bandit (MLogB) (Zhang and Sugiyama, 2023) and multinomial logit contextual bandit (MNL) (Lee and Oh, 2024), we run an online mirror descent algorithm to estimate the parameter $\theta_{h}^{*}$.
-Differently from (4), we use the local Hessian matrix ~$\widetilde{\mathcal{H}}_{k,h}$ to update the estimation instead of the global lower bound.
-Specifically, we estimate $\widetilde{\theta}_{k,h}$ as follows:
-$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$ (9)
-where $\eta$ is the step size, $\mathcal{H}_{k,h}=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d}$ and ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$.
-Note that both $\mathcal{H}_{k,h}$ and ~$\widetilde{\mathcal{H}}_{k,h}$ can be updated incrementally.
-Similar to the update in (4), the optimization problem in (9) can be efficiently solved using the following two-step update:
-$\theta^{\prime}_{k+1,h}=\theta_{k,h}-\eta\widetilde{\mathcal{H}}_{k,h}^{-1}g_{k,h}(\theta_{k,h}),\quad\mbox{and}\quad\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}.$
-This two-step update procedure incurs a computational cost of $\mathcal{O}(d^{3}U)$ and a storage cost of $\mathcal{O}(d^{2})$ per episode, both independent of the episode count $k$.
-Based on this estimator, we can construct the $\kappa$-independent confidence set as follows.
-Lemma 3.
-For any $\delta\in(0,1)$, define the confidence set as
+For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$ and $\lambda=84\sqrt{2}\eta(B+d)$, define
 $\widetilde{\mathcal{C}}_{k,h}=\big{\{}\theta\in\Theta\mid\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}\leq\widetilde{\beta}_{k}\big{\}},$
@@
 Remark 3.
-Zhang and Sugiyama (2023) studied the multiple-parameter MLogB model, where the unknown parameter is a matrix.
-Consequently, the confidence set in Theorem 3 of their work exhibits a polynomial dependence on the number of possible outcomes, which corresponds to the number of reachable states $U$ in our setting.
-This dependence is acceptable in the bandit setting while is not suitable for the MDP setting.
-In contrast, Lee and Oh (2024) focused on the single-parameter MNL model and revisited the self-concordant-like property, demonstrating that the log-loss of the single-parameter MNL model is $3\sqrt{2}$-self-concordant-like (Proposition B.1 in Lee and Oh (2024)).
-This property is crucial for the improved confidence set in Lemma 3 that is independent of $\kappa$ and $U$.
-Algorithm 2 UCRL-MNL-LL
+Compared to the confidence set in Lemma 1, the radius $\widetilde{\beta}_{k}$ in Lemma 2 includes an additional $\log U$ factor.
+This is due to our approximation of the original MLE using the OMD update.
+5.2 Efficient Optimistic Value Function Construction
+Although the optimistic value function in Eq.
+(6) preserves local information effectively and provides strong theoretical guarantees, it is computationally intractable due to the need to solve a non-convex optimization problem.
+To address this challenge, we propose an efficient method in this section.
+The key idea is to use a second-order Taylor expansion to derive a closed-form bonus term, which replaces the operation of taking the maximum over the non-convex confidence set.
+While this idea has been used in bandit settings, fundamental challenges arise when applying it in the MDP setting.
+Specifically, Zhang and Sugiyama (2023) studied the multi-parameter MLogB bandit, where each outcome is associated with a distinct parameter vector.
+In contrast, MNL mixture MDPs involve a single shared parameter vector across all outcomes.
+This distinction leads to a more complex Hessian matrix, necessitating a more sophisticated analysis.
+A direct use of their analysis will leads to a polynomial dependence on the number of reachable states $U$, which is undesirable in the MDP setting.
+Lee and Oh (2024) focused on the single-parameter MNL bandit, which is more closely related to our setting.
+However, they construct the optimistic value function by directly taking the maximum over the confidence set, a computationally intractable approach in the MDP setting.
+As a result, they can apply a second-order Taylor expansion around the ground truth parameter $\theta_{h}^{*}$ in their analysis, while we must apply it around the estimated parameter $\widetilde{\theta}_{k,h}$ to construct the bonus term explicitly.
+For MDPs, we show the value difference arising from the transition estimation error as follows.
+Lemma 3.
+Suppose Lemma 2 holds. For any $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds
+$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}.$
+where
+$\epsilon_{s,a}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\epsilon_{s,a}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
+Based on Lemma 3, we construct the optimistic value function as follows:
+$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]},$ (11)
+where $\widetilde{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\widetilde{Q}_{k,h}(s,a)$.
+In contrast to the value function in Eq.
+(5), which incorporates the term $\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$, the refined value function in Eq.
+(11) replaces it with $\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}$.
+This modification better preserves local information, offering a more accurate estimation error bound.
+5.3 Regret Guarantee
+Algorithm 2 UCRL-MNL-OL
 0:  Step size $\eta$, regularization parameter $\lambda$, confidence width $\widetilde{\beta}_{k}$, confidence parameter $\delta$.
@@
 2:  for $k=1,\ldots,K$ do
-3:     Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in (11).
+3:     Compute $\widetilde{Q}_{k,h}(\cdot,\cdot)$ in a backward way as in Eq. (11).
 4:     for $h=1,\ldots,H$ do
@@
 6:        Update ~$\widetilde{\mathcal{H}}_{k,h}=\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$.
-7:         Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}$.
-8:         Update $\mathcal{H}_{k,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
+7:        Compute $\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}$.
+8:        Update $\mathcal{H}_{k+1,h}=\mathcal{H}_{k,h}+H_{k,h}(\widetilde{\theta}_{k+1,h})$.
 9:     end for
 10:  end for
-5.2 Improved Optimistic Value Function Construction
-Based on the confidence set in Lemma 3, a natural choice for the optimistic value function construction is analogous to (7) and can be expressed as:
-$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\mathcal{H}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (10)
-However, though $\widetilde{\beta}_{K}$ now is independent of $\kappa$ and $\mathcal{H}_{k,h}^{-1}$ preserves the local information in (10), the norm $\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is still in a global manner due to the maximum operation, leading to a $\kappa^{-1/2}$ dependence.
-To address this issue, we propose a new construction of the optimistic value function.
-Specifically, we employ a second-order Taylor expansion to more accurately bound the value difference arising from transition estimation errors.
-Although the idea of using second-order Taylor expansion has been explored in bandits (Périvier and Goyal, 2022; Zhang and Sugiyama, 2023; Lee and Oh, 2024), fundamental differences arise in the MDP setting.
-Specifically, Périvier and Goyal (2022) studied the uniform revenue setting, where the reward is identical for all actions.
-Zhang and Sugiyama (2023) and Lee and Oh (2024) focused on the non-uniform setting, but the rewards for different actions are known a priori.
-However, in the MDP setting, the value function is state-dependent and unknown to the learner, leading to a more challenging problem.
-Moreover, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize the expected revenue, whereas in MNL MDPs, the goal is to choose one action at each stage to maximize the cumulative reward.
-Due to these differences, the analyses used in the bandit setting cannot be directly applied to the MDP setting.
-For MDPs, we show the value difference arising from the transition estimation error as follows.
-Lemma 4.
-Suppose Lemma 3 holds. For any $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds
-$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}.$
-where
-$\epsilon_{s,a}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\epsilon_{s,a}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
-Based on Lemma 4, we construct the optimistic value function as follows:
-$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]}.$ (11)
-In contrast to the value function specified in (10), where the term $\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is utilized, the refined value function introduced in (11) substitutes this term with $\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}$.
-This adjustment better preserves local information, offering a more precise and $\kappa$-independent estimation error bound.
-At state $s_{k,h}$, our algorithm chooses action $a_{k,h}=\max\widetilde{Q}_{k,h}(s_{k,h},\cdot)$.
+The overall algorithm UCRL-MNL-OL is similar to UCRL-MNL-LL, but with the estimator and optimistic value function updated in a computationally efficient manner.
 The detailed algorithm is presented in Algorithm 2.
-We show the regret guarantee of UCRL-MNL-LL in the following theorem.
+We provide the guarantee of UCRL-MNL-OL in the following theorem.
 Theorem 2.
-For any $\delta\in(0,1)$, set $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, $\eta=\frac{1}{2}\log(1+U)+(B+1)$ and $\lambda=84\sqrt{2}\eta(B+d)$, with probability at least $1-\delta$, UCRL-MNL-LL algorithm ensures the following regret guarantee
+For any $\delta\in(0,1)$, set $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(kH/\delta))$, $\eta=\frac{1}{2}\log(1+U)+(B+1)$ and $\lambda=84\sqrt{2}\eta(B+d)$, with probability at least $1-\delta$, UCRL-MNL-LL (Algorithm 2) ensures
 ${\operatorname*{Reg}}(K)\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2}\big{)}.$
 Remark 4.
-The high-order term in Theorem 2 is now independent $\kappa$, significantly improving the statistical efficiency compared with Theorem 1.
-In comparison with the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear cases (Zhou et al., 2021; He et al., 2023), the higher-order term in Theorem 2 only differs by a factor of $H^{1/2}$, almost matching the same computational and statistical efficiency simultaneously.
+UCRL-MNL-OL attains the same regret as UCRL-MNL-LL, but with constant computational cost per episode.
+This is achieved by constructing an efficient online estimation based on OMD and an optimistic value function by closed-form bonus instead of the non-convex optimization.
 6 Lower Bound
-In this section, we establish a lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and the logistic bandit problem.
-Consider the following logistic bandit problem (Faury et al., 2020; Abeille et al., 2021): at each round $k\in[K]$, the learner selects an action $x_{k}\in\mathcal{X}$ and receives a reward $r_{k}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter.
-The learner aims to to minimize the regret:
-${\operatorname*{Reg}}^{\mathtt{LogB}}(K)=\max_{x\in\mathcal{X}}\sum_{k=1}^{K}\mu(x^{\top}\theta^{*})-\sum_{k=1}^{K}\mu(x_{k}^{\top}\theta^{*}).$
-Theorem 3 (Lower Bound).
+In this section, we establish the lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and the logistic bandit problem.
+Consider the following logistic bandit problem (Faury et al., 2020): at each round $t\in[T]$, the learner selects an action $x_{t}\in\mathcal{X}$ and receives a reward $r_{t}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter.
+The learner aims to to minimize the regret: ${\operatorname*{Reg}}^{\mathtt{LogB}}(T)=\max_{x\in\mathcal{X}}\sum_{t=1}^{T}\mu(x^{\top}\theta^{*})-\sum_{t=1}^{T}\mu(x_{t}^{\top}\theta^{*})$.
+Theorem 3.
 For any logistic bandit problem $\mathcal{B}$, there exists an MNL mixture MDP $\mathcal{M}$ such that learning $\mathcal{M}$ is as hard as learning $H/2$ independent instances of $\mathcal{B}$ simultaneously.
-Corollary 1.
+Corollary 1 (Lower Bound).
 For any problem instance $\{\theta_{h}^{*}\}_{h=1}^{H}$ and for $K\geq d^{2}\kappa^{*}$, there exists an MNL mixture MDP with infinite action space such that $\operatorname*{Reg}(K)\geq\Omega(dH\sqrt{K\kappa^{*}})$.
 Remark 5.
-Corollary 1 can be proved by combining Theorem 3 and the $\Omega(d\sqrt{K\kappa^{*}})$ lower bound for logistic bandits with infinite arms by Abeille et al.
+Corollary 1 also implies a problem-independent lower bound of $\Omega(dH\sqrt{K})$ directly.
+Corollary 1 can be proved by combining Theorem 3 and the $\Omega(d\sqrt{T\kappa^{*}})$ lower bound for logistic bandits with infinite arms by Abeille et al.
 (2021).
@@
 However, after the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al.
-(2024) proposed a new reduction that bridges MNL mixture MDPs with linear mixture MDPs by approximating multinomial logit functions to linear functions, employing the mean value theorem.
-Leveraging this new reduction, they established an $\Omega(dH^{3/2}\sqrt{K})$ lower bound for MNL mixture MDPs with the finite action space.
-This achievement confirms that our result is indeed optimal with respect to the dependence on $d$ and $K$, only differing by a $\mathcal{O}(H^{1/2})$ factor.
+(2024) proposed a new reduction that bridges MNL mixture MDPs with linear mixture MDPs by approximating MNL functions to linear functions.
+Leveraging this new reduction, they established a problem-independent $\Omega(dH^{3/2}\sqrt{K})$ lower bound for the finite action setting.
+This achievement confirms that our result is optimal in $d$ and $K$, only loosing by an $\mathcal{O}(H^{1/2})$ factor.
 Dependence on $H$.
-By the discussion in Remark 7, we note that our result is optimal with respect to $d$ and $K$, but loosing by a $\mathcal{O}(H^{1/2})$ factor.
+By the discussion in Remark 5, we note that our result is optimal with respect to $d$ and $K$, but loosing by an $\mathcal{O}(H^{1/2})$ factor.
 We discuss the challenges in improving the dependence on $H$.
@@
 The pioneering work by Ayoub et al.
-(2020) achieved a regret bound of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ for linear mixture MDPs, which matches our results in Theorem 2, differing only on the lower order term.
+(2020) achieved a regret bound of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ for linear mixture MDPs, which matches our results in Theorem 1, differing only on the lower-order term.
 Later, Zhou et al.
@@
 7 Conclusion and Future Work
-In this work, we study MNL mixture MDPs that employ multinomial logit function approximation to ensure valid probability distributions over the state space.
-We address both the computational and statistical challenges for this problem.
-Specifically, we first propose an algorithm based on the online Newton step that attains the $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode.
-Next, we propose an enhanced algorithm that leverages local information to improve the statistical efficiency.
-It not only maintains $\mathcal{O}(1)$ computational and storage costs but also achieves an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, nearly matching the result of linear function approximation from both computational and statistical perspectives.
-Finally, we establish the first lower bound for MNL mixture MDPs, justifying the optimality of our results in $d$ and $K$.
+In this work, we addressing both statistical and computational challenges for MNL mixture MDPs, which leverage MNL function approximation to ensure valid probability distributions.
+Specifically, we propose a statistically efficient algorithm that achieve a regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$, eliminating the dependence on $\kappa^{-1}$ in the dominant term for the first time.
+Then, we introduce a computationally enhanced algorithm that achieves the same regret but with only constant cost.
+Finally, we establish the first lower bound for this problem, justifying the optimality of our results in $d$ and $K$.
 There are several interesting directions for future work.
-First, there still exists a gap between the upper and lower bounds and how to close this gap is an open problem.
-Besides, while this work focuses on stochastic rewards, extending this model to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Zhao et al., 2022; Li et al., 2023) is an another important direction.
+First, there still exists a gap between the upper and lower bounds.
+How to close this gap remains an open problem.
+Besides, we focuses on stationary rewards in this work, extending MNL mixture MDPs to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Zhao et al., 2022; Li et al., 2023) is also an important direction.
 Acknowledgments
 This research was supported by National Science and Technology Major Project (2022ZD0114800) and NSFC (U23A20382, 62206125). Peng Zhao was supported in part by the Xiaomi Foundation.
-Appendix A Properties of the Multinomial Logit Function
+Appendix A Notations
+In this section, we collect the notations used in the paper in Table 2.
+Table 2: Notations used in the regret analysis.
+Notation Definition and description
+$\ell_{k,h}(\theta)$ $\triangleq-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta)$, per-episode loss function at episode $k$ and stage $h$
+$g_{k,h}(\theta)$ $\triangleq\nabla\ell_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, gradient of loss $\ell_{k,h}(\theta)$
+$H_{k,h}(\theta)$ $\triangleq\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime},\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
+$\mathcal{L}_{k,h}(\theta)$ $\triangleq\sum_{i=1}^{k-1}\ell_{i,h}(\theta)+\frac{\lambda_{k}}{2}\lVert\theta\rVert_{2}^{2}$, the cumulative MLE loss
+$\mathcal{G}_{k,h}(\theta)$ $\triangleq\nabla\mathcal{L}_{k,h}(\theta)=\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta)-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}+\lambda_{k}\theta$, gradient of $\mathcal{L}_{k,h}(\theta)$
+$\mathcal{H}_{k,h}(\theta)$ $\triangleq\nabla^{2}\mathcal{L}_{k,h}(\theta)=\sum_{i=1}^{k-1}H_{i,h}(\theta)+\lambda_{k}I_{d}$, Hessian of MLE loss $\mathcal{L}_{k,h}(\theta)$
+$\widehat{\theta}_{k,h}$ $\triangleq\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\mathcal{L}_{k,h}(\theta)$, the MLE estimator at episode $k$ and stage $h$
+$\mathcal{H}_{k,h}$ $\triangleq\mathcal{H}_{k,h}(\widetilde{\theta}_{i+1,h})=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda_{k}I_{d}$, the cumulative look ahead Hessian
+~$\widetilde{\mathcal{H}}_{k,h}$ $\triangleq\mathcal{H}_{k,h}+\eta H_{k,h}(\widetilde{\theta}_{k,h})$, the sum of look ahead Hessian and the Hessian of current loss
+$\widetilde{\theta}_{k+1,h}$ $\triangleq\operatorname*{arg\,min}_{\theta\in\Theta}\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\widetilde{\mathcal{H}}_{k,h}}^{2}$, the OMD estimator
+Appendix B Properties of Multinomial Logit Function
 This section collects several key properties of the multinomial logit function used in the paper.
-Without loss of generality, we assume $\forall\mathcal{S}_{h,s,a},\exists s^{\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi(s^{\prime}\>|\>s,a)=\mathbf{0}$.
-Otherwise, we can always define a new feature mapping $\phi^{\prime}(s^{\prime\prime}\>|\>s,a)=\phi(s^{\prime}\>|\>s,a)-\phi(s^{\prime\prime}\>|\>s,a)$ for any $s^{\prime\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi^{\prime}(s^{\prime}\>|\>s,a)=\mathbf{0}$ and the transition induced by $\phi^{\prime}$ is the same as that induced by $\phi$.
-We denote this state as $\dot{s}_{h,s,a}$ and $\dot{\mathcal{S}}_{h,s,a}=\mathcal{S}_{h,s,a}\backslash\{\dot{s}_{h,s,a}\}$.
-Recall the following definitions in the paper:
-$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),$
-$g_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$
-$H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$
-$\widehat{\Sigma}_{k,h}$ $=\frac{\kappa}{2}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}+\lambda I_{d},$
-$\mathcal{H}_{k,h}$ $=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d},$
+Without loss of generality, we assume $\forall\mathcal{S}_{h,s,a},\exists\dot{s}_{h,s,a}\in\mathcal{S}_{h,s,a}$ such that $\phi(\dot{s}\>|\>s,a)=\mathbf{0}$.
+Otherwise, we can always define a new feature mapping $\phi^{\prime}(s^{\prime\prime}\>|\>s,a)=\phi(s^{\prime}\>|\>s,a)-\phi(s^{\prime\prime}\>|\>s,a)$ for any $s^{\prime\prime}\in\mathcal{S}_{h,s,a}$ such that $\phi^{\prime}(s^{\prime}\>|\>s,a)=\mathbf{0}$ and the transition kernel induced by $\phi^{\prime}$ is the same as that induced by $\phi$.
+Furthermore, We denote the set $\dot{\mathcal{S}}_{h,s,a}=\mathcal{S}_{h,s,a}\backslash\{\dot{s}_{h,s,a}\}$.
+First, we introduce the definition of self-concordant-like functions and demonstrate that the MNL loss function is self-concordant-like.
+Definition 3 (Self-concordant-like function, Tran-Dinh et al. [2015]).
+A convex function $f\in\mathcal{C}^{3}\left(\mathbb{R}^{m}\right)$ is $M$-self-concordant-like function with constant $M$ if:
+$\left|\psi^{\prime\prime\prime}(s)\right|\leqslant M\|\mathbf{b}\|_{2}\psi^{\prime\prime}(s).$
+for $s\in\mathbb{R}$ and $M>0$, where $\psi(s):=f(\mathbf{a}+s\mathbf{b})$ for any $\mathbf{a},\mathbf{b}\in\mathbb{R}^{m}$.
+Proposition 1.
+The per-episode MNL loss $\ell_{k,h}(\theta)$ and the cumulative MNL loss $\mathcal{L}_{k,h}(\theta)$ are both $3\sqrt{2}$-self-concordant-like for all $k\in[K],h\in[H]$.
+Proof.
+By proposition B.1 in Lee and Oh [2024], the per-episode MNL loss function $\ell_{k,h}(\theta)$ is $3\sqrt{2}$-self-concordant-like.
+Then, the cumulative MNL loss function $\mathcal{L}_{k,h}(\theta)$ is the sum of self-concordant-like functions and a quadratic function, it is also $3\sqrt{2}$-self-concordant-like.
+∎
+Lemma 4 (Zhang and Sugiyama [2023, Lemma 1]).
+Let $\ell(\mathbf{z},y)=\sum_{k=0}^{K}\mathbf{1}\{y=k\}\cdot\log\left(\frac{1}{[\sigma(\mathbf{z})]_{k}}\right)$ where $\sigma(\mathbf{z})_{k}=\frac{e^{z_{k}}}{\sum_{j=0}^{K}e^{z_{j}}}$, $\mathbf{a}\in[-C,C]^{K}$, $y\in\{0\}\cup[K]$ and $\mathbf{b}\in\mathbb{R}^{K}$ where $C>0$.
+Then, we have
+$\ell(\mathbf{a},y)\geq\ell(\mathbf{b},y)+\nabla\ell(\mathbf{b},y)^{\top}(\mathbf{a}-\mathbf{b})+\frac{1}{\log(K+1)+2(C+1)}(\mathbf{a}-\mathbf{b})^{\top}\nabla^{2}\ell(\mathbf{b},y)(\mathbf{a}-\mathbf{b}).$
+Then, we show the Hessian of the MNL loss function is positive semi-definite.
 Lemma 5.
@@
 Then, we have
-$H_{k,h}(\theta)$
-=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
-=\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
+$H_{k,h}(\theta)=$ $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}$
 =\ = $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\frac{1}{2}\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\sum_{s^{\prime\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\left(\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}+\phi_{k,h}^{s^{\prime\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\right)$
@@
 where the last inequality holds by the definition of $\kappa$ in Assumption 1. This finishes the proof. ∎
+Next, we show several concentration inequalities commonly used in the analysis.
 Lemma 6.
-Suppose $\lambda\geq 1$, define $\widetilde{\phi}_{k,h}^{s^{\prime}}=\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime\prime}}$, for any $k\in[K],h\in[H]$, the following statements hold:
-$(\textnormal{I})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}^{2}\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)$
-(II$)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)( I$
-$(\textnormal{III})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
-$(\textnormal{IV})\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
-$(\textnormal{V})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
+Suppose $\lambda_{k}\geq 1$, for any $k\in[K],h\in[H]$, for the quantities in Table 2 and define
+$\bar{\phi}_{k,h}^{s^{\prime}}=\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\theta_{h}^{*})\phi_{k,h}^{s^{\prime\prime}},\quad\widetilde{\phi}_{k,h}^{s^{\prime}}=\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime\prime}}.$
+Then, the following statements hold:
+$(\textnormal{I})\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}^{2}\leq 2d\log\left(1+\frac{k}{\lambda_{k}d}\right)$
+(II$)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda_{k}d}\right)( I$
+$(\textnormal{III})\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda_{k}d}\right)$
+$(\textnormal{IV})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}(\theta)}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right),\forall\theta\in\Theta$
+$(\textnormal{V})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right).$
 Proof.
 We prove the five statements individually.
-Proof of statement (I).
-By the definition that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$, we have
-$\det(\widehat{\Sigma}_{k+1,h})=\det(\widehat{\Sigma}_{k,h})\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{k,h}^{-1}}\Big{)}.$
-Then, we get
-$\sum_{i=1}^{k}\log\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{i,h}^{-1}}\Big{)}\leq\log\left(\frac{\det(\widehat{\Sigma}_{k+1,h})}{\det(\widehat{\Sigma}_{1,h})}\right)\leq d\log\left(1+\frac{kU}{\lambda d}\right),$ (12)
-where the last inequality holds by determinant trace inequality in Lemma 16.
-Since $\lambda\geq 1$, we have $\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\leq 1$, thus, we have
-$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\min\bigg{\{}1,\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\bigg{\}}$
-$\leq\frac{4}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\widehat{\Sigma}_{i,h}^{-1}}\Big{)}$
-$\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right),$
-where the first inequality holds by the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$ and the last inequality holds by (12).
-Proof of statement (II).
+Proof of statement (I).  By the definition of $H_{k,h}(\theta)$, we have
+$H_{i,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)p_{i,h}^{s^{\prime\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime\prime}})^{\top}$
+$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}]-\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{(}\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime\prime}}]\big{)}^{\top}$
+$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}\big{[}(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})^{\top}\big{]}$ (12)
+Thus, we have $H_{i,h}(\theta_{h}^{*})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})(\bar{\phi}_{i,h}^{s^{\prime}})(\bar{\phi}_{i,h}^{s^{\prime}})^{\top}$. Then, we get
+$\mathcal{H}_{i+1,h}(\theta_{h}^{*})\succeq\mathcal{H}_{i,h}(\theta_{h}^{*})+\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})(\bar{\phi}_{i,h}^{s^{\prime}})(\bar{\phi}_{i,h}^{s^{\prime}})^{\top}$
+As a result, we have
+$\det(\mathcal{H}_{i+1,h}(\theta_{h}^{*}))\geq\det(\mathcal{H}_{i,h}(\theta_{h}^{*}))\Big{(}1+\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}\Big{)}.$
+Since $\lambda\geq 1$, we have $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}\leq 1$.
+Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$, we get
+$\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}\leq$ $2\sum_{i=1}^{k}\log\Big{(}1+\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}\Big{)}$
+[…]
+[…]
+where the last inequality holds by the determinant inequality in Lemma 14.
+Proof of statement (II). The proof is same as that of (I), except that we replace $\theta_{h}^{*}$ with $\widetilde{\theta}_{i+1,h}$.
+Proof of statement (III).
 By Lemma 5, we have $H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$.
+The remaining proof is the same as the proof of statement (I).
+Proof of statement (IV).
+By Lemma 5, we have $\forall\theta\in\Theta$, it holds that $\mathcal{H}_{k+1,h}(\theta)\succeq\mathcal{H}_{k,h}(\theta)+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$.
+Since $\lambda\geq 1$, we have $\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}(\theta)}\leq\kappa$.
+Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$.
+By a similar analysis as the statement (I), we have
+$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\Big{)}$
+$\leq\frac{2}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\kappa\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\Big{)}$
+$\leq\frac{2}{\kappa}\log\left(\frac{\det(\mathcal{H}_{k+1,h}(\theta))}{\det(\mathcal{H}_{1,h}(\theta))}\right)$
+$\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right).$
+This finishes the proof of statement (IV).
+Proof of statement (V).  By (12), we have
+$H_{i,h}(\widetilde{\theta}_{i+1,h})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\widetilde{\theta}_{i+1,h})(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}.$
 Thus, we have
-$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})p_{k,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{k+1,h})\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
-Then, we get
-$\det(\mathcal{H}_{i+1,h})\geq\det(\mathcal{H}_{i,h})\Big{(}1+\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}\Big{)}.$
-Since $\lambda\geq 1$, we have $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}\leq 1$.
-Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$, we get
-$\sum_{i=1}^{k}\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}$
+$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\widetilde{\phi}_{k,h}^{s^{\prime}}(\widetilde{\phi}_{k,h}^{s^{\prime}})^{\top}.$
+Then, the remaining proof is similar to the proof of statement (III). ∎
+Appendix C Useful Lemmas for MNL Mixture MDPs
+In this section, we present some useful lemmas that are commonly used in the analysis.
+C.1 Useful Lemmas
+Lemma 7.
+For any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$ and positive semi-definite matrix $\Lambda$, suppose $\lVert\theta_{1}-\theta_{2}\rVert_{\Lambda}\leq\beta$.
+Then, for any $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds
+$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{2})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{1st}}+\epsilon_{s,a}^{\mathtt{2nd}}.$
+where
+$\epsilon_{s,a}^{\mathtt{1st}}$ $=H\beta\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{1})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\Lambda^{-1}},$
+$\epsilon_{s,a}^{\mathtt{2nd}}$ $=\frac{5}{2}H\beta^{2}\max_{s^{\prime}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\Lambda^{-1}}^{2}.$
+Lemma 8.
+Suppose $\forall(k,h,s,a)\in K\times[H]\times\mathcal{S}\times\mathcal{A}$ and $\widehat{\theta}_{k,h}\in\mathbb{R}^{d}$, it holds that $\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}$ where
+$\widehat{\mathcal{C}}_{k,h}=\left\{\theta\ \Big{|}\ \bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)V(s^{\prime})}\bigg{|}\leq\Gamma_{k,h,s,a}\right\}.$ (13)
+Define
+$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\operatorname*{arg\,max}_{\theta\in\mathcal{C}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})\bigg{]}_{[0,H]},$ (14)
+or,
+$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{k,h,s,a}\bigg{]}_{[0,H]},$ (15)
+where $\widehat{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s,a)$.
+Select the action as $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
+Then, for any $\delta\in(0,1]$, then it holds that
+$Q_{h}^{*}(s,a)\leq\widehat{Q}_{k,h}(s,a)\leq r_{h}(s,a)+\mathbb{P}_{h}\widehat{V}_{k,h+1}(s,a)+2\Gamma_{k,h,s,a}.$
+Lemma 9.
+Suppose Lemma 8 holds. Then, it holds that
+$\operatorname*{Reg}(K)\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{k,h,s_{k,h},a_{k,h}}+H\sqrt{2KH\log(2/\delta)}.$
+C.2 Proof of Lemma 7
+Proof.
+By the second-order Taylor expansion at $\theta_{1}$, there exists $\bar{\theta}=\nu\theta_{1}+(1-\nu)\theta_{2}$ for some $\nu\in[0,1]$, such that
+${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{2})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})V(s^{\prime})}$
+=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\theta_{1})^{\top}(\theta_{2}-\theta_{1})V(s^{\prime})+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\theta_{2}-\theta_{1})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\theta_{2}-\theta_{1})V(s^{\prime})$
+The gradient of $p_{s,a}^{s^{\prime}}(\theta)$ is given by
+$\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}.$
+For the first-order term, we have
+$\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\theta_{1})^{\top}(\theta_{2}-\theta_{1})V(s^{\prime})$
+=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})(\phi_{s,a}^{s^{\prime}})^{\top}(\theta_{2}-\theta_{1})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{1})(\phi_{s,a}^{s^{\prime\prime}})^{\top}(\theta_{2}-\theta_{1})V(s^{\prime})$
+[…]
+=\ = $H\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}^{+}}p_{s,a}^{s^{\prime}}(\theta_{1})\bigg{(}\Big{(}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{1})\phi_{s,a}^{s^{\prime\prime}}\Big{)}^{\top}\left(\theta_{2}-\theta_{1}\right)\bigg{)}$
 […]
@@
 […]
-Proof of statement (III).  By Lemma 5, we have
-$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k+1,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
-Since $\lambda\geq 1$, we have $\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\leq\kappa$.
-Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$.
-By a similar analysis as the statement (2), we have
-$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\Big{)}$
-$\leq\frac{2}{\kappa}\log\left(\frac{\det(\mathcal{H}_{k+1,h})}{\det(\mathcal{H}_{1,h})}\right)$
-$\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right).$
-This finishes the proof.
-Proof of statement (IV).  By the definition of $H_{k,h}(\theta)$, we have
-$H_{i,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)p_{i,h}^{s^{\prime\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime\prime}})^{\top}$
-$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}]-\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{(}\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{)}^{\top}$
-$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}\big{[}(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})^{\top}\big{]}$
-Thus, we have
-$H_{i,h}(\widetilde{\theta}_{i+1,h})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\widetilde{\theta}_{i+1,h})(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}.$ (13)
-Then, we get
-$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}(\widetilde{\theta}_{k+1,h})(\widetilde{\phi}_{k,h}^{s^{\prime}})(\widetilde{\phi}_{k,h}^{s^{\prime}})^{\top}$
-The remaining proof is similar to the proof of statement (2).
-Proof of statement (V).  By (13), we have
-$H_{i,h}(\widetilde{\theta}_{i+1,h})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\widetilde{\theta}_{i+1,h})(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}.$
-Then, the remaining proof is similar to the proof of statement (3). ∎
-Appendix B Proof of Lemma 1
-B.1 Main Proof
-Proof.
-The main proof follows the proof of Theorem 1 in Zhang et al. [2016]. We define
-$\bar{f}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[f_{k,h}(\theta)\mid\mathcal{F}_{k,h}],\quad\bar{g}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[g_{k,h}(\theta)\mid\mathcal{F}_{k,h}].$
-By Property 1, we have $f_{k,h}(\theta)$ is exponential concave such that
-$f_{k,h}(\widehat{\theta}_{k,h})\leq f_{k,h}(\theta_{h}^{*})+\langle g_{k,h}(\theta_{h}^{*}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert(\widehat{\theta}_{k,h}-\theta_{h}^{*})\rVert_{W_{k,h}}.$
-Taking expectations on both sides, we have
-$\bar{f}_{k,h}(\widehat{\theta}_{k,h})\leq\bar{f}_{k,h}(\theta_{h}^{*})+\langle\bar{g}_{k,h}(\theta_{h}^{*}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert(\widehat{\theta}_{k,h}-\theta_{h}^{*})\rVert_{W_{k,h}}.$ (14)
-Based the property of KL diverge, we ensure the true parameter $\theta_{h}^{*}$ is the minimizer of the expected loss function by the following lemma.
-Lemma 7.
-For any $\theta\in\Theta$, we have $\bar{f}_{k,h}(\theta_{h}^{*})\leq\bar{f}_{k,h}(\theta)$.
-Combining Lemma 7 and Equation (14), we have
-00 0 $\leq\bar{f}_{k,h}(\widehat{\theta}_{k,h})-\bar{f}_{k,h}(\theta_{h}^{*})$
-$\leq\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}$
-$\leq\langle g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle.$ (15)
-By standard analysis of OMD in Lemma 17, it holds
-$2g_{k,h}(\widehat{\theta}_{k,h})(\widehat{\theta}_{k,h}-\theta_{h}^{*})\leq\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}}+\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}-\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}.$ (16)
-Combining Equation (15) and Equation (16), we have
-00 0 $\leq\frac{1}{2}\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}+\frac{1}{2}\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}-\frac{1}{2}\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$
-$\hskip 85.35826pt-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle.$ (17)
-First, we consider the first term . We show the gradient can be bounded by the following lemma.
-Lemma 8.
-For any positive semi-definite matrix $Z$, it holds that
-$\lVert g_{k,h}(\theta)\rVert_{Z}^{2}\leq 4\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{Z}^{2}.$ (18)
-Then, we consider the second term. Note that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}W_{k,h}$, we have
-$\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}=\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{\widehat{\Sigma}_{k,h}}^{2}+\frac{\kappa}{2}\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{W_{k,h}}^{2}.$ (19)
-Next, we bound $\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle$, which is a martingale difference sequence.
-Lemma 9.
-For any $\delta\in(0,1)$ and $\theta_{1},\ldots,\theta_{k},\theta^{*}\in[0,B]^{d}$, with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
-$\sum_{i=1}^{k}\langle\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i}),\theta_{i}-\theta_{h}^{*}\rangle\leq\frac{\kappa}{4}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}+(\frac{4}{\kappa}+8B)\gamma_{k}.$ (20)
-where $\gamma_{k}=\log\frac{2k^{2}H\log(kU)}{\delta}$.
-Combining (17), (18), and (19), we have
-$\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ $\leq\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}-\kappa\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}$
-$\hskip 28.45274pt+2\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle$
-$\leq\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k,h}}^{2}+4\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}+(\frac{8}{\kappa}+16B)\gamma_{k}$ (21)
-where the first inequality holds by rearranging the terms in (17), the second holds by (18), (19) and (20). Summing (21) from $i=1$ to $k$, we have
-$\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ $\leq\lambda B^{2}+4\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i+1,h}^{-1}}^{2}+(\frac{8}{\kappa}+16B)\gamma_{k}$
-$\leq\lambda B^{2}+\frac{8}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)+(\frac{8}{\kappa}+16B)\gamma_{k},$
-where the second inequality holds by $\lVert\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq 1$ and $\widehat{\Sigma}_{i,h}\geq I,\forall i\in[K]$, and the last inequality is by Lemma 6.
-This finishes the proof.
-∎
-B.2 Proof of Auxiliary Lemmas
-In this section, we provide the proofs of the lemmas used in Appendix B.1.
-B.2.1 Proof of Lemma 7
-Proof.
-By the definition of $\bar{f}_{k,h}(\theta)$, we have
-$\bar{f}_{k,h}(\theta)-\bar{f}_{k,h}(\theta_{h}^{*})$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log\frac{p_{k,h}^{s^{\prime}}(\theta_{h}^{*})}{p_{k,h}^{s^{\prime}}(\theta)}\geq 0,$
-where the last inequality is due to $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log\frac{p_{k,h}^{s^{\prime}}(\theta_{h}^{*})}{p_{k,h}^{s^{\prime}}(\theta)}$ is the Kullback-Leibler divergence between $p_{k,h}(\theta_{h}^{*})$ and $p_{k,h}(\theta)$, which always is non-negative.
-∎
-B.2.2 Proof of Lemma 8
-Proof.
-For any positive semi-definite matrix $Z$,
-$(x_{i}-x_{j})^{\top}Z(x_{i}-x_{j})=x_{i}^{\top}Zx_{i}-x_{i}^{\top}Zx_{j}-x_{j}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq 0,\quad\forall x_{i},x_{j}\in\mathbb{R}^{d}.$ (22)
-which implies $x_{i}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq x_{i}^{\top}Zx_{j}+x_{j}^{\top}Zx_{i},\quad\forall x_{i},x_{j}\in\mathbb{R}^{d}$.
-Let $x_{i}=(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
-$\lVert g_{k,h}(\theta)\rVert_{Z}^{2}$
-=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})(p_{k,h}^{s^{\prime\prime}}(\theta)-y_{k,h}^{s^{\prime\prime}})\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime\prime}}$
-=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})^{2}\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime}}$
-$\hskip 56.9055pt+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})(p_{k,h}^{s^{\prime\prime}}(\theta)-y_{k,h}^{s^{\prime\prime}})(\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime\prime}}+\phi_{k,h}^{s^{\prime\prime}}Z\phi_{k,h}^{s^{\prime}})$
+where in the first inequality, we denote $\mathcal{S}_{h,s,a}^{+}$ as the subset of $\mathcal{S}_{h,s,a}$ such that $(\phi_{s,a}^{s^{\prime}})^{\top}(\theta_{2}-\theta_{1})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{2})(\phi_{s,a}^{s^{\prime\prime}})^{\top}(\theta_{2}-\theta_{1})$ is non-negative, the second inequality holds by the Holder’s inequality, and the third inequality is by the condition $\lVert\theta_{1}-\theta_{2}\rVert_{\Lambda}\leq\beta$.
+For the second-order term, let $u_{s,a}^{s^{\prime}}(\theta)=(\phi_{s,a}^{s^{\prime}})^{\top}\theta$ and $p_{s,a}^{s^{\prime}}(u)=\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}}\exp(u_{s,a}^{s^{\prime\prime}})}$, further define
+$F(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})},\quad\widetilde{F}(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})V(s^{\prime})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})}.$
+Then, we have
+$\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\theta_{2}-\theta_{1})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\theta_{2}-\theta_{1})V(s^{\prime})$
+=\ = $\frac{1}{2}\big{(}u(\theta_{2})-u(\theta_{1})\big{)}^{\top}\nabla^{2}\widetilde{F}(u(\bar{\theta}))\big{(}u(\theta_{2})-u(\theta_{1})\big{)}$
+=\ = $\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\big{(}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{)}^{\top}\frac{\partial^{2}\widetilde{F}(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{)}$
 […]
-=\ = $2\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})^{2}\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime}}$
+where the inequality holds by $V(s)\in[0,H],\forall s$.
+According to Lemma 17, we have (omit the subscript $\mathcal{S}_{h,s,a}$ for simplicity):
+$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{\rvert}\cdot\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\cdot\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{\rvert}$
 […]
-This finishes the proof. ∎
-B.2.3 Proof of Lemma 9
-Proof.
-First, notice that $\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$ is a martingale difference sequence. Also, we have
-$\quad\left|\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
-$\leq\left|\left(\bar{g}_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|+\left|\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
-$\leq\lVert\bar{g}_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}+\lVert g_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}$
-$\leq 4\sqrt{2}B,$
-where the last inequality holds by $\lVert\bar{g}_{i,h}(\theta_{i})\rVert_{2}=\lVert\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta)-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq\sqrt{2}$.
-We define the martingale $M_{k,h}=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$, and define the conditional variance $\sigma_{i}^{2}$ as
-$\sigma_{k,h}^{2}$ $=\mathbb{E}[M_{k,h}^{2}\mid\mathcal{F}_{i-1}]$
-$=\sum_{i=1}^{k}\mathbb{E}_{y_{k,h}}\left[\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})^{2}\right]$
-$\leq\sum_{i=1}^{k}\mathbb{E}_{y_{k,h}}\left[\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})^{2}\right]$
-$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})^{2}$
-$=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}$
-$\triangleq A_{k,h}.$
-where the first inequality is due to the fact that $\mathbb{E}[(\xi-\mathbb{E}[\xi])^{2}]\leq\mathbb{E}[\xi^{2}]$ for any random variable $\xi$.
-Note that $A_{k,h}$ is a random variable, so we cannot directly apply the Bernstein inequality to $M_{k,h}$.
-Instead, we consider the following two cases:(i) $A_{k,h}\leq\frac{4B^{2}}{kU}$ and (ii) $A_{k,h}>\frac{4B^{2}}{kU}$.
-Case (i): $A_{k,h}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\leq\frac{4B^{2}}{kU}$. Then, we have
-$M_{k,h}$ $=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$
-$=\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta_{i})-y_{i,h}^{s^{\prime}})(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})$
-$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lvert(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})\rvert$
-$\leq\sqrt{kU\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}((\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*}))^{2}}$
-$\leq 2B.$
-where the second equality is due to the definition of $\bar{g}_{i,h}(\theta_{i})$ and $g_{i,h}(\theta_{i})$, the first inequality holds by $p_{i,h}^{s^{\prime}}(\theta_{i})-y_{i,h}^{s^{\prime}}\in[-1,1]$, the second inequality holds by the Cauchy-Schwarz inequality, and the last inequality holds by the condition of case (i).
-Case (ii): $A_{k,h}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{j,h}}^{2}>\frac{4B^{2}}{kU}$.
-We have both a lower and upper bound for $A_{k,h}$, i.e., $\frac{4B^{2}}{kU}<A_{k,h}\leq 4B^{2}kU$.
-Then, we can use the peeling process to bound $M_{k,h}$ as follows:
-$\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3}\right]$
-=\ = $\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3},\frac{4B^{2}}{kU}<A_{k,h}\leq 4kUB^{2}\right]$
-=\ = $\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3},\frac{4B^{2}}{kU}<A_{k,h}\leq 4kUB^{2},\sigma_{k,h}\leq A_{k,h}\right]$
+$+\frac{3H}{2}\sum_{s^{\prime}}\big{(}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta})).$ (17)
+To bound the first term, by applying the AM-GM inequality, we obtain
+$H\sum_{s^{\prime}}\sum_{s^{\prime\prime}\neq s^{\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{\rvert}\cdot p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}\cdot\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{\rvert}$
 […]
 […]
+$\hskip 56.9055pt+\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{)}^{2}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}$
 […]
-where $m=2\log_{2}(kU)$, and the last inequality follows the Bernstein inequality for martingales.
-Combining above two cases, letting $\tau=\log{\frac{mk^{2}}{\delta/H}}$ and taking the union bound over $k$ and $h\in[H]$, we have with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
-$M_{k,h}$ $=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$
-$\leq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3}+4\sqrt{2}B$
-$\leq 2\sqrt{\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\log\frac{2k^{2}H\log(kU)}{\delta}}+8B\left(1+\log\frac{2k^{2}H\log(kU)}{\delta}\right)$
-$=2\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}+8B\left(1+\gamma_{k}\right),~{}$ (23)
-where $\gamma_{k}=\log\frac{2k^{2}H\log(kU)}{\delta}$.
-Then, applying $uv\leq cu^{2}+v^{2}/(4c)$ for any $c,u,v>0$ with $c=2/\kappa$, we have
-$\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}\leq\frac{2\gamma_{k}}{\kappa}+\frac{\kappa}{8}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}.$ (24)
-Combining (23) and (24), we have
-$\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\leq\frac{\kappa}{4}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}+(\frac{4}{\kappa}+8B)\gamma_{k}.$
-This finishes the proof. ∎
-Appendix C Proof of Lemma 2
+Plugging (18) into (17), we have
+$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{\rvert}\cdot\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\cdot\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{\rvert}$
+[…]
+=\ = $\frac{5H}{2}\sum_{s^{\prime}}\big{(}(\phi_{s,a}^{s^{\prime}})^{\top}(\theta_{2}-\theta_{1})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta}))$
+[…]
+where the last inequality holds by the condition $\lVert\theta_{1}-\theta_{2}\rVert_{\Lambda}\leq\beta$.
+Finally, combining (16) and (19) finishes the proof. ∎
+C.3 Proof of Lemma 8
 Proof.
-The gradient of $p_{s,a}(\theta)$ is given by
-$\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}.$
-By the mean value theorem, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
-$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}$
-=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})\bigg{|}$
-=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})\bigg{|}$
-=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})\Big{(}V(s^{\prime})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})V(s^{\prime\prime})\Big{)}\bigg{|}$
-[…]
-[…]
-[…]
-[…]
-where the first inequality is by $V:\mathcal{S}\to[0,H]$, the second inequality is by the definition of $\phi_{k,h}^{s^{\prime}}$, the third inequality is by the Holder’s inequality, the last inequality is by Lemma 1.
-∎
-Appendix D Proof of Theorem 1
-D.1 Main Proof
-To prove the theorem, we first introduce the following lemma.
-Lemma 10.
-Suppose for all $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds that
-$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\Gamma_{h,s,a}.$ (26)
-Define
-$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{h,s,a}\bigg{]}_{[0,H]}.$ (27)
-Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
-$\operatorname*{Reg}(K)\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+H\sqrt{2KH\log(2/\delta)}.$
-Proof of Theorem 1.
-Substituting $\Gamma_{h,s,a}=H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}$ into Lemma 10, we have
-$\operatorname*{Reg}(K)$ $\leq 2H\widehat{\beta}_{k}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}+H\sqrt{2KH\log(2/\delta)}$
-$\leq 2H\widehat{\beta}_{k}\sum_{h=1}^{H}\sqrt{K\sum_{k=1}^{K}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}^{2}}+H\sqrt{2KH\log(2/\delta)}$
-$\leq 2H^{2}\widehat{\beta}_{k}\sqrt{\frac{4dK}{\kappa}\log\left(1+\frac{KU}{d\lambda}\right)}+H\sqrt{2KH\log(2/\delta)}\leq\widetilde{\mathcal{O}}({\kappa}^{-1}dH^{2}\sqrt{K}),$
-where the second inequality holds by the Cauchy-Schwarz inequality and the third inequality holds by Lemma 6. This finishes the proof. ∎
-D.2 Proof of Auxiliary Lemmas
-In this section, we provide the proofs of the lemmas used in Appendix D.1.
-First, we introduce the following lemma.
-Lemma 11.
-Suppose (26) and (27) in Lemma 10 holds for all $k\in[K],h\in[H]$. Then, for any $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$, it holds that
-$Q_{h}^{*}(s,a)\leq\widehat{Q}_{k,h}(s,a)\leq r_{h}(s,a)+\mathbb{P}_{h}V_{k,h+1}(s,a)+2\Gamma_{h,s,a}.$
+First, we prove the left-hand side of the lemma.
+We prove this by backward induction on $h$.
+For the stage $h=H$, by definition, we have $\widehat{Q}_{k,H}(s,a)=r_{H}(s,a)=Q_{H}^{*}(s,a),\widehat{V}_{k,H+1}(s)=0=V_{H+1}^{*}(s)$.
+Suppose the statement holds for $h+1$, we show it holds for $h$.
+By definition, if $\widehat{Q}_{k,h}(s,a)=H$, this holds trivially.
+Otherwise, we consider two cases:
+For $\widehat{Q}_{k,h}(s,a)$ defined in (14), we have
+$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+\operatorname*{arg\,max}_{\theta\in\mathcal{C}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})$
+$\geq r_{h}(s,a)+\operatorname*{arg\,max}_{\theta\in\mathcal{C}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)V_{k,h+1}^{*}(s^{\prime})$
+$\geq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V_{k,h+1}^{*}(s^{\prime})=Q_{h}^{*}(s,a).$
+where the first inequality is by the induction hypothesis, and the second inequality is due to $\theta_{h}^{*}\in\mathcal{C}_{k,h}$.
+For $\widehat{Q}_{k,h}(s,a)$ defined in (15), we have
+$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{h,s,a}$
+$\geq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}(\widehat{\theta}_{k,h})V_{k,h+1}^{*}(s^{\prime})+\Gamma_{h,s,a}$
+$\geq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}(\theta_{h}^{*})V_{k,h+1}^{*}(s^{\prime})=Q_{h}^{*}(s,a).$
+where the first inequality is by the induction hypothesis, and the second inequality is by (13).
+Then, we prove the right-hand side of the lemma.
+For $\widehat{Q}_{k,h}(s,a)$ defined in (14), we have
+$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+\operatorname*{arg\,max}_{\theta\in\mathcal{C}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})$
+$\leq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{k,h,s,a}$
+$\leq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})\widehat{V}_{k,h+1}(s^{\prime})+2\Gamma_{h,s,a},$
+where the inequality is by (13).
+For $\widehat{Q}_{k,h}(s,a)$ defined in (15), we have
+$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{k,h,s,a}$
+$\leq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})\widehat{V}_{k,h+1}(s^{\prime})+2\Gamma_{k,h,s,a},$
+where the inequality is by (13). ∎
+C.4 Proof of Lemma 9
 Proof.
-First, we prove the left-hand side of the lemma. We prove this by backward induction on $h$. For the stage $h=H$, by definition, we have
-$\widehat{Q}_{k,H}(s,a)=r_{H}(s,a)=Q_{H}^{*}(s,a),\quad\widehat{V}_{k,H+1}(s)=0=V_{H+1}^{*}(s).$
-Suppose the statement holds for $h+1$, we show it holds for $h$. By definition, if $\widehat{Q}_{k,h}(s,a)=H$, this holds trivially. Otherwise, we have
-$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}+\Gamma_{h,s,a}$
-$\geq r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})V_{k}^{*}+\Gamma_{h,s,a}\geq r_{h}(s,a)+p_{s,a}(\theta_{h}^{*})V_{k}^{*}=Q_{h}^{*}(s,a).$
-where the first inequality is by the induction hypothesis, and the second inequality is by (26).
-Then, we prove the right-hand side of the lemma. By the definition of $\widehat{Q}_{k,h}(s,a)$, we have
-$\widehat{Q}_{k,h}(s,a)=r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}+\Gamma_{h,s,a}\leq r_{h}(s,a)+p_{s,a}(\theta_{h}^{*})\widehat{V}_{k,h+1}+2\Gamma_{h,s,a},$
-where the inequality is by (26). ∎
-D.2.1 Proof of Lemma 10
-Proof.
-By the definition of $\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$, we have
+By the definition that $\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$, we have for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
 $\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})=$ $\sum_{k=1}^{K}Q_{1}^{*}(s_{k,1},\pi^{*}(s_{k,1}))-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$
@@
 […]
-where the first inequality is by Lemma 11 and the second is by $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
+where the first inequality is by Lemma 8 and $\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}$ with probability at least $1-\delta$, and the second inequality is by action selection $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
 By the right-hand side of Lemma 8, we have
 $\widehat{Q}_{1}(s_{k,1},a_{k,1})-V_{1}^{\pi_{k}}(s_{k,1})$
-=\ = $r(s_{k,1},a_{k,1})+\mathbb{P}_{1}\widehat{V}_{k,2}(s_{k,1},a_{k,1})++2\Gamma_{h,s_{k,1},a_{k,1}}-r(s_{k,1},a_{k,1})-\mathbb{P}_{1}V_{2}^{\pi_{k}}(s_{k,1},a_{k,1})$
+=\ = $r(s_{k,1},a_{k,1})+\mathbb{P}_{1}\widehat{V}_{k,2}(s_{k,1},a_{k,1})+2\Gamma_{h,s_{k,1},a_{k,1}}-r(s_{k,1},a_{k,1})-\mathbb{P}_{1}V_{2}^{\pi_{k}}(s_{k,1},a_{k,1})$
 […]
@@
 $\widehat{Q}_{1}(s_{k,1},a_{k,1})-V_{1}^{\pi_{k}}(s_{k,1})\leq 2\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+\sum_{h=1}^{H}\mathcal{M}_{k,h}$
-Summing over $k$, we have for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
+Summing over $k$, we have
 $\operatorname*{Reg}(K)\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+H\sqrt{2KH\log(2/\delta)}$
 where the inequality holds by the Azuma-Hoeffding inequality as $\mathcal{M}_{k,h}$ is a martingale difference sequence with $\mathcal{M}_{k,h}\leq 2H$. This finishes the proof. ∎
-Appendix E Proof of Lemma 3
+Appendix D Omitted Proofs for Section 3
+D.1 Proof of Claim 1
 Proof.
-The proof is similar to the proof of Theorem 3 in Zhang and Sugiyama [2023] and Lemma 1 of Lee and Oh [2024]. Define
-$\widetilde{f}_{k,h}(\theta)=f_{k,h}(\theta_{k,h})+\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{H_{k,h}(\widetilde{\theta}_{k,h})},$
-which is a second order approximation of the original function $f_{k,h}(\theta)$ at $\widetilde{\theta}_{k,h}$. Then, the update rule in (9) can be rewritten as
-$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{f}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}.$ (28)
-Then, we present the following lemma that bounds the estimation error of the update rule in (28).
-Lemma 12 (Lemma E.1 of Lee and Oh [2024]).
-For the update rule defined in (28), it holds that
-$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}\leq$ $2\eta\left(\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})\right)+4\lambda B$
+First, by the definition of MNL mixture MDP, we have $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ and $s^{\prime}\in\mathcal{S}_{h,s,a}$, it holds that $p_{s,a}^{s^{\prime}}(\theta)\geq\exp(-B)/(U\exp(B)),\forall\theta\in\mathbb{R}^{d}$, thus $\kappa^{*}\geq\kappa\geq 1/(U\exp(2B))^{2}$.
+Next, consider the state-action pair $(s,a)$ at stage $h$ with the maximum number of reachable states $U$, it is clear that $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})=1$.
+This implies that $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})p_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})=1$.
+Applying the pigeonhole principle, there exists $s^{\prime},s^{\prime\prime}\in\mathcal{S}_{h,s,a}$ such that $p_{s,a}^{s^{\prime}}(\theta_{h}^{*})p_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\leq 1/U^{2}$.
+Thus, we conclude that $\kappa\leq\kappa^{*}\leq 1/U^{2}$.
+This finishes the proof.
+∎
+Appendix E Omitted Proofs for Section 4
+E.1 Useful Lemma
+Lemma 10.
+[…]
+Proof.
+By the multivariate mean value theorem, we have
+$\mathcal{G}_{k,h}(\theta_{1})-\mathcal{G}_{k,h}(\theta_{2})=\nabla\mathcal{L}_{k,h}(\theta_{1})-\nabla\mathcal{L}_{k,h}(\theta_{2})=\int_{0}^{1}\nabla^{2}\mathcal{L}_{k,h}(\theta_{2}+t(\theta_{1}-\theta_{2}))\mathrm{d}t(\theta_{1}-\theta_{2}).$
+Hence, we have
+$\lVert\mathcal{G}(\theta_{1})-\mathcal{G}(\theta_{2})\rVert_{G_{k,h}^{-1}(\theta_{1},\theta_{2})}=\lVert\theta_{1}-\theta_{2}\rVert_{G_{k,h}(\theta_{1},\theta_{2})}.$
+where $G_{k,h}(\theta_{1},\theta_{2})=\int_{0}^{1}\nabla^{2}\mathcal{L}_{k,h}(\theta_{2}+t(\theta_{1}-\theta_{2}))\mathrm{d}t$.
+By self-concordant-like property of $\mathcal{L}_{k,h}$ in Proposition 1, we have $\mathcal{H}_{k,h}(\theta_{1})\preceq({1+3\sqrt{2}})G_{k,h}(\theta_{1},\theta_{2})$.
+As a result, we have
+$\left\|\theta_{1}-\theta_{2}\right\|_{\mathcal{H}_{k,h}\left(\theta_{1}\right)}$ $\leq(1+3\sqrt{2})^{1/2}\left\|\theta_{1}-\theta_{2}\right\|_{G_{k,h}\left(\theta_{1},\theta_{2}\right)}$
+$=(1+3\sqrt{2})^{1/2}\left\|\mathcal{G}_{k,h}\left(\theta_{1}\right)-\mathcal{G}_{k,h}\left(\theta_{2}\right)\right\|_{G_{k,h}^{-1}\left(\theta_{1},\theta_{2}\right)}$
+$\leq(1+3\sqrt{2})\left\|\mathcal{G}_{k,h}\left(\theta_{1}\right)-\mathcal{G}_{k,h}\left(\theta_{2}\right)\right\|_{\mathcal{H}_{k,h}^{-1}\left(\theta_{1}\right)}.$
+This finishes the proof. ∎
+E.2 Proof of Lemma 1
+Proof.
+Since $\widehat{\theta}_{k,h}$ minimizes $\mathcal{L}_{k,h}(\theta)$, we have $\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})=\bm{0}$. Thus, we have
+$\mathcal{G}_{k,h}(\theta_{h}^{*})-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})=\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta_{h}^{*})-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}+\lambda_{k}\theta_{h}^{*}.$
+Therefore, since $\lVert\theta_{h}^{*}\rVert\leq B$ and $\mathcal{H}_{k,h}(\theta_{h}^{*})\succeq\lambda_{k}I$, for any $\delta\in(0,1]$, with probability at least $1-\frac{\delta}{H}$,
+$\lVert\mathcal{G}_{k,h}(\theta_{h}^{*})-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}\leq$ $\Big{\lVert}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta_{h}^{*})-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}+\sqrt{\lambda_{k}}B$
+[…]
+where the last inequality holds by the Bernstein-type concentration inequality in Lemma 13.
+Then, by the determinant inequality in Lemma 14, we have $\det(\mathcal{H}_{k,h}(\theta_{h}^{*}))\leq\left(\lambda_{k}+\frac{k}{d}\right)^{d}$.
+Thus, we obtain
+$\lVert\mathcal{G}_{k,h}(\theta_{h}^{*})-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}\leq\left(B+\frac{1}{2}\right)\sqrt{\lambda_{k}}+\frac{2}{\sqrt{\lambda_{k}}}\log\left(\frac{4}{\delta}\left(1+\frac{kH}{d\lambda_{k}}\right)\right).$
+By the configuration that $\lambda_{k}=d\log(kH/\delta)$ and applying the union bound for $h\in[H]$, we have with probability at least $1-\delta$, for all $h\in[H]$ simultaneously, it holds that
+$\lVert\mathcal{G}_{k,h}(\theta_{h}^{*})-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}\leq(B+3)\sqrt{d\log(kH/\delta)}.$
+This finishes the proof. ∎
+E.3 Proof of Theorem 1
+Proof.
+By Lemma 10, with probability at least $1-\delta$, it holds that
+$\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}$ $\leq(1+3\sqrt{2})\lVert\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})-\mathcal{G}_{k,h}(\theta_{h}^{*})\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}\leq(1+3\sqrt{2})\widehat{\beta}_{k}.$
+where the last inequality holds the confidence set $\widehat{\mathcal{C}}_{k,h}$ in Lemma 1. Then, by Lemma 7, we have
+$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{1st}}+\epsilon_{s,a}^{\mathtt{2nd}}.$
+where
+$\epsilon_{s,a}^{\mathtt{1st}}$ $=(1+3\sqrt{2})H\widehat{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})},$
+$\epsilon_{s,a}^{\mathtt{2nd}}$ $=90H\widehat{\beta}_{k}^{2}\max_{s^{\prime}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}^{2}.$
+By Lemma 9, we have
+$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$ $\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}(\epsilon_{k,h}^{\mathtt{1st}}+\epsilon_{k,h}^{\mathtt{2nd}})+H\sqrt{2KH\log(2/\delta)}$ (20)
+Next, we bound $\epsilon_{k,h}^{\mathtt{1st}}$ and $\epsilon_{k,h}^{\mathtt{2nd}}$ respectively.
+Bounding $\epsilon_{k,h}^{\mathtt{1st}}$.  By statement (I) of Lemma 6, we have
+$\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\theta_{h}^{*})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}$
+[…]
+[…]
+By the configuration that $\widehat{\beta}_{k}=(B+3)\sqrt{d\log(k/\delta)}$, we have
+$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{1st}}\leq(1+3\sqrt{2})(B+3)dH^{2}\sqrt{K\log\left(1+\frac{K}{d\lambda_{K}}\right)\log\left(\frac{k}{\delta}\right)}.$ (21)
+Bounding $\epsilon_{k,h}^{\mathtt{2nd}}$.  By statement (IV) of Lemma 6, we have
+$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{2nd}}$ $=90H^{2}\sum_{k=1}^{K}\widehat{\beta}_{k}^{2}\max_{s^{\prime}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}^{2}$
+$\leq\frac{180}{\kappa}H^{2}\widehat{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda_{K}}\right)$
+$\leq\frac{180}{\kappa}(B+3)^{2}d^{2}H^{2}\log\left(1+\frac{K}{d\lambda_{K}}\right)\log\left(\frac{K}{\delta}\right)$ (22)
+where the first inequality holds by $sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}(\theta)}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right),\forall\theta\in\Theta$ in Lemma 6 and the second inequality holds by the configuration of $\widehat{\beta}_{k}$.
+Combining (20), (21), and (22), we have with probability at least $1-\delta$,
+$\operatorname*{Reg}(K)$ $\leq\sqrt{2dK\log\left(1+\frac{K}{d\lambda_{K}}\right)}+\frac{180}{\kappa}(B+3)^{2}d^{2}H^{2}\log\left(1+\frac{K}{d\lambda_{K}}\right)\log\left(\frac{K}{\delta}\right)$
+$\leq\widetilde{\mathcal{O}}\left(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2}\right).$
+This finishes the proof. ∎
+Appendix F Omitted Proofs for Section 5
+F.1 Useful Lemma
+Lemma 11.
+For any $k\in[K]$, $h\in[H]$, define the second-order approximation of the loss function $\ell_{k,h}(\theta)$ at the estimator $\widetilde{\theta}_{k,h}$ as ~$\widetilde{\ell}_{k,h}(\theta)=\ell_{k,h}(\theta_{k,h})+\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{H_{k,h}(\widetilde{\theta}_{k,h})}^{2}$.
+Then, for the following update rule
+$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{\ell}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}^{2},$
+it holds that
+$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}\leq$ $2\eta\left(\sum_{i=1}^{k}\ell_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}\ell_{i,h}(\widetilde{\theta}_{i+1,h})\right)+4\lambda B$
+$+12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}-\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}.$
+Proof.
+Based on the analysis of (implicit) OMD update (see Lemma 16), for any $i\in[K]$, we have
+~$\big{\langle}\nabla\widetilde{\ell}_{i,h}(\widetilde{\theta}_{i+1,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}\leqslant\frac{1}{2\eta}\left(\|\widetilde{\theta}_{i,h}-\theta_{h}^{*}\|_{\mathcal{H}_{i,h}}^{2}-\|\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\|_{\mathcal{H}_{i,h}}^{2}-\|\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\|_{\mathcal{H}_{i,h}}^{2}\right)$
+According to Lemma 4, we have
+$\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\ell_{i,h}\left(\theta_{h}^{*}\right)\leqslant\big{\langle}\nabla\ell_{i,h}(\widetilde{\theta}_{i+1,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}-\frac{1}{\zeta}\left\|\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\right\|_{\nabla^{2}\ell_{i,h}(\widetilde{\theta}_{i+1,h})}^{2},$
+where $\zeta=\log(K+1)+4$. Then, by combining the above two inequalities, we have
+$\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\ell_{i,h}(\theta_{h}^{*})$ $\leqslant\langle\nabla\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\nabla\widetilde{\ell}_{i,h}(\widetilde{\theta}_{i+1,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\rangle$
+$+\frac{1}{\zeta}\Big{(}\|\widetilde{\theta}_{i,h}-\theta_{h}^{*}\|_{\mathcal{H}_{i,h}}^{2}-\|\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\|_{\mathcal{H}_{i+1,h}}^{2}-\|\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\|_{\mathcal{H}_{i,h}}^{2}\Big{)}.$
+We can further bound the first term of the right-hand side as:
+$\big{\langle}\nabla\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\nabla\widetilde{\ell}_{i,h}(\widetilde{\theta}_{i+1,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}$
+=\ = $\big{\langle}\nabla\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\nabla\ell_{i,h}(\widetilde{\theta}_{i,h})-\nabla^{2}\ell_{i,h}(\widetilde{\theta}_{i,h})(\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}$
+=\ = $\big{\langle}D^{3}\ell_{i,h}({\xi}_{i+1})[\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}](\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}$
+[…]
+[…]
+where the second equality holds by the mean value theorem, the first inequality holds by the self-concordant-like property of $\ell_{i,h}(\cdot)$ in Proposition 1, and the last inequality holds by $\widetilde{\theta}_{i+1,h}$ and $\theta_{h}^{*}$ belong to $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, and ${\nabla^{2}\ell_{i,h}({\xi}_{i+1})}\preceq I_{d}$.
+Then, by taking the summation over $i$ and rearranging the terms, we obtain
+$\big{\|}\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\big{\|}_{\mathcal{H}_{k+1,h}}^{2}\leqslant$ $\zeta\left(\sum_{s=1}^{k}\ell_{s,h}\left(\theta_{h}^{*}\right)-\sum_{s=1}^{k}\ell_{s,h}(\widetilde{\theta}_{i+1,h})\right)+\|\widetilde{\theta}_{1,h}-\theta_{h}^{*}\|_{\mathcal{H}_{1,h}}^{2}$
+$+6\sqrt{2}B\zeta\sum_{s=1}^{k}\big{\|}\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\big{\|}_{2}^{2}-\sum_{s=1}^{k}\big{\|}\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\big{\|}_{\mathcal{H}_{i,h}}^{2}$
+[…]
+$+6\sqrt{2}B\zeta\sum_{s=1}^{k}\big{\|}\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\big{\|}_{2}^{2}-\sum_{s=1}^{k}\big{\|}\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\big{\|}_{\mathcal{H}_{i,h}}^{2},$
+where the last inequality holds by $\|\widetilde{\theta}_{1,h}-\theta_{h}^{*}\|_{\mathcal{H}_{1,h}}^{2}\leq\lambda\|\widetilde{\theta}_{1,h}-\theta_{h}^{*}\|_{2}^{2}\leq 4\lambda B$.
+Plugging $\zeta=2\eta$ finishes the proof.
+∎
+F.2 Proof of Lemma 2
+Proof.
+According to Lemma 11, we have
+$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}\leq$ $2\eta\left(\sum_{i=1}^{k}\ell_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}\ell_{i,h}(\widetilde{\theta}_{i+1,h})\right)+4\lambda B$
 $+12\sqrt{2}B\eta\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{2}^{2}-\sum_{i=1}^{k}\lVert\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\rVert_{\mathcal{H}_{i,h}}^{2}.$
 We bound the right-hand side of the above lemma separately in the following.
-First, we bound the first term. Define the softmax function as follows.
-$[\sigma_{k,h}(z)]_{s}=\frac{\exp([z])_{s}}{1+\sum_{s\in\dot{\mathcal{S}}_{k,h}}\exp([z])_{s}},\quad\forall s\in\mathcal{S}_{k,h}.$
-Then, the loss function in (2) can be rewritten as
-$f_{k,h}(z_{k,h},y_{k,h})=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\operatorname{\mathds{1}}[y_{k,h}^{s^{\prime}}=1]\log\left(\frac{1}{[\sigma_{k,h}(z_{k,h})]_{s^{\prime}}}\right).$
-Define a pseudo-inverse function of $\sigma_{k,h}(\cdot)$ as
-$[\sigma_{k,h}^{-1}(p)]_{s^{\prime}}=\log\left(\frac{[p]_{s^{\prime}}}{1-\lVert p\rVert_{1}}\right),\forall p\in\{p\in[0,1]^{N_{k,h}}\mid\lVert p\rVert_{1}<1\}.$
-Then, we decompose the first term as follows.
-$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})$
-=\ = $\underbrace{\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})}_{(a)}+\underbrace{\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})}_{(b)}$
+The most challenging part is to bound the term $\sum_{i=1}^{k}\ell_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}\ell_{i,h}(\widetilde{\theta}_{i+1,h})$.
+At first glance, this term appears straightforward to control, as it can be observed that $\theta_{h}^{*}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\bar{\ell}_{h}(\theta)\triangleq\mathbb{E}{y_{i,h}}[\ell_{i,h}(\theta)]$, where the instantaneous loss $\ell_{i,h}(\theta)$ serves as an empirical observation of ¯$\bar{\ell}_{h}(\theta)$.
+Consequently, the loss gap term seemingly can be bounded using appropriate concentration results.
+However, a caveat lies in the fact that the update of the estimator $\widetilde{\theta}_{i+1,h}$ depends on the information $\ell_{i,h}$, or more precisely $y_{i,h}$, making it difficult to directly apply such concentration results.
+To address this issue, we decompose the loss gap into two components by introducing an intermediate term.
+Specifically, we define the softmax function as $[\sigma_{k,h}(z)]_{s}=\frac{\exp([z])_{s}}{1+\sum_{s\in\dot{\mathcal{S}}_{k,h}}\exp([z])_{s}},\forall s\in\mathcal{S}_{k,h}$.
+Using this definition, the loss function can be rewritten as:
+$\ell_{k,h}(z_{k,h},y_{k,h})=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\operatorname{\mathds{1}}[y_{k,h}^{s^{\prime}}=1]\log\left(\frac{1}{[\sigma_{k,h}(z_{k,h})]_{s^{\prime}}}\right).$
+Define a pseudo-inverse function of $\sigma_{k,h}(\cdot)$ as $[\sigma_{k,h}^{-1}(p)]_{s^{\prime}}=\log\left(\frac{[p]_{s^{\prime}}}{1-\lVert p\rVert_{1}}\right),\forall p\in\{p\in[0,1]^{S_{k,h}}\mid\lVert p\rVert_{1}<1\}$.
+Then, the loss gap term can be decomposed into two parts as follows.
+$\sum_{i=1}^{k}\left(\ell_{i,h}(\theta_{h}^{*})-\ell_{i,h}(\widetilde{\theta}_{i+1,h})\right)$
+=\ = $\underbrace{\sum_{i=1}^{k}\left(\ell_{i,h}(\theta_{h}^{*})-\ell_{i,h}(z_{i,h},y_{i,h})\right)}_{\mathtt{term~{}(a)}}+\underbrace{\sum_{i=1}^{k}\left(\ell_{i,h}(z_{i,h},y_{i,h})-\ell_{i,h}(\widetilde{\theta}_{i+1,h})\right)}_{\mathtt{term~{}(b)}}$
 where $z_{k,h}=\sigma^{-1}_{k,h}(\mathbb{E}_{\theta\sim P_{k,h}}[\sigma_{k,h}((\phi_{k,h}^{s^{\prime}})^{\top}\theta)_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}])$, $P_{k,h}\triangleq\mathcal{N}(\widetilde{\theta}_{k,h},(1+c\mathcal{H}_{k,h}^{-1}))$ is the Gaussian distribution with mean $\widetilde{\theta}_{k,h}$ and covariance $(1+c\mathcal{H}_{k,h}^{-1})$ where $c$ is a constant to be specified later.
-First, we bound the term $(a)$ as follows.
-Lemma 13 (Lemma E.2 of Lee and Oh [2024]).
-Let $\delta\in(0,1]$ and $\lambda\geq 1$, for all $k\in[K],h\in[H]$, with probability at least $1-\delta$, we have
-$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})$
+The design of the intermediate term was originally proposed by Zhang and Sugiyama [2023] in their study of the multiclass logistic bandit problem and was subsequently applied to the multinomial logit bandits problem by Lee and Oh [2024].
+Notably, the intermediate loss is independent of the information contained in $y_{i,h}$, enabling the application of concentration results.
+Specifically, based on Lemma F.2 and Lemma F.3 of Lee and Oh [2024], we obtain the following upper bounds for them.
+For term (a), let $\delta\in(0,1]$ and $\lambda\geq\max\{2,72cd\}$, for all $k\in[K],h\in[H]$, with probability at least $1-\delta$, we have
+$\mathtt{term~{}(a)}$
 […]
-Then, we bound the term $(b)$ as follows.
-Lemma 14 (Lemma E.3 of Lee and Oh [2024]).
-For any $c\geq 0$, let $\lambda\geq\max\{2,72cd\}$, then for all $k\in[K],h\in[H]$, we have
-$\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})\leq\frac{1}{2c}\sum_{i=1}^{k}\left\|\widetilde{\theta}_{i,h}-\theta_{i+1,h}\right\|_{\mathcal{H}_{i,h}}^{2}+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right).$
-Now, we are ready to prove Lemma 3. Combining Lemma 12, Lemma 13, and Lemma 14, we have
+For term (b), for all $k\in[K],h\in[H]$, we have
+$\mathtt{term~{}(b)}\leq\frac{1}{2c}\sum_{i=1}^{k}\left\|\widetilde{\theta}_{i,h}-\theta_{i+1,h}\right\|_{\mathcal{H}_{i,h}}^{2}+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right).$
+Combining term (a) and term (b), we have
 $\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}$
@@
 This finishes the proof. ∎
-Appendix F Proof of Lemma 4
+F.3 Proof of Lemma 3
 Proof.
-By the second-order Taylor expansion, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
-${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}$
-=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})^{\top}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})$
-The gradient of $p_{s,a}(\theta)$ is given by
-$\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}.$
-For the first-order term, we have
-$\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})^{\top}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})$
-=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})$
-[…]
-[…]
-[…]
-[…]
-where in the first inequality, we denote $\mathcal{S}_{h,s,a}^{+}$ as the subset of $\mathcal{S}_{h,s,a}$ such that $\phi_{s,a}^{s^{\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})$ is non-negative, the second inequality holds by the Holder’s inequality, and the third inequality is by the confidence set in Lemma 3.
-For the second-order term, we define $u_{s,a}^{s^{\prime}}(\theta)=\phi_{s,a}^{s^{\prime}}\theta$, and $p_{s,a}^{s^{\prime}}(u)=\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}}\exp(u_{s,a}^{s^{\prime\prime}})}$, further define
-$F(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})},\widetilde{F}(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})V(s^{\prime})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})}.$
-Then, we have
-$\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})$
-=\ = $\frac{1}{2}\big{(}u(\widehat{\theta}_{k,h})-u(\theta_{h}^{*})\big{)}^{\top}\nabla^{2}\widetilde{F}(u(\bar{\theta}))\big{(}u(\widehat{\theta}_{k,h})-u(\theta_{h}^{*})\big{)}$
-=\ = $\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\big{(}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{)}^{\top}\frac{\partial^{2}\widetilde{F}(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{)}$
-[…]
-where the inequality holds by $V(s)\in[0,H],\forall s$.
-According to Lemma 18, we have (omit the subscript $\mathcal{S}_{h,s,a}$ for simplicity):
-$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
-[…]
-$+\frac{3H}{2}\sum_{s^{\prime}}\big{(}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta})).$ (30)
-To bound the first term, by applying the AM-GM inequality, we obtain
-$H\sum_{s^{\prime}}\sum_{s^{\prime\prime}\neq s^{\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
-[…]
-[…]
-$\hskip 56.9055pt+\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}$
-[…]
-Plugging (31) into (30), we have
-$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
-[…]
-=\ = $\frac{5H}{2}\sum_{s^{\prime}}\big{(}(\phi_{s,a}^{s^{\prime}})^{\top}(\widehat{\theta}_{k,h}-\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta}))$
-[…]
-where the last inequality holds by Lemma 3. Combining (29) and (32) finishes the proof. ∎
-Appendix G Proof of Theorem 2
+Lemma 3 follows directly by substituting the confidence set $\widehat{\mathcal{C}}_{k,h}$ defined in Lemma 2, into Lemma 7. This finishes the proof. ∎
+F.4 Proof of Theorem 2
 Proof.
@@
 […]
-=\ = $\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(c)}+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(d)}$
-$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(e)}.$
+=\ = $\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{\mathtt{term~{}(c)}}+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{\mathtt{term~{}(d)}}$
+$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{\mathtt{term~{}(e)}}.$
 We bound these terms separately in the following.
@@
 […]
-=\ = $\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\Big{\lvert}\Big{(}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime}}-p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime\prime}}\Big{)}^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\Big{\rvert}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
+=\ = $\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\Big{\lvert}\Big{(}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime}}-p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime\prime}}\Big{)}^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\Big{\rvert}\cdot\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
 […]
-$\hskip 28.45274pt+\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\big{\lvert}(\phi_{k,h}^{s^{\prime\prime\prime}})^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\big{\rvert}$
+$\hskip 28.45274pt+\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\cdot\big{\lvert}(\phi_{k,h}^{s^{\prime\prime\prime}})^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\big{\rvert}$
 […]
@@
 […]
-where and the fifth inequality is by Cauchy-Schwarz inequality and the fourth inequality is because by Lemma 17, since ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
-$\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}}\leq\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\widetilde{\mathcal{H}}_{k,h}}\leq 2\eta\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{\widetilde{\mathcal{H}}_{k,h}^{-1}}\leq\frac{2\eta}{\sqrt{\lambda}}\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2},$
-and since $g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
-$\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2}\leq\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}+\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}\leq 2\max_{s^{\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime}}\big{\rVert}_{2}\leq 2.$
+where and the fourth inequality is because by Lemma 15 and the fact ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
+$\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}}\leq\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\widetilde{\mathcal{H}}_{k,h}}\leq 2\eta\lVert\nabla\ell_{k,h}(\widetilde{\theta}_{k,h})\rVert_{\widetilde{\mathcal{H}}_{k,h}^{-1}}\leq\frac{2\eta}{\sqrt{\lambda}}\lVert\nabla\ell_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2},$
+and since $\nabla\ell_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
+$\lVert\nabla\ell_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2}\leq\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}+\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}\leq 2\max_{s^{\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime}}\big{\rVert}_{2}\leq 2.$
 Therefore, we have
-$\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\frac{4\eta}{\sqrt{\lambda}}\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}$
+$\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\frac{8\eta}{\sqrt{\lambda}}\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}$
 […]
 […]
-where the last inequality holds by Lemma 6.
+where the last inequality holds by $\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right)$ in Lemma 6.
 For the term $(d)$, by similar analysis, we have
@@
 […]
-$\hskip 56.9055pt+\frac{2\eta}{\sqrt{\lambda}}\frac{\left(\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\right)^{2}+\left(\max_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\right)^{2}}{2}$
+$\hskip 56.9055pt+\frac{2\eta}{\sqrt{\lambda}}\left(\Big{(}\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\Big{)}^{2}+\Big{(}\max_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\Big{)}^{2}\right)$
 […]
@@
 […]
-where the last inequality holds by Lemma 6.
+where the last inequality holds by $\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right)$ in Lemma 6.
 Finally, we bound the term $(e)$ as follows.
@@
 $\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{fst}}$ $=H\widetilde{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
-$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{16\eta}{\kappa\sqrt{\lambda}}d\log(1+\frac{K}{d\lambda})\right)$ (36)
+$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{32\eta}{\kappa\sqrt{\lambda}}d\log\left(1+\frac{K}{d\lambda}\right)\right)$ (26)
 For the second-order term, by Lemma 6, we have
-$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}$ $=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}$
-$\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (37)
-where the last inequality holds by Lemma 6.
+$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (27)
+where the last inequality holds by $\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right)$ in Lemma 6.
 Combining (26) and (27), we have
 $\operatorname*{Reg}(K)$ $\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}(\epsilon_{k,h}^{\mathtt{fst}}+\epsilon_{k,h}^{\mathtt{snd}})+H\sqrt{2KH\log(2/\delta)}$
-$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{16\eta}{\kappa\sqrt{\lambda}}d\log(1+\frac{K}{d\lambda})\right)+\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right)$
-$\hskip 28.45274pt+H\sqrt{2KH\log\Big{(}\frac{2}{\delta}\Big{)}}$
-$\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1}\big{)}$
+$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{32\eta}{\kappa\sqrt{\lambda}}d\log\left(1+\frac{K}{d\lambda}\right)\right)+\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right)$
+$\hskip 28.45274pt+H\sqrt{2KH\log\frac{2}{\delta}}$
+$\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1}\big{)}.$
 This finishes the proof. ∎
-Appendix H Proof of Theorem 3
+Appendix G Omitted Proofs for Section 6
+G.1 Proof of Theorem 3
 Proof.
-Our proof is similar to the proof of the lower bound for adversarial linear mixture MDPs with the unknown transition in Zhao et al.
-[2023].
-We prove this lemma by reducing the MNL mixture MDP problem to a sequence of binary logistic bandit problems.
-Specifically, we use each three layers to construct a block.
+Our proof is similar to adversarial linear mixture MDPs with the unknown transition in Zhao et al. [2023]. We prove this lemma by reducing MNL mixture MDPs to a sequence of logistic bandits.
+We use each three layers to construct a block.
 Note that the third layer of block $i$ is also the first layer of block $i+1$ and hence there are total $H/2$ blocks.
@@
 Further, for any $a\in\mathcal{A}$, let $\phi(s_{i,1},a,s_{i,2})=0$.
-The probability of transferring to state $s_{i,2}^{*}$ when executing action $a$ at state $s_{i,1}$ is $\rho_{a}$, in particular,
-$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a))^{\top}\theta_{i,1}^{*})}=\rho_{a},$
-$\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)$ $=\frac{1}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=1-\rho_{a}.$
-For the second layer, the MDP instance satisfies that $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$.
-The reward function satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second layer for all $a\in\mathcal{A}$.
-Consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from Bernoulli distributed with mean
-$\mu(x^{\top}\theta^{*})=\frac{1}{1+\exp(x^{\top}\theta^{*})}.$
-By this configuration, we can see that learning in this block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$.
-It is also clear that the optimal policy at state $s_{i,1}$ is to select action $a_{i,1}^{*}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$.
+The transition probability is defined as follows:
+$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=\rho_{a},\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)=1-\rho_{a}.$
+For the second layer, it satisfies $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$.
+The reward satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second for all $a\in\mathcal{A}$.
+Then, consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from the Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(x^{\top}\theta^{*}))^{-1}$.
+By this configuration, we can see that learning in each block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$.
 Thus, learning this MNL mixture MDP equals to learning $H/2$ logistic bandit problems.
@@
 ∎
-Appendix I Supporting Lemmas
+Appendix H Supporting Lemmas
 In this section, we provide several supporting lemmas used in the proofs of the main results.
@@
 For any $t\geq 1$, define
-$V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top}\quad S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}.$
-Then, for any $\delta>0$, with probability at least $1-\delta$, for all $t\geq 0$,
-$\left\|S_{t}\right\|_{V_{t}^{-1}}^{2}\leq 2R^{2}\log\left(\frac{\operatorname{det}\left(V_{t}\right)^{1/2}\operatorname{det}(V)^{-1/2}}{\delta}\right).$
+$V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top},\quad S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}.$
+Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all $t\geq 1$,
+$\left\|S_{t}\right\|_{V_{t}^{-1}}\leq R\sqrt{2\log\bigg{(}\frac{\operatorname{det}\left(V_{t}\right)^{1/2}\operatorname{det}(V)^{-1/2}}{\delta}\bigg{)}}.$
+Lemma 13 (Périvier and Goyal [2022, Theorem 4]).
+Let $\{\mathcal{F}_{t}\}_{t=0}^{\infty}$ be a filtration.
+Let $\{\delta_{t}\}_{t=1}^{\infty}$ be an $\mathbb{R}^{N}$-valued stochastic process such that $\delta_{t}$ is $\mathcal{F}_{t}$-measurable one-hot vector.
+Furthermore, assume $\mathbb{E}[\delta_{t}|\mathcal{F}_{t-1}]=p_{t}$ and define $\varepsilon_{t}=p_{t}-\delta_{t}$.
+Let $\{X_{t}\}_{t=1}^{\infty}$ be a sequence of $\mathbb{R}^{N\times d}$-valued stochastic process such that $X_{t}$ is $\mathcal{F}_{t-1}$-measurable and $\lVert X_{t,i}\rVert_{2}\leq 1,\forall i\in[N]$.
+Let $\{\lambda_{t}\}_{t=1}^{\infty}$ be a sequence of non-negative scalars.
+Define
+$H_{t}=\sum_{i=1}^{t-1}\left(\sum_{j=1}^{N}p_{j}X_{i,j}X_{i,j}^{\top}-\sum_{j=1}^{N}\sum_{k=1}^{N}p_{j}p_{k}X_{i,j}X_{i,k}^{\top}\right)+\lambda_{t}I_{d},\quad S_{t}=\sum_{i=1}^{t-1}\sum_{j=1}^{N}\varepsilon_{i,j}X_{i,j}.$
+Then, for any $\zeta\in(0,1)$, with probability at least $1-\zeta$, for all $t\geq 1$,
+$\lVert S_{t}\rVert_{H_{t}^{-1}}\leq\frac{\sqrt{\lambda_{t}}}{4}+\frac{4}{\sqrt{\lambda_{t}}}\log\left(\frac{2^{d}\det\left(H_{t}\right)^{\frac{1}{2}}\lambda_{t}^{-\frac{d}{2}}}{\zeta}\right).$
 Lemma 14 (Abbasi-Yadkori et al. [2011, Lemma 10]).
@@
 Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$.
-Then, for any $1\leq s\leq t$, we have
-$\det(V_{t})\leq\left(\lambda+\frac{tL^{2}}{d}\right)^{d}.$
+Then, we have $\det(V_{t})\leq\left(\lambda+{tL^{2}}/{d}\right)^{d}$.
 Lemma 15 (Orabona [2019, Lemma 6.9]).
-Let $Z$ be a positive define matrix and $\mathcal{X}$ be a convex set,
-$\widetilde{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\langle\mathbf{g},\mathbf{x}\rangle+\frac{1}{2\eta}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{Z}^{2}.$
-Then, for all $\mathbf{x}^{*}\in\mathcal{X}$, we have
-$\lVert\widetilde{\mathbf{x}}-\widehat{\mathbf{x}}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}.$
-and
-$\langle\mathbf{g},\widehat{\mathbf{x}}-\mathbf{x}^{*}\rangle\leq\frac{\eta}{2}\lVert\mathbf{g}\rVert_{Z^{-1}}^{2}+\frac{1}{2\eta}\lVert\widehat{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}-\frac{1}{2\eta}\lVert\widetilde{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}.$
+Let $Z$ be a positive define matrix and $\mathcal{W}$ be a convex set, define $\mathbf{w}_{t+1}$ as the solution of
+$\mathbf{w}_{t+1}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{W}}\Big{\{}\langle\mathbf{g},\mathbf{w}\rangle+\frac{1}{2\eta}\lVert\mathbf{w}-\mathbf{w}_{t}\rVert_{Z}^{2}\Big{\}}.$
+Then we have
+$\lVert\mathbf{w}_{t+1}-\mathbf{w}_{t}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}.$
+Lemma 16 (Campolongo and Orabona [2020, Proposition 4.1]).
+Define $\mathbf{w}_{t+1}$ as the solution of
+$\mathbf{w}_{t+1}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{V}}\big{\{}\eta\ell_{t}(\mathbf{w})+\mathcal{D}_{\psi}\left(\mathbf{w},\mathbf{w}_{t}\right)\big{\}},$
+[…]
+$\left\langle\eta_{t}\mathbf{g}_{t}^{\prime},\mathbf{w}_{t+1}-\mathbf{u}\right\rangle\leq\left\langle\nabla\psi\left(\mathbf{w}_{t}\right)-\nabla\psi\left(\mathbf{w}_{t+1}\right),\mathbf{w}_{t+1}-\mathbf{u}\right\rangle$
+for any $\mathbf{u}\in\mathcal{W}$.
 Lemma 17 (Lee and Oh [2024, Lemma D.3]).